### v1

# On the generalized coloring numbers11footnotemark: 1

Abstract
The coloring number $\operatorname{col}(G)$ of a graph $G$, which is equal to the degeneracy of $G$ plus one, provides a very useful measure for the uniform sparsity of $G$. The coloring number is generalized by three series of measures, the generalized coloring numbers. These are the $r$-admissibility $\operatorname{adm}_{r}(G)$, the strong $r$-coloring number $\operatorname{col}_{r}(G)$ and the weak $r$-coloring number $\operatorname{wcol}_{r}(G)$, where $r$ is an integer parameter. The generalized coloring numbers measure the edge density of bounded-depth minors and thereby provide an even more uniform measure of sparsity of graphs. They have found many applications in graph theory and in particular play a key role in the theory of bounded expansion and nowhere dense graph classes introduced by Nešetřil and Ossona de Mendez. We overview combinatorial and algorithmic applications of the generalized coloring numbers, emphasizing new developments in this area. We also present a simple proof for the existence of uniform orders and improve known bounds, e.g., for the weak coloring numbers on graphs with excluded topological minors.
keywords:
Generalized coloring numbers, graph sparsity, nowhere denseness, bounded expansion.
journal: arXiv
1 Introduction
The notions of bounded expansion and nowhere dense graph classes were introduced by Nešetřil and Ossona de Mendez as models for uniformly sparse graph classes [112, 116]. These are abstract and robust notions of sparsity, and many of the commonly studied sparse graph classes have bounded expansion or at least are nowhere dense. Examples include the class of all planar graphs, classes excluding a minor, classes with bounded degree, classes excluding a topological minor, classes with locally bounded treewidth, and many more. Bounded expansion and nowhere dense graph classes were originally defined by restricting the edge densities of bounded-depth minors that may appear in these classes [^2] A graph $H$ is a depth-$r$ minor of $G$ if there is a map $M$ that assigns to every vertex $v\in V(H)$ a connected subgraph $M(v)\subseteq G$ of radius at most $r$ and to every edge $e\in E(H)$ an edge $M(e)\in E(G)$ such that $M(u)$ and $M(v)$ are vertex disjoint for distinct vertices $u,v\in V(H)$, and if $e=uv\in E(H)$, then $M(e)=u^{\prime}v^{\prime}\in E(G)$ for vertices $u^{\prime}\in M(u)$ and $v^{\prime}\in M(v)$, see Figure 2. A class of graphs has bounded expansion (is nowhere dense) if the depth-$r$ minors of its members have edge density (clique number) bounded by $f(r)$ for some function $f$.. It turns out, however, that they can equivalently be characterized in many different ways, each characterization highlighting specific combinatorial properties and leading to specific algorithmic tools.
One particular useful characterization of bounded expansion and nowhere dense graph classes is in terms of the so-called generalized coloring numbers. Before introducing these, let us consider the classical coloring number. The coloring number $\operatorname{col}(G)$ of a graph $G$, first considered by Erdős and Hajnal, is the minimum integer $k$ such that there exists a linear order $\pi$ of the vertices of $G$ such that every vertex $v$ has back-degree at most $k-1$, i.e., at most $k-1$ neighbors $u$ with $u<_{\pi}v$ [59]. The coloring number of $G$ is one more than the degeneracy of $G$, which is the minimum integer $\ell$ such that every subgraph $H\subseteq G$ has a vertex of degree at most $\ell$, as observed by Lick and White [103]. To see this, assume first that $G$ has coloring number $k$ and let $H\subseteq G$ be a subgraph of $G$. Let $\pi$ be an order of $V(G)$ witnessing that $\operatorname{col}(G)\leq k$, and let $v$ be the largest vertex of $H$ with respect to $\pi$. All neighbors of $v$ in $H$ are hence smaller than $v$, and, by assumption, there are at most $k-1$ of them. Hence, as $H$ was chosen arbitrarily, $G$ is $k-1$ degenerate. Conversely, assume that the degeneracy of $G$ is equal to $\ell$. Observe that also every subgraph $H\subseteq G$ has degeneracy at most $\ell$ and that the statement is trivial for single vertex graphs. Now, if $G$ has more than one vertex, we can find an order witnessing that $\operatorname{col}(G)\leq\ell+1$ by taking a vertex $v$ of $G$ of degree at most $\ell$, finding inductively an order with back-degree at most $\ell$ for the graph $H=G-v$, and inserting $v$ into the order as the largest vertex, see Figure 1.
Figure 1: Graphs with coloring number $k$ can have vertices with arbitrarily large degree, however, the back-degree of every vertex is bounded by $k-1$. A simple greedy procedure leads to a proper vertex coloring.
An $\ell$-degenerate graph $G$ is uniformly sparse in the sense that every $n$-vertex subgraph $H\subseteq G$ has at most $\ell\cdot n$ edges. Hence, despite the name, which suggests a strong connection with the chromatic number $\chi$, the coloring number is a parameter that measures uniform sparsity of graphs. To its rescue, the coloring number does have a connection to the chromatic number, as a simple greedy coloring algorithm with the degeneracy order provides a bound for the chromatic number: $\chi(G)\leq\operatorname{col}(G)$. To see this, let $\pi$ be an order of $V(G)$ witnessing that $\operatorname{col}(G)\leq k$, and let $v$ be the largest vertex of $G$ with respect to $\pi$. By induction, we obtain a proper vertex coloring with at most $k$ colors of $G-v$. As $v$ has at most $k-1$ neighbors, there is a color that is not used by any of its neighbors and we can complete the coloring of $G-v$ to a coloring of $G$ with at most $k$ colors. If $G$ is chordal, this bound is tight, while bipartite graphs are $2$-colorable and can have arbitrarily large coloring number.
Several generalizations of the coloring number have been studied in the literature, including the arrangeability by Chen and Schelp [23] (connected to Ramsey problems), the game chromatic number by Faigle et al. [63], the game coloring number by Zhu [157], the rank by Kierstead [91] (related to the game chromatic number), and the admissibility of a graph by Kierstead and Trotter [92] (used to bound both the arrangeability and the game chromatic number). These parameters are yet generalized by the following three series of parameters, called the generalized coloring numbers. For a positive integer $r$, the strong $r$-coloring number $\operatorname{col}_{r}(G)$ and weak $r$-coloring number $\operatorname{wcol}_{r}(G)$ were introduced by Kierstead and Yang [93], and the $r$-admissibility $\operatorname{adm}_{r}(G)$ was introduced by Dvořák [50]. All of these generalizations rely on linear orders and certain back-connectivity properties (formal definition will be given below).
Intuitively, the generalized coloring numbers capture structural properties not only of subgraphs of a graph, as degeneracy does, but also of its bounded-depth minors. Bounded-depth minors are the key concept in the theory of bounded expansion and nowhere dense graph classes, introduced by Nešetřil and Ossona de Mendez. The (non trivial) connection between the generalized coloring numbers and classes with bounded expansion was made by Zhu [158], and this connection naturally extends to nowhere dense classes as observed by Nešetřil and Ossona de Mendez [116].
Figure 2: Graph $H$ (on the right) is a depth-$2$ minor of $G$ (on the left).
On the other hand, the strong $r$-coloring numbers can be seen as gradations between the coloring number $\operatorname{col}(G)$ and the treewidth $\operatorname{tw}(G)$ of $G$, and the weak $r$-coloring numbers can be seen as gradations between $\operatorname{col}(G)$ and the treedepth $\operatorname{td}(G)$ of $G$. Just as treewidth captures the global connectivity of a graph $G$, the generalized coloring numbers capture its local connectivity. By now, the generalized coloring numbers have found many applications in graph theory, and play a key role in the theory of bounded expansion and nowhere dense classes of graphs.
Bounded expansion and nowhere dense classes are intimately linked to first-order logic and many of the developments were driven by the question to solve the first-order model checking problem on these classes. Structural decompositions derived from the generalized coloring numbers lead to quantifier-elimination, and as a consequence to efficient model checking as shown by Dvořák, Král’ and Thomas [53], and enumeration and query counting as shown by Kazana and Segoufin [89], for first-order logic on classes with bounded expansion. Game characterizations and sparse neighborhood covers lead to tractable first-order model checking on nowhere dense classes as shown by Grohe et al. [77]. In fact, for monotone (that is, subgraph-closed) classes of graphs nowhere denseness turns out to be the dividing line between tractable and intractable first-order model checking. Hence, the classification of tractable first-order model checking on monotone graph classes is complete and leaves the question of a classification of tractability on hereditary graph classes.
The deep connections of nowhere dense classes to classical model theory (stability theory) are currently guiding an exciting development of dense but structurally sparse graph classes. Stability is one of the key notions from classical model theory (see the monumental work of Shelah [143]) and it turns out that nowhere dense classes are stable, as observed by Adler and Adler [3], building on work of Podewski and Ziegler [133]. This connection is highly fruitful and many notions from model theory can be appropriately translated to graph theory and turn out to be dense analogs of well known concepts from sparsity theory.
Another recent exciting development evolves around the graph product structure theory for planar graphs, initiated by Dujmović et al. [47], which was inspired by the techniques developed for the construction of low treedepth colorings for classes with excluded minors by Pilipczuk and Siebertz [127].
Figure 3: The map of the (structurally) sparse universe.
In this survey, we overview combinatorial and algorithmic applications of the generalized coloring numbers, emphasizing new developments in this area. We provide the proofs of a few key results that connect the relevant concepts and will allow the reader to obtain a deeper understanding. The first part of the survey deals with upper and lower bounds for the generalized coloring numbers in general and on restricted graph classes. We give background on graph theory and formally define the generalized coloring numbers in Section 2. We also give some background on first-order logic to later give an outlook on the theory of structural sparsity, however, this is not the focus of this survey. We present upper bounds for the generalized coloring numbers and make the connection to the theory of bounded expansion and nowhere dense graph classes in Section 3. We also present a simple proof for the existence of uniform orders in Section 3.3; the existence of such orders was first proved by Van den Heuvel and Kierstead [148]. In Section 4 we turn our attention to tree decomposable graphs and make the connection to the recently emerging graph product structure theory. We present lower bounds in Section 5.
In Section 6 we study structural decompositions of graphs, such as low treedepth decompositions and their generalizations. In Section 7 we consider game characterizations of bounded expansion and nowhere dense graph classes and highlight their connections to the generalized coloring numbers. We study further applications, such as neighborhood complexity, vc-density, sparse neighborhood covers, approximations and parameterized algorithms for the distance-$d$ dominating set problem in Section 8. Further selected topics are presented in Section 9. We give more background and references to the discussed concepts in the respective sections.
2 The generalized coloring numbers
Unless explicitly stated differently, all graphs in this paper are finite, undirected and simple. We follow the notation of Diestel’s textbook [34] and refer to it for all graph theoretic notations that are left undefined here. An order of the vertex set $V(G)$ of an $n$-vertex graph $G$ is a permutation $\pi=(v_{1},\ldots,v_{n})$. We say that $v_{i}$ is smaller than $v_{j}$ and write $v_{i}<_{\pi}v_{j}$ if $i<j$. We write $\Pi(G)$ for the set of all orders of $V(G)$. We write $\mathcal{P}(X)$ for the powerset of a set $X$.
2.1 Treewidth and the strong r-coloring numbers
The generalized coloring numbers $\operatorname{col}_{r}(G)$ and $\operatorname{wcol}_{r}(G)$ were introduced by Kierstead and Yang [93] as generalizations of the usual coloring number. Here, we take a different perspective and introduce the strong $r$-coloring number $\operatorname{col}_{r}(G)$ as a local version of treewidth, while the weak $r$-coloring number $\operatorname{wcol}_{r}(G)$ will be introduced in the next subsection as a local version of treedepth. This approach supports our view of these parameters as measures for the local connectivity properties of a graph, and sets the stage for the study of tree decomposable graphs.
Treewidth, as a measure of the global connectivity of a graph, plays an important role in graph theory. Treewidth and tree decompositions were introduced by Robertson and Seymour [136] as part of their graph minors project, however, several strongly related or equivalent notions were invented independently [10, 79]. We refer to [13, 81] for surveys on various characterizations of treewidth and equivalent parameters and to [14, 28] for surveys on algorithmic applications.
Definition 2.1.
A tree decomposition of a graph $G$ is a pair $\mathcal{T}=(T,\mathrm{Bag})$, where $T$ is a tree and $\mathrm{Bag}\colon V(T)\to\mathcal{P}(V(G))$ is a mapping that assigns each node $x$ of $T$ to its bag $\mathrm{Bag}(x)\subseteq V(G)$ so that the following conditions are satisfied:
  1. 1.
    For each $u\in V(G)$, the set $\{x\mid u\in\mathrm{Bag}(x)\}$ is non-empty and induces a connected subtree of $T$.
  2. 2.
    For every edge $uv\in E(G)$, there is a node $x\in V(T)$ such that $\{u,v\}\subseteq\mathrm{Bag}(x)$.
Let $\mathcal{T}=(T,\mathrm{Bag})$ be a tree decomposition of $G$. The width of $\mathcal{T}$ is the maximum bag size minus $1$, i.e., $\max_{x\in V(T)}|\mathrm{Bag}(x)|-1$. The treewidth of a graph $G$ is the minimum possible width of a tree decomposition of $G$. For $X\subseteq V(T)$ we let $\mathrm{Bag}(X)=\bigcup_{x\in X}\mathrm{Bag}(x)$.
Figure 4: A tree decomposition of a graph. No edge can cross the intersection of two bags, which form separators in $G$.
In the following, it will be convenient to assume that $T$ is a rooted tree. A rooted tree is a tree $T$ with a designated root vertex $r$. Assigning a root imposes the standard ancestor/descendant relation $\leq_{T}$ in $T$ where the root is the $\leq_{T}$ minimal element: a node $v$ is a descendant of all the nodes that appear on the unique path leading from $v$ to the root. Note that we treat a vertex as an ancestor and descendant of itself. The ancestors and descendants of  $v$ excluding $v$ are its true ancestors and true descendants, respectively. The parent of a non-root node is its closest true ancestor. A rooted forest $F$ is a disjoint union of rooted trees. We write $\leq_{F}$ for the partial order that coincides with $\leq_{T}$ on each rooted tree in $F$.
Tree decompositions capture the global connectivity properties of graphs, as made precise next.
Definition 2.2.
Let $G$ be a graph and let $U,W\subseteq V(G)$. A set $S\subseteq V(G)$ separates $U$ and $W$, if every path from a vertex of $U$ to a vertex of $W$ contains a vertex of $S$.
Lemma 2.3 ([136]).
Let $G$ be a graph and let $\mathcal{T}=(T,\mathrm{Bag})$ be a tree decomposition of $G$. For an edge $e=xy\in E(T)$ let $T_{1},T_{2}$ be the components of $T-e$. Then $\mathrm{Bag}(x)\cap\mathrm{Bag}(y)$ separates $\mathrm{Bag}(T_{1})$ and $\mathrm{Bag}(T_{2})$.
Let $\mathcal{T}=(T,\mathrm{Bag})$ be a tree decomposition of $G$. This decomposition induces a quasi-order on $V(G)$ as follows. For $v\in V(G)$, let $x(u)$ be the unique $\leq_{T}$-minimal node of $T$ with $v\in\mathrm{Bag}(x)$. Define the quasi-order $\leq_{\mathcal{T}}$ on $V(G)$ by $u\leq_{\mathcal{T}}v$ if and only if $x(u)\leq_{T}x(v)$. Any linearization of $\leq_{\mathcal{T}}$ has very nice properties, which we will exploit later. In fact, we can associate a width to any linear order of $V(G)$ and thereby provide a characterization of treewidth.
Definition 2.4.
Let $G$ be a graph. A vertex $v\in V(G)$ is simplicial if its neighborhood induces a clique in $G$. An order $\pi=(v_{1},\ldots,v_{n})$ of $V(G)$ is a perfect elimination order if for every $1\leq i\leq n$, $v_{i}$ is simplicial in $G[V_{i}]$, where $V_{i}=\{v_{1},\ldots,v_{i}\}$.
Definition 2.5.
A graph $G$ is chordal or triangulated if all its cycles with four or more vertices have a chord, that is, an edge that is not part of the cycle but connects two vertices of the cycle. Equivalently, $G$ is chordal if every induced cycle in $G$ has exactly three vertices. A triangulation of $G$ is a chordal supergraph $H$ of $G$ on the same vertex set as $G$.
Lemma 2.6 ([64]).
A graph $G$ is chordal if and only if it admits a perfect elimination order.
Let $G$ be a graph and let $\pi=(v_{1},\ldots,v_{n})$ be an order of $V(G)$. We construct the graph $G^{\pi}_{\textit{fill}}$, called the fill-in graph of $G$ with respect to the order $\pi$, by the following fill-in procedure. We let $G_{n}\coloneqq G$. Now, for $i=n-1,\ldots,1$, assume that the graph $G_{i+1}$ has been constructed. We define $G_{i}$ as the graph obtained from $G_{i+1}$ by adding all edges between non-adjacent neighbors of $v_{i+1}$ that are smaller than $v_{i}$. Finally, we let $G^{\pi}_{\textit{fill}}\coloneqq G_{1}$. It is immediate from the definition that $\pi$ is a perfect elimination order of $G^{\pi}_{\textit{fill}}$.
Figure 5: An elimination order and fill-in graph. The neighbors of $v_{i}$ that are smaller than $v_{i}$ are marked by red edges in $G_{i-1}$, the fill-in edges that are added between these neighbors are marked by green edges.
Recall that the back-degree of a vertex $v$ with respect to an order $\pi$ is the number of neighbors of $v$ that are smaller than $v$ with respect to $\pi$.
Definition 2.7.
Let $G$ be a graph and let $\pi=(v_{1},\ldots,v_{n})$ be an order of $V(G)$. The width of $\pi$ on $G$ is the maximum back-degree of any vertex $v$ in the fill-in graph $G^{\pi}_{\textit{fill}}$ (with respect to the order $\pi$).
Lemma 2.8 (see e.g. [13]).
Let $G$ be a graph. The treewidth $\operatorname{tw}(G)$ of $G$ is equal to the minimum width over all orders $\pi$ of $V(G)$.
An easy induction shows that for $1\leq i<j\leq n$, there is an edge between $v_{i}$ and $v_{j}$ in the fill-in graph $G^{\pi}_{\textit{fill}}$ if and only if there exists a path $P$ between $v_{i}$ and $v_{j}$ such that $v_{i}$ is the smallest vertex on $P$ and all internal vertices of $P$ are larger than $v_{j}$ with respect to $\pi$. We say that $v_{i}$ is strongly reachable from $v_{j}$. For convenience, we also consider each vertex to be strongly reachable from itself and we write $\operatorname{SReach}[G,\pi,v]$ for the set of vertices that are strongly reachable from $v$. The above observation is the motivation for the following definition of local strong reachability.
Definition 2.9.
Let $G$ be a graph and $r$ a positive integer. Let $\pi$ be a linear order of $V(G)$. We say that a vertex $u\in V(G)$ is strongly $r$-reachable with respect to $\pi$ from a vertex $v\in V(G)$ if $u\leq_{\pi}v$ and there exists a path $P$ (possibly of length 00) between $u$ and $v$ of length at most $r$ with $w>_{\pi}v$ for all internal vertices $w\in V(P)$. The set of vertices strongly $r$-reachable by $v$ with respect to the order $\pi$ is denoted $\operatorname{SReach}_{r}[G,\pi,v]$. We define
$\operatorname{col}_{r}(G,\pi)\coloneqq\max_{v\in V(G)}|\operatorname{SReach}_{r}[G,\pi,v]|,$
and the strong $r$-coloring number $\operatorname{col}_{r}(G)$ as
$\operatorname{col}_{r}(G)\coloneqq\min_{\pi\in\Pi(G)}\max_{v\in V(G)}|\operatorname{SReach}_{r}[G,\pi,v]|.$
Figure 6: Vertices $v_{3}$ and $v_{4}$ strongly $2$-reach vertex $v_{1}$. This is witnessed by the path along $v_{5}$.
Note that since we allow the path $P$ to be of length 00 every vertex strongly $r$-reaches itself. It is immediate from the definitions that
$\operatorname{col}(G)=\operatorname{col}_{1}(G)\leq\operatorname{col}_{2}(G)\leq\ldots\leq\operatorname{col}_{n}(G)=\operatorname{tw}(G)+1.$
Hence, the strong $r$-coloring numbers can be seen as gradations between the coloring number $\operatorname{col}(G)$ and the treewidth $\operatorname{tw}(G)$ of $G$. As such, the strong $r$-coloring numbers capture local separation properties of $G$, and many applications are based on lemmas similar to the following.
Lemma 2.10.
Let $G$ be a graph, let $\pi$ be an order of $V(G)$, and let $r$ be a positive integer. Let $u,v\in V(G)$ such that $u<_{\pi}v$. Then every path $P$ of length at most $r$ connecting $u$ and $v$ intersects $\operatorname{SReach}_{r}[G,\pi,v]\setminus\{v\}$.
Observe that if $u\in\operatorname{SReach}_{r}[G,\pi,v]$, then the statement of the lemma is trivial.
Proof.
Let $P$ be any path of length at most $r$ connecting $u$ and $v$. Traverse $P$ starting at the endpoint $v$. Then the first vertex in this traversal of $P$ that is strictly smaller than $v$ lies in $\operatorname{SReach}_{r}[G,\pi,v]\setminus\{v\}$. ∎
There is some ambiguity in the literature on whether a vertex $v$ itself should be included in the set $\operatorname{SReach}_{r}[G,\pi,v]$. We have chosen to include it as some arguments become a bit nicer. In fact, some people argue that one should not subtract $1$ in the definition of treewidth.
2.2 Treedepth and the weak r-coloring numbers
The notion of treedepth was introduced by Nešetřil and Ossona de Mendez in [113], and again, equivalent notions were studied before under different names. We refer to [111] for a discussion on various equivalent parameters. Where intuitively treewidth measures the similarity of a graph with a tree, treedepth measures the similarity of a graph with a star.
The depth of a vertex $v$ in a rooted forest $F$ is the number of vertices on the path from $v$ to the root (of the tree to which $v$ belongs). The depth of $F$ is the maximum depth of the vertices of $F$.
Definition 2.11.
Let $G$ be a graph. The treedepth $\operatorname{td}(G)$ of $G$ is the minimum depth of a rooted forest $F$ on the same vertex set as $G$ such that whenever $uv\in E(G)$, then $u\leq_{F}v$ or $v\leq_{F}u$.
Figure 7: The treedepth of a path $P_{n}$ with $n$ vertices is $\lceil\log_{2}(n+1)\rceil$.
Similar to the width of an order $\pi$, we can define its depth via a corresponding notion of reachability as follows.
Definition 2.12.
Let $G$ be a graph and let $\pi$ be an order of $V(G)$. We say that a vertex $u\in V(G)$ is weakly reachable with respect to $\pi$ from a vertex $v\in V(G)$ if $u\leq_{\pi}v$ and there exists a path $P$ (possibly of length 00) between $u$ and $v$ with $w>_{\pi}u$ for all internal vertices $w\in V(P)$. We write $\operatorname{WReach}[G,\pi,v]$ for the set of vertices that are weakly reachable from $v$. The depth of $\pi$ on $G$ is the maximum over all vertices $v$ of $G$ of $|\operatorname{WReach}[G,\pi,v]|$.
Lemma 2.13 (see e.g. [111], Lemma 6.5).
Let $G$ be a graph. The treedepth of $G$ is equal to the minimum depth over all orders $\pi$ of $V(G)$.
And again we can naturally define a local version of weak reachability.
Definition 2.14.
Let $G$ be a graph and $r$ a positive integer. Let $\pi$ be a linear order of $V(G)$. We say that a vertex $u\in V(G)$ is weakly $r$-reachable with respect to $\pi$ from a vertex $v\in V(G)$ if $u\leq_{\pi}v$ and there exists a path $P$ (possibly of length 00) between $u$ and $v$ of length at most $r$ with $w>_{\pi}u$ for all internal vertices $w\in V(P)$. The set of vertices weakly $r$-reachable by $v$ with respect to the order $\pi$ is denoted $\operatorname{WReach}_{r}[G,\pi,v]$. We define
$\operatorname{wcol}_{r}(G,\pi)\coloneqq\max_{v\in V(G)}|\operatorname{WReach}_{r}[G,\pi,v]|,$
and the weak $r$-coloring number $\operatorname{wcol}_{r}(G)$ as
$\operatorname{wcol}_{r}(G)\coloneqq\min_{\pi\in\Pi(G)}\max_{v\in V(G)}|\operatorname{WReach}_{r}[G,\pi,v]|.$
Figure 8: Vertex $v$ weakly $4$-reaches the vertices $u_{1},u_{2},u_{3}$ (and itself).
Note that again we also consider the path of length 00 in the definition of weak reachability, hence, each vertex is weakly reachable from itself. And again, it is immediate from the definitions that
$\operatorname{col}(G)=\operatorname{wcol}_{1}(G)\leq\operatorname{wcol}_{2}(G)\leq\ldots\leq\operatorname{wcol}_{n}(G)=\operatorname{td}(G).$
Hence, the weak $r$-coloring numbers can be seen as gradations between the coloring number $\operatorname{col}(G)$ and the treedepth $\operatorname{td}(G)$ of $G$. The weak $r$-coloring numbers capture local separation properties of $G$ as follows.
Lemma 2.15.
Let $G$ be a graph, let $\pi$ be an order of $V(G)$ and let $r$ be a positive integer. Let $u,v\in V(G)$ such that $u<_{\pi}v$. Then every path $P$ of length at most $r$ connecting $u$ and $v$ intersects $\operatorname{WReach}_{r}[G,\pi,v]\cap\operatorname{WReach}_{r}[G,\pi,u]$.
Proof.
Let $P$ be a path of length at most $r$ connecting $u$ and $v$. Then the minimum vertex of $P$ lies both in $\operatorname{WReach}_{r}[G,\pi,v]$ and in $\operatorname{WReach}_{r}[G,\pi,u]$. ∎
And again, there is some ambiguity in the literature on whether a vertex $v$ itself should be included in the set $\operatorname{WReach}_{r}[G,\pi,v]$ and we have chosen to include it.
For the strong $r$-reachability, we follow paths that stop at the first point smaller than their starting vertex. For the weak $r$-reachability, we consider paths that can make up to $r$ hops to smaller and smaller vertices. The following gradation between the weak and strong coloring numbers was introduced by Cortés et al. in [25].
Definition 2.16.
Let $G$ be a graph and $r$ a positive integer. Let $\pi$ be a linear order of $V(G)$. We say that a vertex $u\in V(G)$ is $\ell$-hop $r$-reachable with respect to $\pi$ from a vertex $v\in V(G)$ if $u\leq_{\pi}v$ and there exists a path $P=v_{0}v_{1}\ldots v_{s}$ with $v_{0}=v$, $v_{s}=u$ for some $s\leq r$ such that $u<_{\pi}v_{i-1}$ for every $i\in[s]$ and such that |$|\{j\leq s\mid v_{j}<_{\pi}v_{i-1}$ for every $i\leq j\}|\leq\ell$. The set of vertices that are $\ell$-hop $r$-reachable from $v$ with respect to the order $\pi$ is denoted $\operatorname{GReach}_{\ell,r}[G,\pi,v]$. We define the $\ell$-hop $r$-coloring number of $\pi$ as
$\operatorname{gcol}_{\ell,r}(G,\pi)\coloneqq\max_{v\in V(G)}|\operatorname{GReach}_{\ell,r}[G,\pi,v]|,$
and the $\ell$-hop $r$-coloring number $\operatorname{gcol}_{\ell,r}(G)$ as
$\operatorname{gcol}_{\ell,r}(G)\coloneqq\min_{\pi\in\Pi(G)}\max_{v\in V(G)}|\operatorname{GReach}_{\ell,r}[G,\pi,v]|.$
Observe that $\operatorname{col}_{r}(G)=\operatorname{gcol}_{1,r}(G)$ and $\operatorname{wcol}_{r}(G)=\operatorname{gcol}_{r,r}(G)$. The paths from $\operatorname{gcol}{\ell,r}(G)$ can be split into shorter paths whenever they reach a local minimum with respect to the order. In particular,
$\operatorname{gcol}_{\ell+\ell^{\prime},r+r^{\prime}}(G)\leq\operatorname{gcol}_{\ell,r}(G)\cdot\operatorname{gcol}_{\ell^{\prime},r^{\prime}}(G).$
2.3 The r-admissibility
Finally, we define one last related measure, the $r$-admissibility of $G$, which was introduced by Dvořák [50] as a generalization of the admissibility [92].
Definition 2.17.
Let $G$ be a graph, $v\in V(G)$ and $A\subseteq V(G)$. A $v$-$A$ fan is a set of paths $P_{1},\ldots P_{k}$ with one endpoint in $v$ and the other endpoint in $A$ and which are internally vertex disjoint from $A$ such that $V(P_{i})\cap V(P_{j})=\{v\}$ for all $i\neq j$.
Note that if $v\in A$, then we may have the path of length 00 consisting only of $v$ in a $v$-$A$ fan. We remark that this differs from the standard definition of $v$-$A$ fans, which requires paths with one endpoint in $v$ and the other endpoint in $A\setminus\{v\}$. We have chosen this definition for consistency with the strong and weak coloring numbers, where each vertex is reachable from itself.
Definition 2.18.
Let $G$ be a graph and let $\pi=(v_{1},\ldots,v_{n})$ be a linear order of $V(G)$. For $1\leq i\leq n$, we write $V_{i}$ for the set $\{v_{1},\ldots,v_{i}\}$. For $1\leq i\leq n$, let $\operatorname{adm}_{\infty}[G,\pi,v_{i}]$ be the maximum size of a $v_{i}$-$V_{i}$ fan. We define the $\infty$-admissibility $\operatorname{adm}_{\infty}(G)$ of $G$ as
$\operatorname{adm}_{\infty}(G)\coloneqq\min_{\pi\in\Pi(G)}\max_{v\in V(G)}\operatorname{adm}_{\infty}[G,\pi,v].$
Again, we define a local version of admissibility.
Definition 2.19.
Let $G$ be a graph and let $r$ be a positive integer. Let $v\in V(G)$ and $A\subseteq V(G)$. A depth-$r$ $v$-$A$ fan is a set of paths $P_{1},\ldots P_{k}$ of length at most $r$ with one endpoint in $v$ and the other endpoint in $A$ and which are internally vertex disjoint from $A$ such that $V(P_{i})\cap V(P_{j})=\{v\}$ for all $i\neq j$.
Definition 2.20.
Let $G$ be a graph and let $r$ be a positive integer. Let $\pi=(v_{1},\ldots,v_{n})$ be a linear order of $V(G)$. For $1\leq i\leq n$, let $\operatorname{adm}_{r}[G,\pi,v_{i}]$ be the maximum size of a depth-$r$ $v_{i}$-$V_{i}$ fan. We define
$\operatorname{adm}_{r}(G,\pi)\coloneqq\max_{v\in V(G)}\operatorname{adm}_{r}[G,\pi,v],$
and the $r$-admissibility of $G$, $\operatorname{adm}_{r}(G)$, as
$\operatorname{adm}_{r}(G)\coloneqq\min_{\pi\in\Pi(G)}\max_{v\in V(G)}\operatorname{adm}_{r}[G,\pi,v].$
One last time, it is immediate from the definitions that
$\operatorname{col}(G)=\operatorname{adm}_{1}(G)\leq\operatorname{adm}_{2}(G)\leq\ldots\leq\operatorname{adm}_{n}=\operatorname{adm}_{\infty}(G).$
We will discuss the limit parameter $\operatorname{adm}_{\infty}(G)$ later, it turns out that graphs classes with finite $\operatorname{adm}_{\infty}(G)$ are those whose graphs are tree decomposable over torsos that have a bounded number of high degree vertices.
The three series $\operatorname{adm}_{r}(G),\operatorname{col}_{r}(G)$ and $\operatorname{wcol}_{r}(G)$ are generally referred to as the generalized coloring numbers. We have the following inequalities, which enable us to switch between them according to which is more suitable for a particular need. The inequalities of the first item are immediate from the definitions. For the proofs of the other items, we refer to [144].
Lemma 2.21.
Let $G$ be a graph, let $r$ be a positive integer. Then
  1. 1.
    $\operatorname{adm}_{r}(G)\leq\operatorname{col}_{r}(G)\leq\operatorname{wcol}_{r}(G)$.
  2. 2.
    $\operatorname{col}_{r}(G)\leq(\operatorname{adm}_{r}(G)-1)\cdot(\operatorname{adm}_{r}(G)-2)^{r-1}+1\leq\operatorname{adm}_{r}(G)^{r}$.
  3. 3.
    $\operatorname{wcol}_{r}(G)\leq\operatorname{col}_{r}(G)^{r}$.
  4. 4.
    $\operatorname{wcol}_{r}(G)\leq\operatorname{adm}_{r}(G)^{r}$.
We remark that the above relations involving the $r$-admissibility are stated differently in the work of Dvořák [50], due to the fact that the set of vertices weakly or strongly reachable from a vertex $v$ exclude the vertex $v$ in his work.
In our study of specific graph classes below, we will also see that these bounds are almost tight.
2.4 Logic and transductions
Let us recall some basics from first-order model theory. For extensive background we refer to Hodges’ textbook [82]. A signature is a collection of relation and function symbols, each with an associated arity. Let $\sigma$ be a signature. A $\sigma$-structure $\mathbb{A}$ consists of a non-empty set $A$, the universe of $\mathbb{A}$, together with an interpretation of each $k$-ary relation symbol $R\in\sigma$ as a $k$-ary relation $R^{\mathbb{A}}\subseteq A^{k}$ and an interpretation of each $k$-ary function symbol $f\in\sigma$ as a $k$-ary function $f^{\mathbb{A}}:A^{k}\rightarrow A$. In this work we assume that all structures are finite (i.e. have a finite universe and a finite signature).
We now define first-order logic FO and monadic second-order logic MSO (over the signature $\sigma$). We assume an infinite supply $\textsc{Var}_{1}$ of first-order variables (which will range over elements) and an infinite supply $\textsc{Var}_{2}$ of monadic second-order variables (which will range over sets of elements). Every variable is a term, and if $t_{1},\ldots,t_{k}$ are terms and $f\in\sigma$ is a $k$-ary function symbol, then also $f(t_{1},\ldots,t_{k})$ is a term. First-order formulas are built from the atomic formulas $t_{1}=t_{2}$, where $t_{1}$ and $t_{2}$ are terms, and $R(t_{1},\ldots,t_{k})$, where $R\in\sigma$ is a $k$-ary relation symbol and $t_{1},\ldots,t_{k}$ are terms, by closing under the Boolean connectives $\neg$¬$\wedge$ and $\vee$, and by existential and universal quantification $\exists x$ and $\forall x$. Monadic second-order formulas are defined as first-order formulas, but further allow the use of monadic quantifiers $\exists X$ and $\forall X$, and of a membership atomic formula $x\in X$, where $x$ is a first-order variable and $X$ a monadic second-order variable.
A variable $x$ not in the scope of a quantifier is a free variable (we do not consider formulas with free set second-order variables). A formula without free variables is a sentence. The quantifier rank $\mathrm{qr}(\varphi)$ of a formula $\varphi$ is the maximum nesting depth of quantifiers in $\varphi$. A formula without quantifiers is called quantifier-free.
If $\mathbb{A}$ is a $\sigma$-structure with universe $A$, then an assignment of the variables in $\mathbb{A}$ is a mapping $\bar{a}:\textsc{Var}\rightarrow A$. We use the standard notation $(\mathbb{A},\bar{a})\models\varphi(\bar{x})$ or $\mathbb{A}\models\varphi(\bar{a})$ to indicate that $\varphi$ is satisfied in $\mathbb{A}$ when the free variables $\bar{x}$ of $\varphi$ have been assigned by $\bar{a}$. We write $\models\varphi$ to express that $\varphi$ is a valid sentence, that is, $\varphi$ holds in every structure (of an appropriate signature). For a formula $\varphi(\bar{x})$ we define $\varphi(\mathbb{A}):=\{\bar{a}\in A^{|\bar{x}|}\mid\mathbb{A}\models\varphi(\bar{a})\}$.
In this work we consider the following types of structures:
  • 1.
    Colored graphs are $\sigma$-structures, where $\sigma$ consists of a single binary relation $E$ and unary relations, with the property that $E$ is symmetric and anti-reflexive.
  • 2.
    Guided pointer structures are $\sigma$-structures, where $\sigma$ consists of a single binary relation $E$, unary relations, and unary functions, with the property that $E$ is symmetric and anti-reflexive, and that every function $f\in\sigma$ is guided, meaning that if $f(u)=v$, then $u=v$ or $uv\in E(G)$.
  • 3.
    Ordered graphs are $\sigma$-structures, where $\sigma$ consists of a two binary relations $E$ and $<$, with the property that $E$ is symmetric and anti-reflexive, and $<$ is a linear order.
  • 4.
[…]
Let $\sigma,\tau$ be relational signatures and let $\mathcal{L}$ be one of $\mathrm{FO}$ or $\mathrm{MSO}$. An $\mathcal{L}$-interpretation $\mathsf{I}$ of $\sigma$-structures in $\tau$-structures is a tuple $\mathsf{I}=(\nu(\bar{x}),(\rho_{R}(\bar{x}_{1},\ldots,\bar{x}_{ar(R)}))_{R\in\sigma})$, where $\nu(\bar{x})$ and $\rho_{R}(\bar{x}_{1},\ldots,\bar{x}_{ar(R)})$ are $\mathcal{L}$-formulas, and $|\bar{x}|=|\bar{x}_{i}|$ for all $1\leq i\leq ar(R)$ for all $R\in\sigma$, where $ar(R)$ is the arity of $R$. For every $\tau$-structure $\mathbb{B}$, the $\sigma$-structure $\mathbb{A}=\mathsf{I}(\mathbb{B})$ has the universe $\nu(\mathbb{B})$ and each relation $R_{\mathbb{A}}$ is interpreted as $\rho_{R}(\mathbb{B})$. We say that $\nu$ defines the universe and $\rho_{R}$ defines the relation $R$ of $\mathsf{I}(\mathbb{B})$. The number of free variables $|\bar{x}|$ of $\nu$ is the dimension of the interpretation. A $1$-dimensional interpretation is called a simple interpretation. In particular, a simple interpretation $\mathsf{I}$ of graphs in $\tau$-structures is a pair $\mathsf{I}=(\nu,\eta)$, where $\nu(x)$ and $\eta(x,y)$ are formulas (such that $\models\forall x\forall y(\eta(x,y)\rightarrow\eta(y,x)\wedge\neg\eta(x,x))$). For a $\tau$-structure $\mathbb{B}$ with universe $B$, the graph $\mathsf{I}(\mathbb{B})$ has vertex set $\nu(\mathbb{B})=\{u\in B\mid\mathbb{A}\models\nu(u)\}$ and edge set $\eta(\mathbb{A})=\{uv\mid\mathbb{A}\models\eta(u,v)\}$.
A monadic lift of a $\tau$-structure $\mathbb{B}$ is a $\tau^{+}$-expansion $\mathbb{B}^{+}$ of $\mathbb{B}$, where $\tau^{+}$ is the union of $\tau$ and a set $\Sigma$ of unary relation symbols (also called colors). We also say that $\mathbb{B}^{+}$ is a $\Sigma$-colored $\tau$-structure.
A transduction $\mathsf{T}$ of $\sigma$-structures from $\tau$-structures is a pair $\mathsf{T}=(\Sigma,\mathsf{I})$, where $\Sigma$ is a finite set of colors and $\mathsf{I}$ is a simple interpretation of $\sigma$-structures in $\Sigma$-colored $\tau$-structures.
A $\sigma$-structure $\mathbb{A}$ can be $\mathsf{T}$-transduced from a $\tau$-structure $\mathbb{B}$ if there exists a $\Sigma$-coloring $\mathbb{B}^{+}$ of $\mathbb{B}$ such that $\mathbb{A}=\mathsf{I}(\mathbb{B}^{+})$. A class $\mathscr{C}$ of $\sigma$-structures can be $\mathsf{T}$-transduced from a class $\mathscr{D}$ of $\tau$-structures if for every structure $\mathbb{A}\in\mathscr{C}$ there exists a structure $\mathbb{B}\in\mathscr{D}$ such that $\mathbb{A}$ can be $\mathsf{T}$-transduced from $\mathbb{B}$. A class $\mathscr{C}$ of $\sigma$-structures can be transduced from a class $\mathscr{D}$ of $\tau$-structures if it can be $\mathsf{T}$-transduced from $\mathscr{D}$ for some transduction $\mathsf{T}$. An FO/MSO-transduction is a transduction whose interpretation uses FO/MSO formulas.
For a relational signature $\sigma$, the incidence graph of a $\sigma$-structure $\mathbb{A}$ with universe $A$ is the colored graph with vertices $A$ and one vertex for each tuple $\bar{v}$ appearing in a relation $R_{\mathbb{A}}$ (we take multiple copies if a tuple appears in several relations). If $\bar{v}=(v_{1},\ldots,v_{k})$ and $\bar{v}\in R_{\mathbb{A}}$, then the vertex $\bar{v}$ in the incidence graph is marked with a color $R$ and connected with an edge of color $i$ with the vertex $v_{i}$ for $1\leq i\leq k$. Note that $\mathbb{A}$ is interpretable by a simple interpretation from its incidence graph, but vice versa, in general we need a higher-dimensional interpretation to interpret the incidence graph of a structure from the structure.
We define the structural complexity of a class of structures as the structural complexity of the class of its incidence graphs. In particular, we can define the above defined concepts of treedepth and treewidth via transductions as follows. We write $\mathscr{T}_{d}$ for the class of trees of depth at most $d$ and $\mathscr{T}$ for the class of all trees.
  • 1.
    A class $\mathscr{C}$ of graphs has bounded treedepth if and only if the class of incidence graphs of graphs from $\mathscr{C}$ can be FO- or MSO-transduced from $\mathscr{T}_{d}$ for some $d\geq 1$ (follows from the work of Ganian et al. [72, 73]).
  • 2.
    A class $\mathscr{C}$ of graphs has bounded treewidth if and only if the class of incidence graphs of graphs from $\mathscr{C}$ can be MSO-transduced from $\mathscr{T}$ by a classical result of Courcelle [27].
  • 3.
    A class $\mathscr{C}$ of graphs has bounded treewidth if and only if the class of incidence graphs of graphs from $\mathscr{C}$ can be FO-transduced from the class of all (finite) tree orders as shown by Colcombet [24].
Transductions allow to define dense but structurally sparse classes from sparse classes of graphs (or structures). For example, we can define the dense analogs of classes with bounded treedepth and classes with bounded treewidth via transductions.
  • 1.
    A class $\mathscr{C}$ of graphs has bounded shrubdepth if and only if $\mathscr{C}$ can be FO- or MSO-transduced from $\mathscr{T}_{d}$ for some $d\geq 1$ as shown by Ganian et al. [72, 73].
  • 2.
    A class $\mathscr{C}$ of graphs has bounded cliquewidth if and only if $\mathscr{C}$ can be MSO-transduced from $\mathscr{T}$ as proved by Courcelle [27].
  • 3.
    A class $\mathscr{C}$ of graphs has bounded cliquewidth if and only if $\mathscr{C}$ can be FO-transduced from the class of all (finite) tree orders as shown by Colcombet [24].
We will revisit the concept of structural sparsity in later sections. We conclude this section by stating the following basic quantifier-elimination result, which is a key ingredient for the efficient first-order model checking algorithm on classes with bounded expansion by Dvořák, Král’ and Thomas [53].
Theorem 2.22 ([53]).
For every FO-formula $\varphi(\bar{x})$ and every class $\mathscr{C}$ of colored graphs with bounded treedepth there exists a quantifier-free formula $\tilde{\varphi}(\bar{x})$ and a linear-time computable map $Y$ such that, for every $G\in\mathscr{C}$, $Y(G)$ is a guided expansion of $G$ such that for all tuples of vertices $\bar{v}$
$G\models\varphi(\bar{v})\quad\iff\quad Y(G)\models\tilde{\varphi}(\bar{v}).$
3 General bounds for the generalized coloring numbers
In this section, we present lower and upper bounds for the generalized coloring numbers. It turns out that the numbers are intimately linked to the density of bounded-depth minors and topological minors. These are the central notions in the theory of bounded expansion and nowhere dense graph classes, which were introduced by Nešetřil and Ossona de Mendez [114, 116]. After introducing these concepts, we present simple lower bounds for the generalized coloring numbers in terms of the density of depth-$r$ minors, and depth-$r$ topological minors. Throughout this work when $r$ is not explictly quantified it denotes a non-negative integer.
We then present a characterization of graphs with bounded $r$-admissibility in terms of forbidden substructures, which is due to Dvořák [50]. We derive the fixed-parameter linear time algorithm of Dvořák [50] to compute $\operatorname{adm}_{r}(G)$. Based on the forbidden substructure characterization we also derive an upper bound on $\operatorname{adm}_{r}(G)$ in terms of the density of depth-$r$ topological minors, which is due to Grohe et al. [76] and prove the existence of universal orders, that is, orders $\pi$ that witness $\operatorname{adm}_{r}(G)\leq f(r)\cdot\operatorname{adm}_{r}(G,\pi)$ for all values of $r$ and some function $f$. The existence of such orders was first proved by Van den Heuvel and Kierstead [148].
We then discuss the approach of Grohe et al. [77] to approximate the weak $r$-coloring numbers via transitive fraternal augmentations, introduced by Nešetřil and Ossona de Mendez [122].
3.1 Graph expansion and simple lower bounds
The following notions of bounded depth minors and bounded depth topological minors are the fundamental definitions in the theory of bounded expansion and nowhere dense graph classes introduced by Nešetřil and Ossona de Mendez [114, 116].
Definition 3.1.
[…]
  1. 1.
    if $u,v\in V(H)$ with $u\not=v$, then $M(v)$ and $M(u)$ are vertex disjoint, and
  2. 2.
    if $e=uv\in E(H)$, then $M(e)=u^{\prime}v^{\prime}\in E(G)$ for vertices $u^{\prime}$ of $M(u)$ and $v^{\prime}$ of $M(v)$.
Figure 9: Graph $H$ (on the right) is a depth-$2$ minor of $G$ (on the left).
The set $M(v)$ for a vertex $v\in V(H)$ is called the branch set or model of $v$ in $G$. The map $M$ is called the model of $H$ in $G$.
[…]
Definition 3.2.
[…]
  1. 1.
    if $u,v\in V(H)$ with $u\not=v$, then $T(v)\neq T(u)$,
  2. 2.
    if $e=uv\in E(H)$, then $T(e)$ is a path with endpoints $T(u)$ and $T(v)$, and
  3. 3.
    if $e,e^{\prime}\in E(H)$ with $e\neq e^{\prime}$, then $T(e)$ and $T(e^{\prime})$ are internally vertex disjoint.
[…]
Figure 10: Graph $H$ (on the right) is a depth22-2- 2 topological minor of $G$ (on the left).
[…]
Definition 3.3.
Let $G$ be a graph. The edge density of $G$ is
$\nabla(G)=|E(G)|/|V(G)|.$
The greatest reduced average density $\nabla_{r}(G)$ of $G$ with rank $r$ is
[…]
and the greatest topological reduced average density ~$\widetilde{\nabla}_{r}(G)$ of $G$ with rank $r$ is
[…]
Similarly, we define
[…]
as the largest clique that appears as a depth-$r$ minor, and
[…]
as the largest clique that appears as a topological depth-$r$ minor.
Note that $\nabla_{0}(G)$ is equal to the maximum edge density over all subgraphs of $G$. In particular, the degeneracy of $G$ is equal to $2\nabla_{0}(G)$ and we derive the following lemma for the chromatic number $\chi$ of $G$.
Lemma 3.4.
[…]
We refer to the functions $r\mapsto\nabla_{r}(G)$ and $r\mapsto\widetilde{\nabla}_{r}(G)$ as the expansion and topological expansion of $G$, respectively. As proved by Dvořák [48], these measures satisfy
~$\widetilde{\nabla}_{r}(G)\leq\nabla_{r}(G)\leq 4\big{(}4\widetilde{\nabla}_{r}(G)\big{)}^{(r+1)^{2}}.$
Hence, expansion and topological expansion are functionally equivalent. The key observation for the proof is that in the branch set of a high-degree vertex in a bounded depth minor model, we will always find a vertex of large degree. We can use such a vertex as the branching vertex of a high-degree vertex in a topological minor model.
The functions $r\mapsto\omega_{r}(G)$ and $r\mapsto\widetilde{\omega}_{r}(G)$ are called the $\omega$-expansion and topological $\omega$-expansion of $G$, respectively. As shown by Nešetřil and Ossona de Mendez [116] also the $\omega$-expansion and topological $\omega$-expansion are functionally related, see [129] for the following bounds.
$\widetilde{\omega}_{r}(G)\leq\omega_{r}(G)\leq 1+\widetilde{\omega}_{3r+1}(G)^{2r+2}.$
The above notions extend to classes $\mathscr{C}$ of graphs by defining $\nabla_{r}(\mathscr{C})=\sup_{G\in\mathscr{C}}\nabla_{r}(G)$, ~$\widetilde{\nabla}_{r}(\mathscr{C})=\sup_{G\in\mathscr{C}}\widetilde{\nabla}_{r}(G)$, $\omega(\mathscr{C})=\sup_{G\in\mathscr{C}}\omega_{r}(G)$, and $\widetilde{\omega}_{r}(\mathscr{C})=\sup_{G\in\mathscr{C}}\widetilde{\omega}_{r}(G)$.
We can now define bounded expansion and nowhere dense graph classes.
Definition 3.5.
A class $\mathscr{C}$ of graphs has bounded expansion if there exists a function $d$ such that for every non-negative integer $r$ we have $\nabla_{r}(\mathscr{C})\leq d(r)$.
The class $\mathscr{C}$ has polynomial expansion if the function $d$ is polynomially bounded.
Definition 3.6.
A class $\mathscr{C}$ of graphs is nowhere dense if there exists a function $t$ such that for every non-negative integer $r$ we have $\omega_{r}(\mathscr{C})\leq t(r)$.
Classes of bounded expansion are classes of sparse graphs by definition. It turns out that also nowhere dense classes are sparse.
Theorem 3.7 ([116]).
A class $\mathscr{C}$ of graphs is nowhere dense if and only if for all non-negative integers $r$ and all real $\varepsilon>0$, all $n$-vertex graphs $G\in\mathscr{C}$ satisfy $\nabla_{r}(G)\in\mathcal{O}_{r,\varepsilon}(n^{\varepsilon})$.
Here, the notation $\mathcal{O}_{r,\varepsilon}$ hides constant factors depending only on $r$ and $\varepsilon$. By rescaling $\varepsilon$, we can equivalently say that a class $\mathscr{C}$ is nowhere dense if and only if for all real $\varepsilon>0$ and all positive integers $r$ all sufficiently large $n$-vertex graphs $G\in\mathscr{C}$ satisfy $\nabla_{r}(G)\leq n^{\varepsilon}$.
Many familiar classes of sparse graphs have bounded expansion or are nowhere dense.
Example 3.8.
The following classes of graphs have bounded expansion.
  • 1.
    Every class $\mathscr{C}$ that excludes a fixed graph $H$ as a minor. For such classes, there exists an absolute constant $c$ such that for all $G\in\mathscr{C}$ and all positive integers $r$ we have $\nabla_{r}(G)\leq c$. In particular, every class that excludes a fixed minor has polynomial expansion. Special cases are the class of planar graphs, every class of graphs that can be drawn with a bounded number of crossings, see [120], and every class of graphs that embeds into a fixed surface.
  • 2.
    Every class of intersection graphs of low-density objects in low dimensional Euclidean space has polynomial expansion [80].
  • 3.
    Every class $\mathscr{C}$ that excludes a fixed graph $H$ as a topological minor. For such classes, there exists an absolute constant $c$ such that for all $G\in\mathscr{C}$ and all positive integers $r$ we have ~$\widetilde{\nabla}_{r}(G)\leq c$. Every class that excludes $H$ as a minor also excludes $H$ as a topological minor. Further special cases are classes of bounded degree and classes of graphs that can be drawn with a linear number of crossings, see [120].
  • 4.
    Every class of graphs that can be drawn with a bounded number of crossings per edge [120].
  • 5.
    Every class of graphs with bounded queue-number, bounded stack-number or bounded non-repetitive chromatic number [120].
  • 6.
    The class of Erdös-Rényi random graphs with constant average degree $d/n$, $G(n,d/n)$, has asymptotically almost surely bounded expansion [120].
Example 3.9.
The class of graphs $G$ that satisfy $\mathrm{girth}(G)\geq\Delta(G)$ is nowhere dense and does not have bounded expansion [111].
We will next establish a connection between the generalized coloring numbers and the expansion and topological expansion of a graph, proving that classes of bounded expansion are exactly those classes that have bounded generalized coloring numbers for every value of $r$. We first observe the following simple lower bounds, which are variations of Lemma 3.3 of [158] proved by Zhu.
Lemma 3.10 ([158]).
Let $G$ be a graph and let $r$ be a positive integer. Then $\nabla_{r}(G)\leq\operatorname{wcol}_{2r+1}(G)$.
Proof.
Let $d\coloneqq\operatorname{wcol}_{2r+1}(G)$. We show that for every depth-$r$ minor $H$ of $G$ we can find an orientation $\vec{H}$ such that the out-degree of every vertex $v\in V(\vec{H})$ is at most $d$. By adding the out-degrees of all vertices in $\vec{H}$ we count the edges of $H$, and thereby show that the ratio $|E(H)|/|V(H)|$ is bounded by $d$.
Let $H$ be a depth-$r$ minor of $G$ and let $M$ be a depth-$r$ minor model of $H$ in $G$. Further, let $\pi$ be an order of $V(G)$ with $\operatorname{wcol}_{2r+1}(G,\pi)\leq d$, that is, $|\operatorname{WReach}_{2r+1}[G,\pi,v]|\leq d$ for each $v\in V(G)$. For every vertex $u\in V(H)$, let $c(u)$ be a central vertex of $M(u)$, that is, a vertex such that the distance in $M(u)$ between $c(u)$ and every other vertex $v$ of $M(u)$ is at most $r$. For $uv\in E(H)$, let $P(u,v)$ be a path of length at most $12r+$ 1 between $c(u)$ and $c(v)$ that uses only vertices of $M(u)$ and $M(v)$, more precisely, such that $P(u,v)=(c(u)=v_{0},\ldots,v_{\ell}=c(v))$, $11\leq\ell\leq 2r+$ 1, and there is an index $i$ such that $v_{0},\ldots,v_{i}$ are vertices of $M(u)$ and $v_{i+1},\ldots,v_{\ell}$ are vertices of $M(v)$.
To obtain the orientation $\vec{H}$, we orient an edge $uv\in E(H)$ towards $v$ if the smallest vertex $m(u,v)$ (with respect to $\pi$) of $P(u,v)$ lies in $M(v)$, and towards $u$ otherwise. Observe that if $uv$ is oriented towards $v$, then $m(u,v)\in\operatorname{WReach}_{2r+1}[G,\pi,c(u)]$. Furthermore, observe that if $uv$ and $uw$ are different edges that are oriented away from $u$, then $m(u,v)$ and $m(u,w)$ are different vertices (from $M(v)$ and $M(w)$, respectively, which are disjoint by definition). This implies that the out-degree of $u$ in $\vec{H}$ is bounded by $|\operatorname{WReach}_{2r+1}[G,\pi,c(u)]|\leq d$. Since we considered an arbitrary vertex $u$ of $H$, we have found the desired orientation $\vec{H}$. ∎
Lemma 3.11.
Let $G$ be a graph and let $r$ be a positive integer. Then $\nabla_{r}(G)\leq\operatorname{col}_{4r+1}(G)$.
Proof.
Let $d\coloneqq\operatorname{col}_{4r+1}(G)$, let $H$ be a depth-$r$ minor of $G$ and let $M$ be a depth-$r$ minor model of $H$ in $G$. As in the previous proof, we show that we can find an orientation $\vec{H}$ of $H$ such that the out-degree of every vertex $v\in V(\vec{H})$ is at most $d$. Let $\pi$ be an order of $V(G)$ with $\operatorname{col}_{4r+1}(G,\pi)\leq d$, that is, $|\operatorname{SReach}_{4r+1}[G,\pi,v]|\leq d$ for every $v\in V(G)$.
For every vertex $u\in V(H)$, let $m(u)$ be the smallest, with respect to $\pi$, vertex of the branch set $M(u)$. To obtain the orientation $\vec{H}$, we orient an edge $uv\in E(H)$ towards $v$ if $m(v)<_{\pi}m(u)$, and towards $u$ otherwise.
For every edge $uv\in E(H)$ there exists a path $P(u,v)$ of length at most $14r+$ 1 connecting $m(u)$ and $m(v)$ and which uses only vertices of $M(u)$ and $M(v)$, more precisely, such that $P(u,v)=(c(u)=v_{0},\ldots,v_{\ell}=c(v))$, $11\leq\ell\leq 4r+$ 1, and there is an index $i$ such that $v_{0},\ldots,v_{i}$ are vertices of $M(u)$ and $v_{i+1},\ldots,v_{\ell}$ are vertices of $M(v)$. Observe that if $uv$ is oriented towards $v$, then the first vertex $w(u,v)$ of $P(u,v)$ that lies in $M(v)$ and satisfies $w<_{\pi}c(u)$ (such a vertex must exist, as $c(v)$ satisfies these properties) satisfies $w\in\operatorname{SReach}_{4r+1}[G,\pi,c(u)]$. Observe that if $uv$ and $uy$ are different edges that are oriented away from $u$, then the vertices $w(u,v)$ and $w(u,y)$ are different. This implies that the out-degree of $u$ in $\vec{H}$ is bounded by $\operatorname{SReach}_{4r+1}[G,\pi,c(u)]$. Since $u$ is an arbitrary vertex of $H$, we have found the desired orientation $\vec{H}$. ∎
Finally, the following lemma is proved exactly as Lemma 3.10, by deriving a depth-$r$ minor model $M$ from a topological minor model and using that all branch sets of $M$ are subdivided stars.
Lemma 3.12.
Let $G$ be a graph and let $r$ be a positive integer. Then ~$\widetilde{\nabla}_{r}(G)\leq\operatorname{adm}_{2r+1}(G)$.
3.2 Bounding r-admissibility
We now turn our attention to upper bounds for the generalized coloring numbers. Recall Definition 2.19 of an $a$-$A$ fan.
Definition 3.13.
Let $G$ be a graph and let $r$ and $k$ be positive integers. A $(k,r)$-fan set in $G$ is a set $A\subseteq V(G)$ such that for every $a\in A$, there exists a depth-$r$ $a$-$A$ fan of order $k$ in $G$.
Figure 11: A $(3,3)$-fan set.
The following theorem is implicit in the work of Dvořák [50].
Theorem 3.14 ([50]).
Let $G$ be a graph and let $r$ be a positive integer. Then $\operatorname{adm}_{r}(G)$ is equal to the largest number $k$ such that there exists a $(k,r)$-fan set in $G$.
Proof.
If $A$ is a $(k,r)$-fan set in $G$, then $\operatorname{adm}_{r}(G)\geq k$. To see this, let $\pi$ be any order of $V(G)$ and let $a$ be the maximal element of $A$ with respect to $\pi$. As $A$ is a $(k,r)$-fan set, there exist paths $P_{1},\ldots,P_{k-1}$ of length at most $r$ with one endpoint in $a$ and the other endpoint in $A$ that are internally disjoint from $A$ and such that $V(P_{i})\cap V(P_{j})=\{a\}$ for all $i\neq j$. For every path $P_{i}$, there exists an element $b_{i}\in V(P_{i})$ that is smaller than or equal to $a$ with respect to $\pi$ by our choice of the vertex $a$. Hence, the vertex $a$ witnesses that $\operatorname{adm}_{r}(G,\pi)\geq k$. As $\pi$ was arbitrary, we conclude that $\operatorname{adm}_{r}(G)\geq k$.
Vice versa, assume that $k$ is the largest integer such that there exists a $(k,r)$-fan set in $G$. We consider the following iterative procedure to construct an order $\pi=(v_{1},\ldots,v_{n})$ of $V(G)$ with $\operatorname{adm}_{r}(G,\pi)\leq k$. For a set $S\subseteq V(G)$ and a vertex $v\in S$, denote by $b_{r}(S,v)$ the maximum size of a depth-$r$ $v-S$ fan.
We initialize $S_{n}\coloneqq V(G)$. Now, for $i=n,\ldots,1$, if $v_{i+1},\ldots,v_{n}$ have already been ordered, choose $v_{i}$ such that if $S_{i}=\{v_{1},\ldots,v_{i}\}$, then $p_{i}\coloneqq b_{r}(S_{i},v_{i})$ is minimal. Clearly, the $r$-admissibility of the resulting order is the maximum of the values $p_{i}$ occurring in its construction. Furthermore, observe that the set $S_{i}$ is a $(p_{i},r)$-fan set. Since $k$ is the largest number such that there exists a $(k,r)$-fan set in $G$, we conclude that $\operatorname{adm}_{r}(G,\pi)\leq k$. ∎
The iterative procedure described in the proof of Theorem 3.14 can be turned into an algorithm to compute the $r$-admissibility in a straightforward manner. A problematic step is to compute the values $b_{r}(S_{i},v_{i})$, since in general graphs it is NP-complete to determine the value $b_{r}(S,v)$ for a vertex $v$ and set $S$, for values of $r$ greater than $4$, as shown by Alon et al. [85]. However, Dvořák presents a simple approach to approximate the values $b_{r}(S,v)$, and hence the $r$-admissibility, up to a factor of $r$. This approach can be easily implemented and runs in time $\mathcal{O}(mn^{3})$. Using a sophisticated data structure presented by Dvořák, Král’ and Thomas [53], Dvořák showed how to compute the exact values in fixed-parameter linear time on every class of bounded expansion [50].
Theorem 3.15 ([50]).
Let $\mathscr{C}$ be a class of bounded expansion and let $r$ be a positive integer. There exists a function $f$ and an algorithm that given an $n$-vertex graph $G\in\mathscr{C}$ computes an order $\pi$ with $\operatorname{adm}_{r}(G,\pi)=\operatorname{adm}_{r}(G)$ in time $f(r)\cdot n$.
In view of Lemma 2.21, the algorithm of Theorem 3.15 also gives linear time computable approximations for the strong $r$-coloring numbers and weak $r$-coloring numbers. It was shown by Grohe et al. [76] that for $r\geq 3$, computing exact values for $\operatorname{wcol}_{r}(G)$ is NP-complete. Similarly, it is shown by Dvořák [48] that deciding whether a graph contains a depth-$1$ minor of density at least $d$ is NP-complete for $d\geq 2$ and by Muzi et al. [107] that deciding whether a graph contains a $1$-subdivision of density at least $d$ for $d\geq 2$ is also NP-complete.
Problem 3.16.
Let $r\geq 2$ be a positive integer. Does there exist an algorithm that, given an $n$-vertex graph $G$, computes $\operatorname{col}_{r}(G)$ and $\operatorname{wcol}_{r}(G)$ in time $n^{f(r,\operatorname{col}_{r}(G))}$, or $n^{f(r,\operatorname{wcol}_{r}(G))}$, for some function $f$?
How can it possibly be hard to compute the values $b_{r}(S_{i},v_{i})$? According to Menger’s theorem the number of vertex disjoint paths between two vertices $s$ and $t$ is equal to the minimum cardinality of an $s$-$t$-cut, that is, of a set of vertices in $V(G)\setminus\{s,t\}$ hitting all paths between $s$ and $t$ (and such cuts can be efficiently computed). However, this fact is not true for paths of bounded length as observed by Adámek and Koubek [2]: a hitting set for all $s$-$t$-paths of length at most $r$ can be larger by factor $r$ than the number of disjoint $s$-$t$ paths of length at most $r$.
We now turn our attention to upper bounds for the generalized coloring numbers. The connection between the generalized coloring numbers and the expansion of a graph was first made by Zhu [158].
Theorem 3.17 (Lemma 3.4 of [158]).
Let $G$ be a graph and let $r$ be a positive integer. Then
$\operatorname{col}_{r}(G)\leq 1+q_{r},$
where $q_{1}=2\nabla_{\left\lceil(r-1)/2\right\rceil}(G)$, and for $i\geq 1$, $q_{i+1}$ is inductively defined as $q_{i+1}=q_{1}\cdot q_{i}^{2i^{2}}$.
The currently best known bounds for $\operatorname{adm}_{r}(G),\operatorname{col}_{r}(G)$ and $\operatorname{wcol}_{r}(G)$ follow from the following theorem presented by Grohe et al. [76] combined with Lemma 2.21.
Theorem 3.18 ([76]).
Let $G$ be a graph and let $r$ be a positive integer. Then $G$ does not contain a $(k,r)$-fan set for $k>6r(\widetilde{\nabla}_{r-1}(G))^{3}$. Consequently,
$\operatorname{adm}_{r}(G)\leq 6r\cdot\widetilde{\nabla}_{r-1}(G)^{3},$
and
$\operatorname{col}_{r}(G)\leq\operatorname{wcol}_{r}(G)\leq(6r)^{r}\cdot\widetilde{\nabla}_{r-1}(G)^{3r}.$
Proof.
Let $G$ be a graph with ~$\widetilde{\nabla}_{r-1}(G)\leq c$, and let $k\coloneqq 6rc^{3}+1$. Assume towards a contradiction that $\operatorname{adm}_{r}(G)>k$. By Theorem 3.14 there exists a $(k+1,r)$-fan set $A$. Let $s\coloneqq|A|$. For $v\in S$ let $\mathcal{P}_{v}$ be a set of at least $k$ paths of length at most $r$ between $v$ and $S\setminus\{v\}$ such that $V(P)\cap V(Q)=\{v\}$ for $P\neq Q\in\mathcal{P}_{v}$ (we start with at least $k+1$ paths by definition of a $(k+1,r)$-fan and by dropping the path of length 00 we are left with at least $k$ paths with one endpoint in $A\setminus\{v\}$).
[…]
By Lemma 3.4 $H$ is $(2c+1)$-colorable and hence contains an independent set $R$ of size at least $\left\lceil s/(2c+1)\right\rceil$. For $v\in R$, let $\mathcal{Q}_{v}$ be the set of initial segments of paths in $\mathcal{P}_{v}$ between $v$ and a vertex in $(M\cup A)\setminus\{v\}$ with all internal vertices in $V(G)\setminus(M\cup A)$. Observe that for $u,v\in R$ the paths in $\mathcal{Q}_{v}$ and $\mathcal{Q}_{u}$ are internally vertex disjoint: if $P\in\mathcal{Q}_{u}$ and $Q\in\mathcal{Q}_{v}$ had an internal vertex in common, then $P\cup Q$ would contain a path of length at most $22r-$ 2 that is internally disjoint from all paths in $\mathcal{P}$, contradicting the maximality of $\mathcal{P}$.
Let $G^{\prime}$ be the subgraph of $G$ induced by the vertices of the paths in $\mathcal{P}$ and the paths in $\mathcal{Q}_{v}$ for $v\in R$. Let $H^{\prime}$ be the graph obtained from $G^{\prime}$ by contracting for each $v\in R$ all paths in $\mathcal{Q}_{v}$ to single edges. Then $H^{\prime}\preceq^{t}_{r-1}G$. We have $|V(H^{\prime})|\leq s+m\leq s+s\cdot c\cdot(2r-2)\leq s\cdot c\cdot(2r-1)$ and at least $|E(H^{\prime})|\geq\left\lceil s/(2c+1)\right\rceil\cdot k$ edges. Thus, the edge density of $H^{\prime}$ is
$\frac{|E(H^{\prime})|}{|V(H^{\prime})|}\geq\frac{\left\lceil s/(2c+1)\right\rceil\cdot k}{s\cdot c\cdot(2r-1)}\geq\frac{s\cdot k}{s\cdot c\cdot(2c+1)\cdot(2r-1)}\geq\frac{6rc^{3}+1}{(2c^{2}+c)\cdot 2r}\geq\frac{6rc^{3}+1}{3c^{2}\cdot 2r}>c.$
A contradiction, showing that $k\leq 6rc^{3}$. ∎
Corollary 3.19.
A class $\mathscr{C}$ of graphs has bounded expansion if and only if there exist functions $a,c$ and $w$ such that for every positive integer $r$ and every $G\in\mathscr{C}$ we have $\operatorname{adm}_{r}(G)\leq a(r)$, $\operatorname{col}_{r}(G)\leq c(r)$ and $\operatorname{wcol}_{r}(G)\leq w(r)$.
Observe that the function bounding the $r$-admissibility in Theorem 3.14 is polynomial in $r$ and ~$\widetilde{\nabla}_{r}(G)$. One may wonder if a similar relation holds between $\operatorname{col}_{r}(G)$ and $\nabla_{r}(G)$. The following problem is attributed to Joret and Wood in [60].
Problem 3.20.
Does there exist a polynomial $p(x,y)$ such that for all graphs $G$ and all positive integers $r$ we have $\operatorname{col}_{r}(G)\leq p(r,\nabla_{r}(G))$?
The problem appears already in bounded degree graphs.
Problem 3.21.
Does there exist a polynomial $p(x,y,z)$ such that for all graphs $G$ with maximum degree $\Delta$ and all positive integers $r$ we have $\operatorname{col}_{r}(G)\leq p(r,\nabla_{r}(G),\Delta)$?
3.3 Universal orders
According to Corollary 3.19, if $\mathscr{C}$ has bounded expansion, then there exists a function $f$ such that for every value of $r$ and every $G\in\mathscr{C}$ there exists an order $\pi$ of $V(G)$ such that $\operatorname{adm}_{r}(G,\pi)\leq f(r)$. A universal order is an order that works for every value of $r$ at the same time.
Definition 3.22.
Let $G$ be a graph and let $f$ be a function. An order $\pi$ of $V(G)$ is a universal order bounded by $f$ if $\operatorname{adm}_{r}(G,\pi)\leq f(r)$ for every non-negative integer $r$.
By Lemma 2.21 we could equivalently ask that $\operatorname{col}_{r}(G,\pi)\leq f(r)$ or $\operatorname{wcol}_{r}(G,\pi)\leq f(r)$ in the above definition. It was shown by Van den Heuvel and Kierstead [148] that all classes of bounded expansion admit universal orders. In this section, we provide an alternative proof.
For this, we consider a variant of the iterative procedure presented in the proof of Theorem 3.14 to construct the order $\pi=(v_{1},\ldots,v_{n})$. In the procedure of Theorem 3.14, Theorem 3.18 guarantees that in step $i$ we will always find a vertex $v_{i}$ with back-connectivity $b_{r}(S_{i},v_{i})\leq 6r(\widetilde{\nabla}_{r}(G))^{3}$. We are going to prove that we do not only find one vertex with small back-connectivity, but in fact, a large fraction of the vertices have a back-connectivity that is only slightly larger than the optimal back-connectivity. For every value of $r$, there will hence be only a few vertices violating the $r$-admissibility constraint, so that we are still left with at least one vertex that can be inserted as $v_{i}$ without violating the constraint for any value of $r$. The proof of the following lemma is a variation of the proof of Theorem 3.18 (only the numbers need to be adapted), which is presented by Pilipczuk et al. [131] (the arXiv version of [132]).
Lemma 3.23 ([131], Lemma 14).
Let $G$ be a graph and $S\subseteq V(G)$. Let $\varepsilon>0$ and let $r$ be a positive integer. Then at most $\varepsilon\cdot|S|$ vertices $v\in S$ have $b_{r}(S,v)>6\cdot\varepsilon^{-1}\cdot r\cdot\bigl{(}\widetilde{\nabla}_{r-1}(G)\bigr{)}^{3}$.
We are ready to prove the existence of universal orders.
Theorem 3.24.
Let $G$ be a graph and let $f(r)\coloneqq 2^{r+1}\cdot 6r\cdot\bigr{(}\widetilde{\nabla}_{r-1}(G)\bigl{)}^{3}$. Then $G$ admits an order $\pi$ with $\operatorname{adm}_{r}(G,\pi)\leq f(r)$ for every positive integer $r$.
Proof.
We adapt the iterative procedure presented in the proof of Theorem 3.14 to an order $\pi=(v_{1},\ldots,v_{n})$ of $V(G)$ with $\operatorname{adm}_{r}(G,\pi)\leq f(r)$ for all positive integers $r$.
We initialize $S_{n}\coloneqq V(G)$. Now, for $i=n,\ldots,1$, if $v_{i+1},\ldots,v_{n}$ have already been ordered, choose $v_{i}$ such that if $S_{i}=\{v_{1},\ldots,v_{i}\}$, then $b_{r}(S_{i},v_{i})\leq f(r)$ for all positive integers $r$. Clearly, if such a choice for $v_{i}$ is possible, which remains to be shown, then $\pi$ has the desired properties.
According to Lemma 3.23, for any set $S$ and any positive integer $r$, at most $1/(2^{r+1})\cdot|S|$ vertices have back-connectivity $b_{r}(S,v)>f(r)$. Hence, at most $|S|/2$ vertices conflict $f$ for the value $r=0$, at most $|S|/4$ vertices conflict $f$ for the value $r=1$, and so on. Hence, at most
$\sum_{r=0}^{n}\frac{1}{2^{r+1}}\cdot|S|<|S|,$
vertices of $S$ are conflicting $f$ for some value of $r$. We conclude that among all vertices of $S$, we find at least one vertex $v_{i}$ satisfying $b_{r}(S_{i},v_{i})\leq f(r)$ for all values of $r$. ∎
Corollary 3.25.
Let $\mathscr{C}$ be a class of bounded expansion. Then there exists a function $f$ such that for all $G\in\mathscr{C}$ there exists an order $\pi$ with $\operatorname{adm}_{r}(G,\pi)\leq f(r)$ for every positive integer $r$.
We can use the approach of Dvořák to approximate the values $b_{r}(S,v)$ up to a factor of $r$ and thereby, for every fixed class $\mathscr{C}$ of bounded expansion obtain an algorithm that computes a universal order in polynomial time.
Corollary 3.26.
Let $G$ be a graph and let $f(r)\coloneqq 2^{r+1}\cdot 6r^{2}\cdot\bigr{(}\widetilde{\nabla}_{r-1}(G)\bigl{)}^{3}$. Then we can compute in polynomial time an order $\pi$ of $G$ with $\operatorname{adm}_{r}(G,\pi)\leq f(r)$ for every positive integer $r$.
3.4 Transitive fraternal augmentations
We now come to a second approach to approximate the generalized coloring numbers, which is based on Nešetřil and Ossona de Mendez’s transitive fraternal augmentation technique, see Chapter 7.4 of [111]. In the following, for clarity, we will call directed edges $(u,v)$ arcs.
Definition 3.27.
Let $G$ be a graph and let $r$ be a positive integer. An $r$-fraternity function is a function $w\colon V(G)\times V(G)\rightarrow\mathbb{N}\cup\{\infty\}$ such that for all distinct $u,v\in V(G)$ at least one of $w(u,v)$ and $w(v,u)$ is $\infty$, such that for all $uv\in E(G)$ either $w(u,v)=1$ or $w(v,u)=1$, and such that for all distinct $u,v\in V(G)$
  • 1.
    either $\min\{w(u,v),w(v,u)\}=1$,
  • 2.
    or $\min\{w(u,v),w(v,u)\}=\min_{z\in V\setminus\{u,v\}}(w(z,u)+w(z,v))$,
  • 3.
    or $\min\{w(u,v),w(v,u)\}>r$ and $\min_{z\in V\setminus\{u,v\}}(w(z,u)+w(z,v))>r$.
If $w$ is an $r$-fraternity function for every value of $r$, then $w$ is called a fraternity function.
Figure 12: A fraternity function on a path $P_{6}$. All missing pairs are assigned value $\infty$.
Definition 3.28.
Let $G$ be a graph, let $r$ be a positive integer and let $w$ be an $r$-fraternity function for $G$. For $1\leq i\leq r$, the directed graph $\vec{G}^{w}_{\leq i}$ on the same vertex set as $G$ containing all arcs $(u,v)$ with $w(u,v)\leq i$ is called an $i$-fraternal augmentation of $G$.
Obviously we have $\vec{G}^{w}_{\leq 1}\subseteq\vec{G}^{w}_{\leq 2}\subseteq\ldots\subseteq\vec{G}^{w}_{\leq r}$. The arcs of $\vec{G}^{w}_{\leq r}$ that are not arcs of $\vec{G}^{w}_{\leq 1}$ are called fraternal arcs.
We remark that fraternal augmentations in [111] are defined so that the fraternal arcs in the second and third item above are introduced between vertices that have a common in-neighbor, that is, in that work, it is required that $\min\{w(u,v),w(v,u)\}=\min_{z\in V\setminus\{u,v\}}(w(u,z)+w(v,z))$, or if $\min\{w(u,v),w(v,u)\}>r$ then $\min_{z\in V\setminus\{u,v\}}(w(u,z)+w(v,z))>r$. We prefer our definition, as it highlights the connection to the generalized coloring numbers. Let us make this connection explicit.
Lemma 3.29.
Let $G$ be a graph, let $r$ be a positive integer and let $\pi$ be a linear order of $V(G)$. Then $w\colon V(G)\times V(G)\rightarrow\mathbb{N}\cup\{\infty\}$ with
$w(u,v)=\begin{cases}i&\text{if $i\leq r$ and $v\in\operatorname{SReach}_{i}[G,\pi,u]\setminus\operatorname{SReach}_{i-1}[G,\pi,u]$}\\ \infty&\text{if $v\not\in\operatorname{SReach}_{r}[G,\pi,u]$}.\end{cases}$
is an $r$-fraternity function.
Proof.
It is clear that for all distinct $u,v\in V(G)$ at least one of $w(u,v)$ and $w(v,u)$ is $\infty$ and that for all $uv\in E(G)$ either $w(u,v)=1$ or $w(v,u)=1$. Now assume $w(u,v)=i$ for some $i\geq 2$, hence, $v\in\operatorname{SReach}_{i}[G,\pi,u]\setminus\operatorname{SReach}_{i-1}[G,\pi,u]$. Then there exists a path $P$ of length $i$ between $u$ and $v$ such that $v$ is minimum among $V(P)$ with respect to $\pi$ and such that all internal vertices of $P$ are larger than $u$. Let $z$ be the smallest vertex of $V(P)\setminus\{u,v\}$ and let $\ell_{1}$ be the distance between $z$ and $v$ on $P$ and let $\ell_{2}$ be the distance between $z$ and $u$ on $P$. Then $v\in\operatorname{SReach}_{\ell_{1}}[G,\pi,z]$ and $u\in\operatorname{SReach}_{\ell_{2}}[G,\pi,z]$. Furthermore, there does not exist $z^{\prime}$ and $d_{1},d_{2}$ with $d_{1}+d_{2}<i$ such that $v\in\operatorname{SReach}_{d_{1}}[G,\pi,z^{\prime}]$ and $u\in\operatorname{SReach}_{d_{2}}[G,\pi,z^{\prime}]$, as otherwise, we could concatenate the paths witnessing this to find a path of length $i-1$ between $u$ and $v$ such that $v$ is minimum among $V(P)$ and such that all internal vertices of $P$ are larger than $u$. Such a path does not exist as $v\not\in\operatorname{SReach}_{i-1}[G,\pi,u]$. In particular, $v\not\in\operatorname{SReach}_{\ell_{1}-1}[G,\pi,z]$ and $u\not\in\operatorname{SReach}_{\ell_{2}-1}[G,\pi,z]$, hence, $w(z,u)=\ell_{1}$ and $w(z,v)=\ell_{2}$ and there is no $z^{\prime}$ with $w(z^{\prime},u)+w(z^{\prime},v)<i$. ∎
Using Theorem 3.18, we get the following corollary.
Corollary 3.30.
Let $G$ be a graph and let $r$ be a positive integer. Then there exists an $r$-fraternity function $w$ such that $G^{w}_{\leq r}$ has maximum out-degree at most $(6r)^{r}\cdot\widetilde{\nabla}_{r-1}(G)^{3r}$.
Vice versa, we can turn fraternal augmentations into weak coloring orders as follows. We need the following property of fraternal augmentations that was (in variations) noted several times [53, 77, 111, 134, 144].
Lemma 3.31.
Let $G$ be a graph, let $r$ be a positive integer and let $w$ be an $r$-fraternity function of $G$. Let $u,v\in V(G)$ and assume that $u=v_{0},v_{1},\ldots,v_{\ell}=v$ is a path of length $\ell\leq r$ between $u$ and $v$ in $G$. Then one of the following holds.
  1. 1.
    There are indices $0<i_{1}<\ldots<i_{k}<\ell$ such that $u,v_{i_{1}},\ldots,v_{i_{k}},v$ is a directed path from $u$ to $v$ in $G^{w}_{\leq r}$ and for all $0\leq i_{m}<i_{n}\leq\ell$ we have $w(v_{i_{m}},v_{i_{n}})\leq i_{n}-i_{m}$.
  2. 2.
    There are indices $0<i_{1}<\ldots<i_{k}<\ell$ such that $v,v_{i_{k}},\ldots,v_{i_{1}},u$ is a directed path from $v$ to $u$ in $G^{w}_{\leq r}$ and for all $0\leq i_{m}<i_{n}\leq\ell$ we have $w(v_{i_{n}},v_{i_{m}})\leq i_{n}-i_{m}$.
  3. 3.
    There are indices $0<i_{1}<\ldots<i_{k}<\ell$ and an index $i_{j}\in\{i_{1},\ldots,i_{k}\}$ such that $u,v_{i_{1}},\ldots,v_{i_{j}}$ and $v,v_{i_{k}},\ldots,v_{i_{j}}$ are directed paths in $G^{w}_{\leq r}$ from $u$ to $v_{i_{j}}$ and from $v$ to $v_{i_{j}}$, respectively. Furthermore, for all $0\leq i_{m}<i_{n}\leq i_{j}$ we have $w(v_{i_{m}},v_{i_{n}})\leq i_{n}-i_{m}$ and for all $i_{j}\leq i_{m}<i_{n}\leq\ell$ we have $w(v_{i_{n}},v_{i_{m}})\leq i_{n}-i_{m}$.
Proof.
We prove the statement by induction on $\ell$. For $\ell=1$, we have $(u,v)\in E(\vec{G}^{w}_{\leq 1})$ if and only if $w(u,v)=1$, hence, one of the first two items is true. Now assume that $\ell>1$ and assume the statements hold for $v_{0},v_{1},\ldots,v_{\ell-1}$. Assume we are in the first case and there are indices $10<i_{1}<\ldots<i_{k}<\ell-$ 1 such that $u,v_{i_{1}},\ldots,v_{i_{k}},v_{\ell-1}$ is a directed path from $u$ to $v_{\ell-1}$ in $G^{w}_{\leq r}$ and that for all $10\leq i_{m}<i_{n}\leq\ell-$ 1 we have $w(v_{i_{m}},v_{i_{n}})\leq i_{n}-i_{m}$. We distinguish two cases. If $w(v_{\ell-1},v)=1$, $u,v_{i_{1}},\ldots,v_{i_{k}},v_{\ell-1},v$ is a directed path in $G^{w}_{r}$, hence the first item holds for $u=v_{0},v_{1},\ldots,v_{\ell}=v$. If $w(v,v_{\ell-1})=1$, then we set $v_{j}$ in the third item to $v_{\ell-1}$ and the third item holds for $u=v_{0},v_{1},\ldots,v_{\ell}=v$. The statements about the weight function follow directly from the hypothesis.
Now assume we are in the second case and there are indices $10<i_{1}<\ldots<i_{k}<\ell-$ 1 such that $v_{\ell-1},v_{i_{k}},\ldots,v_{i_{1}},u$ is a directed path from $v_{\ell-1}$ to $u$ in $G^{w}_{\leq r}$ and that for all $10\leq i_{m}<i_{n}\leq\ell-$ 1 we have $w(v_{i_{n}},v_{i_{m}})\leq i_{n}-i_{m}$. Again we distinguish two cases. If $w(v,v_{\ell-1})=1$, then the first case holds for $u=v_{0},v_{1},\ldots,v_{\ell}=v$. The additional statement about the weight function follows directly from the hypothesis. If $w(v_{\ell-1},v)=1$, then we have $w(v,v_{i_{k}})\leq\ell-i_{k}$ or $w(v_{i_{k}},v)\leq\ell-i_{k}$, as $w$ is a fraternity function and by assumption, we have $w(v_{\ell-1},v_{i_{k}})\leq\ell-i_{k}-1$. In the former case we have a directed path $v,v_{i_{k}},\ldots,v_{i_{1}},u$ from $v$ to $u$ in $G^{w}_{\leq r}$, that is, the second case holds for $u=v_{0},v_{1},\ldots,v_{\ell}=v$. In the latter case, we apply the argument again to the sequence $v,v_{i_{k}},\ldots,v_{i_{1}},u$ ($v_{i_{k}}$ taking the role of $v_{\ell-1}$). Again, we either get a directed path $v,v_{i_{k-1}},\ldots,v_{i_{1}},u$ from $v$ to $u$ in $G^{w}_{\leq r}$, or we apply the argument again. In some step, we either get a directed path $v,v_{i_{m}},\ldots,v_{i_{1}},u$ from $v$ to $u$ in $G^{w}_{\leq r}$ for some $1\leq i\leq k$, or a directed arc $(u,v)$ in $G^{w}_{\leq r}$. The additional property on the weight function is easily verified in each step.
The argument in the third case is analogous to the argument in the second case. ∎
Compare Lemma 3.31 with Lemma 2.15, which describes the local separation properties of the weak coloring numbers. We now use Lemma 3.31 to approximate $\operatorname{wcol}_{r}(G)$ by $r$-fraternal augmentations as follows.
Lemma 3.32 (adaptation of Lemma 6.7 of [77]).
Let $G$ be a graph, let $r$ be a positive integer and let $w$ be an $r$-fraternity function of $G$. Assume that $G^{w}_{\leq r}$ has maximum out-degree $\Delta$ and let $d=r\cdot\Delta^{r}$. Then $\operatorname{wcol}_{r}(G)\leq 4d^{2}$.
Proof.
Let $\vec{H}$ be the directed graph that we obtain by adding all edges $(u,v)$ such that there is a directed path of (weighted) length at most $r$ from $u$ to $v$ in $G^{w}_{\leq r}$. As $G^{w}_{\leq r}$ has maximum out-degree $\Delta$, $\vec{H}$ has out-degree at most $\sum_{i=1}^{r}\Delta^{i}\leq r\cdot\Delta^{r}=d$. Furthermore, Lemma 3.31 implies that for all paths $u=v_{0},v_{1},\ldots,v_{\ell}=v$ of length $\ell\leq r$ between $u$ and $v$ in $G$, either
  1. 1.
    $(u,v)\in E(\vec{H})$, or
  2. 2.
    $(v,u)\in E(\vec{H})$, or
  3. 3.
    there is an index $0<i<\ell$ such that $(u,v_{i}),(v,v_{i})\in E(\vec{H})$.
As $\vec{H}$ has out-degree at most $d$, the underlying undirected graph $H$ is $2d$-degenerate and there exists an order $\pi$ of $V(H)$ such that each vertex has at most $2d$ smaller neighbors. We claim that $\operatorname{wcol}_{r}(G,\pi)\leq 4d^{2}$.
To see this, we count for each vertex $u\in V(G)$ the number of end-vertices of paths of length at most $r$ from $u$ such that the end-vertex is the smallest vertex of the path. This number bounds $|\operatorname{WReach}_{r}[G,\pi,u)]|$.
By our above observation, for each such path $P$ with end-vertex $v$, we either have an edge $(u,v)$ or an edge $(v,u)$ or there is $z$ on the path and we have arcs $(u,z),(v,z)$ in $\vec{H}$.
By construction of the order, there are at most $2d$ edges $(u,v)$ or $(v,u)$ such that $v<_{\pi}u$. Furthermore, we have at most $d$ arcs $(u,z)$ in $\vec{H}$, as $u$ has out-degree at most $d$. For each such $z$ there are at most $2d$ edges $(v,z)$ such that $v<_{\pi}z$ by construction of the order. As $v\in\operatorname{WReach}_{r}[G,\pi,u)]$ requires that $v$ is the minimum vertex on $P$, these are the only pairs we have to count. Hence, in total we have $|\operatorname{WReach}_{r}[G,<,u]|\leq 2d+2d^{2}\leq 4d^{2}$. ∎
We can hence use Lemma 3.32 to find orders with good weak coloring properties, if we are able to find good $r$-fraternity functions (without applying Lemma 3.29 of course). As shown by Nešetřil and Ossona de Mendez [122], a simple greedy procedure yields good results.
Let $G$ be an undirected graph and let $\vec{G}_{1}$ be any orientation of $G$. We construct a sequence $\vec{G}_{1}\subseteq\vec{G}_{2}\subseteq\dots\subseteq G_{r}$ of directed graphs together with an $r$-fraternity function $w\colon V(G)\times V(G)\rightarrow\mathbb{N}\cup\{\infty\}$. We let $w(u,v)=1$ if $(u,v)\in E(\vec{G}_{1})$ and $w(u,v)=\infty$ if $(v,u)\in E(\vec{G}_{1})$. Now, assume $\vec{G}_{d}$ for $d\geq 1$ has been constructed and assume that $w$ is defined for all pairs $u,v$ with either $(u,v)$ or $(v,u)\in E(\vec{G}_{d})$ and undefined for all other pairs. We obtain the edge set $E(\vec{G}_{d+1})$ by adding the following edges to $E(\vec{G}_{d})$ as follows.
  • 1.
    If $u,v,z\in V(G)$ are such that $(z,u)\in E(\vec{G}_{i})$ and $(z,v)\in E(\vec{G}_{j})$ and $i+j=d+1$, and neither $(u,v)$ nor $(v,u)\in E(\vec{G}_{d})$, then we introduce exactly one of $(u,v)\in E(\vec{G}_{d+1})$ or $(v,u)\in E(\vec{G}_{d+1})$.
We define $w(u,v)=d+1$ if $(u,v)\in E(\vec{G}_{d+1})$ and $w(u,v)=\infty$ if $(v,u)\in E(\vec{G}_{d+1})$. Finally, we define $w(u,v)=\infty$ for all $u,v$ with $(u,v),(v,u)\not\in E(\vec{G}_{r})$.
Observation 3.33.
The function $w$ as defined above is an $r$-fraternity function.
Proof.
It is immediately from the construction that for all $u,v\in V(G)$ at least one of $w(u,v)$ and $w(v,u)$ is $\infty$. It remains to show that for all distinct $u,v\in V(G)$
  • 1.
    either $\min\{w(u,v),w(v,u)\}=1$,
  • 2.
    or $\min\{w(u,v),w(v,u)\}=\min_{z\in V\setminus\{u,v\}}(w(z,u)+w(z,v))$,
  • 3.
    or $\min\{w(u,v),w(v,u)\}>r$ and $\min_{z\in V\setminus\{u,v\}}(w(z,u)+w(z,v))>r$.
We prove the statement for all $u,v$ with $\min\{w(u,v),w(v,u)\}\leq d\leq r$, by induction on $d$. It is immediate that for all $uv\in E(G)$ either $w(u,v)=1$ or $w(v,u)=1$ and that there are no other pairs for which $w$ takes the value $1$. Now assume the statement holds for $1\leq d<r$ and fix some pair $u,v$ with $\min\{w(u,v),w(v,u)\}=d+1$, say $w(u,v)=d+1$. Then $(u,v)\in E(\vec{G}_{d+1})\setminus E(\vec{G}_{d})$. Hence, there is $z\in V(G)$ and $i,j$ with $i+j=d+1$ such that $(z,u)\in E(\vec{G}_{i})$ and $(z,v)\in E(\vec{G}_{j})$. Furthermore, there is no $z\in V(G)$ and $i,j$ with $i+j<d+1$ such that $(z,u)\in E(\vec{G}_{i})$ and $(z,v)\in E(\vec{G}_{j})$, as otherwise by construction, we would have $(u,v)$ or $(v,u)\in\vec{G}_{d}$. Hence, by induction hypothesis, $\min\{w(u,v),w(v,u)\}=\min_{z\in V\setminus\{u,v\}}(w(z,u)+w(z,v))$. With the same argument, we now get that $\min\{w(u,v),w(v,u)\}>r$ implies $\min_{z\in V\setminus\{u,v\}}(w(z,u)+w(z,v))>r$. ∎
It remains to specify how we obtain the orientation $\vec{G}_{1}$ and how we decide whether we introduce the edge $(u,v)$ or $(v,u)$ in each augmentation step. We can choose the orientation $\vec{G}_{1}$ to be the acyclic ordering derived from the degeneracy ordering of $G$. This is optimal up to factor $2$ if we aim to minimize the maximum out-degree $\Delta^{+}(\vec{G}_{1})$ of $\vec{G}_{1}$. Second, in any step $d+1$, we can orient the fraternal edges added in step $d+1$ by first collecting all potential fraternal edges in an auxiliary undirected graph $H_{d+1}$ and then again compute an acyclic orientation $\vec{H}_{d+1}$ that is optimal up to factor $2$. We then insert the arcs into $\vec{G}_{d+1}$ according to their orientation in $\vec{H}_{d+1}$.
It remains to prove that we obtain good bounds for the maximum out-degree of $\vec{G}_{r}$. The simplest way to establish such bounds is via shallow packings, which we introduce next.
Definition 3.34.
Let $G$ be a graph and let $k,t$ be positive integers. A collection $\mathcal{F}$ of subsets of $V(G)$ is a $(k,t)$-packing if $G[F]$ is connected and has radius at most $t$ for every $F\in\mathcal{F}$, and every vertex appears in at most $k$ sets of $\mathcal{F}$. The elements of $\mathcal{F}$ are called clusters. The induced packing graph $G[\mathcal{F}]$ has $\mathcal{F}$ as the set of vertices and two clusters $F,F^{\prime}\in\mathcal{F}$ are connected by an edge if they share a vertex or there are vertices $u\in F$ and $v\in F^{\prime}$ such that $uv\in E(G)$.
Definition 3.35.
Let $G,H$ be graphs. The lexicographic product of $G$ with $H$, denoted $G\bullet H$ is defined by
$V(G\bullet H)$ $=V(G)\times V(H)$
$E(G\bullet H)$ =$=\{(u,x)(v,y):uv\in E(G)\text{ or $u=v$ and $xy\in E(H)$}\}.$
Observe that $G\bullet K_{k}$ corresponds to the packing graph $G[\mathcal{F}]$, where $\mathcal{F}$ is the $(k,0)$-packing that contains $k$ copies of each vertex $v\in V(G)$. Vice versa, if $\mathcal{F}$ is a $(k,t)$-packing, then the induced packing graph $G[\mathcal{F}]$ is a depth-$t$ minor of the lexicographic product $G\bullet K_{k}$.
Nešetřil and Ossona de Mendez [111] and Har-Peled and Quanrud [80] proved that shallow packings of sparse graphs remain sparse. The following bounds are from [129].
Lemma 3.36 ([129]).
Let $G$ be a graph and let $\mathcal{F}$ be a $(k,t)$-packing of $G$. Then for every positive integer $r$,
$\nabla_{r}(G[\mathcal{F}])\leq\frac{k-1}{2}+(2(k-1)(2rt+r+t+1)+1)\cdot\nabla_{2rt+t+r}(G).$
In particular,
$\nabla_{r}(G\bullet K_{k})\leq\frac{k-1}{2}+(2(k-1)(r+1)+1)\cdot\nabla_{r}(G)\leq 2(k-1)(r+3)\cdot\nabla_{r}(G).$
Let us give some intuition how to now derive bounds for the fraternal augmentations. Let $\vec{G}_{1}\subseteq\vec{G}_{2}\subseteq\ldots\subseteq\vec{G}_{d}$ be defined as above and assume that $\vec{G}_{i}$ has maximum out-degree $d_{i}$. Recall that the orientation $\vec{G}_{1}$ is obtained from the acyclic ordering derived from the degeneracy ordering of $G$, hence, the out-degree $d_{1}$ is bounded. For $1<i\leq d$ and for each $v\in V(G)$ let $F^{i}_{v}\coloneqq\{u\leavevmode\nobreak\ :\leavevmode\nobreak\ (v,u)\in E(\vec{G}_{i})\cup\{v\}$. Let $\mathcal{F}^{i}=\{F^{i}(v):v\in V(G)\}$. For $k+\ell=d+1$ let $G^{k,\ell}$ be the packing graph $G[\mathcal{F}^{k}\cup\mathcal{F}^{\ell}]$. Observe that $\mathcal{F}^{k}\cup\mathcal{F}^{\ell}$ is a $(d+1,d_{k}+d_{\ell})$-packing. Let $G_{d+1}=\bigcup_{k+\ell=d+1}G^{k,\ell}$. Then $G_{d+1}$ has all fraternal augmentations of length $d+1$ as undirected edges. Now use Lemma 3.36 to derive that $\nabla_{0}(G_{d+1})$ is bounded and hence we can again find an orientation with small out-degree for the fraternal edges of $G_{d+1}$.
A stronger analysis of the augmentation process gives the following bounds. We remark that Reidl in [134] does not only consider fraternal but also transitive augmentations.
Theorem 3.37 (Theorem 16 of [134], reformulated).
For every graph $G$ the above augmentation process yields fraternal augmentation sequence $\vec{G}_{1},\vec{G}_{2},\ldots$ such that
$\Delta^{+}(\vec{G}_{d+1})\leq 2^{8\cdot 5^{d}}d^{6^{d}}(d+1)^{2^{d}}\cdot\big{(}\widetilde{\nabla}_{d+1}(G)\cdot\Delta^{+}(\vec{G}_{1})\big{)}^{3\cdot 5^{d-1}}.$
3.5 Pointer-structures and quantifier elimination
Lemma 3.31 makes quite clear that transitive fraternal augmentations are just the weak coloring numbers in disguise, and most current research works directly with the generalized coloring numbers instead of with augmentations. Nevertheless, the augmentations have been very important, especially in the highly influential quantifier elimination schema for first-order logic on classes of bounded expansion of Dvořák, Král’ and Thomas [53]. The recent formulation of results of this type is in terms of pointer structures (a structure with a signature that consists of unary relation and unary function symbols, see Section 2.4) that are guarded by [53], founded in the graph [74] or guided by the graph [119], and have found many applications at the intersection of sparse and structurally sparse graphs and logic, see e.g. [36, 43, 53, 74, 89, 119, 132, 140, 153]. Quantifier elimination over pointer structures in nowhere dense classes does not work as for bounded expansion classes as proved by Grobler et al. [74]. We will revisit the concept of quantifier elimination in Section 6 when discussing low treedepth colorings.
4 Tree-decomposable graphs
In this section, we turn our attention to upper bounds for the generalized coloring numbers on special graph classes. Special focus is put on tree-decomposable graph classes. Let us define the used concepts.
Let $\mathcal{T}=(T,\mathrm{Bag})$ be a rooted tree decomposition of $G$. For a non-root node $x$ with parent $y$, we define the adhesion of $x$ as $\mathrm{Adh}(x)=\mathrm{Bag}(x)\cap\mathrm{Bag}(y)$. If $x$ is the root, then we set $\mathrm{Adh}(x)=\emptyset$ by convention. The adhesion of the tree decomposition $\mathcal{T}=(T,\mathrm{Bag})$ is the maximum size of an adhesion set in $\mathcal{T}$, i.e., $\max_{x\in V(T)}|\mathrm{Adh}(x)|$. The torso of a node $x$ is the graph $\mathrm{Torso}(x)$ on vertex set $\mathrm{Bag}(x)$ where two vertices $u,v\in\mathrm{Bag}(x)$ are adjacent if and only if $uv\in E(G)$ or if there exists $y\neq x$ such that $u,v\in\mathrm{Bag}(y)$. Equivalently, $\mathrm{Torso}(x)$ is obtained from $G[\mathrm{Bag}(x)]$ by turning the adhesions of $x$ and of all children of $x$ into cliques. The margin of a node $x$ is the set $\mathrm{Margin}(x)=\mathrm{Bag}(x)\setminus\mathrm{Adh}(x)$. We say that $\mathcal{T}$ is a tree decomposition over a class $\mathscr{D}$ of graphs if $\mathrm{Torso}(x)\in\mathscr{D}$ for every node $x$ of $T$.
Figure 13: A tree decomposition of a graph $G$. The torso of $x$ is obtained from the graph induced by $\mathrm{Bag}(x)$ by turning the adhesions into cliques.
We are going to use the separator properties of tree-decomposable graphs described in Lemma 2.3 to construct good generalized coloring orders for classes $\mathscr{C}$ that admit tree decompositions of small adhesion over other graph classes $\mathscr{D}$. The orders for graphs from the class $\mathscr{D}$ can be naturally combined with the partial order induced by the tree decomposition to obtain good orders for graphs from $\mathscr{C}$.
Weak reachability for vertices from different bags of a decomposition is guided by the interaction through the adhesion sets. The base case capturing this situation, which also guides the general situation of tree decomposable graphs, is the case of bounded treewidth. It was proved by Grohe et al. [76] that the weak $r$-coloring numbers are bounded polynomially in $r$ on every class of bounded treewidth.
Theorem 4.1 ([76]).
Let $G$ be a graph of treewidth at most $k$ and let $r$ be a positive integer. Then $\operatorname{wcol}_{r}(G)\leq\binom{r+k}{k}$.
By combining the partial order induced by the tree decomposition with the orders for graphs from a class $\mathscr{C}$, we then obtain the following result.
Theorem 4.2.
Let $\mathscr{C}$ be a class of graphs that is tree decomposable with adhesion at most $k$ over a class $\mathscr{D}$. Let $G\in\mathscr{C}$ and let $r$ be a positive integer. Then
  • 1.
    $\operatorname{adm}_{r}(G)\leq k+\operatorname{adm}_{r}(\mathscr{D})$,
  • 2.
    $\operatorname{col}_{r}(G)\leq k+\operatorname{col}_{r}(\mathscr{D})$, and
  • 3.
    $\operatorname{wcol}_{r}(G)\leq\binom{k+r}{k}\cdot(\operatorname{wcol}_{r}(\mathscr{D})+k)$.
In the rest of this section, we present a proof of Theorem 4.2, which is based on a construction of Pilipczuk and Siebertz [127]. In the next section, we present a construction of Van den Heuvel et al. for $H$-minor free graphs [150], showing that the generalized coloring numbers are polynomially bounded on such classes. Based on the decomposition theorem for $H$-topological-minor-free graphs we give new bounds for the weak $r$-coloring numbers on $H$-topological-minor-free graphs. Finally, we present Dvořák’s characterization of graphs $G$ with bounded $\operatorname{adm}_{\infty}(G)$ [49].
Let $\mathcal{T}=(T,\mathrm{Bag})$ be a tree decomposition of $G$. For every vertex $u$ of $G$, let $x(u)$ be the unique $\leq_{T}$-minimal node of $T$ with $u\in\mathrm{Bag}(x)$. We define a quasi-order $\leq_{\mathcal{T}}$ on the vertex set of $G$ by $u\leq_{\mathcal{T}}v$ if and only if $x(u)\leq_{T}x(v)$. We can use any linearization of $\leq_{\mathcal{T}}$ to obtain an order of $V(G)$ achieving the bounds of Theorem 4.1 for the weak coloring numbers. We now extend this order to tree-decomposable graphs.
Lemma 4.3.
For every vertex $u$ of $G$, the node $x(u)$ is the unique node of $T$ whose margin contains $u$.
Note that by the lemma, $\{\mathrm{Margin}(x)\}_{x\in V(T)}$ is a partition of the vertex set of $G$ and the margins of nodes of $T$ are exactly the classes of equivalence in the quasi-order $\leq_{\mathcal{T}}$ on $V(G)$. Now assume that $G$ admits a tree decomposition over a class $\mathscr{D}$. Then, for each torso $\Gamma(x)$ there exists a linear order $\pi_{x}$ with $\operatorname{adm}_{r}[\Gamma(x),\pi_{x}]\leq\operatorname{adm}_{r}(\mathscr{D})$, $\operatorname{col}_{r}[\Gamma(x),\pi_{x}]\leq\operatorname{col}_{r}(\mathscr{D})$, and $\operatorname{wcol}_{r}[\Gamma(x),\pi_{x}]\leq\operatorname{wcol}_{r}(\mathscr{D})$, respectively (these could be different for each measure, but for simplicity of notation we assume they are the same). We let $\pi^{\prime}_{x}$ be the restriction of these orders to $\mathrm{Margin}(x)$. We define $\pi$ such that it linearizes $\leq_{\mathcal{T}}$ and respects the orders $\pi^{\prime}_{x}$.
We now prove that the constructed order has the claimed properties. The proof follows a proof from  [127]. We need one more concept.
Definition 4.4.
The skeleton of $G$ over $\mathcal{T}$ is the directed graph with vertex set $V(G)$ and arcs $(u,v)$ such that there is $x\in V(T)$ with $u\in\mathrm{Margin}(x)$ and $v\in\mathrm{Adh}(x)$.
Note that if $(u,v)$ is an arc in the skeleton of $G$, then in particular $v<_{\mathcal{T}}u$, equivalently $x(v)<_{T}x(u)$. This implies that the skeleton is always acyclic (that is, it is a DAG).
Now we observe that weak reachability across adhesions implies reachability in the skeleton.
Lemma 4.5 (Lemma 23 of [127]).
Let $u,v$ be vertices of $G$ with $v<_{\mathcal{T}}u$ and let $P$ be a path in $G$ with endpoints $u$ and $v$ such that every vertex $w$ of $P$ apart from $v$ satisfies $v<_{\mathcal{T}}w$. Then there exists a directed path $Q$ in the skeleton leading from $u$ to $v$ and satisfying $V(Q)\subseteq V(P)$.
Now the key observation is that if the considered tree decomposition has small adhesion, then every vertex reaches only a small number of vertices via short paths in the skeleton. The argument is essentially the same as Grohe et al.’s argument that graphs of treewidth $k$ have polynomial weak $r$-coloring number [76].
Lemma 4.6 (Lemma 24 of [127]).
Let $u\in V(G)$ and $r\in\mathbb{N}$. Then there exist at most $\binom{r+k}{k}$ vertices of $G$ that are reachable by a directed path of length at most $r$ from $u$ in the skeleton of $G$ over $\mathcal{T}$.
We can now use Lemma 4.6 to prove Theorem 4.2.
Proof of Theorem 4.2.
We first show the claimed bounds for $\operatorname{adm}_{r}[G,\pi]$ and $\operatorname{col}_{r}[G,\pi]$, where $\pi$ is the above constructed order with corresponding vertex enumeration $v_{1},\ldots,v_{n}$. Consider any $v_{i}\in V(G)$ and let $x=x(v_{i})$. Then a maximal $v_{i}-V_{i}$-fan has its ends in $\mathrm{Margin}(x)\cup\mathrm{Adh}(x)$. Observe that if a path in $G$ leaves $\mathrm{Margin}(x)$ to a larger bag $y$, i.e., $y$ is a child of $x$, and enters again so that it ends in a smaller vertex, then this is also a path in the torso $\Gamma(x)$, not leaving $\mathrm{Margin}(x)$. Hence, we have $\operatorname{adm}_{r}[G,\pi,v]\leq\operatorname{adm}_{r}(\mathscr{D})+k$. The same argument shows that $\operatorname{col}_{r}[G,\pi,v]\leq\operatorname{col}_{r}(\mathscr{D})+k$ for all $v\in V(G)$.
Finally, for $\operatorname{wcol}_{r}(G)$, let $v\in V(G)$ and let $P$ be a path of length at most $r$ between $u$ and $v$ such that $u$ is minimum on $P$. Let $z$ be minimum on $P$ such that every vertex $w$ of $P$ apart from $z$ satisfies $z<_{\mathcal{T}}w$, or $z=u$ if no such $z$ exists. Then $u\in\operatorname{WReach}_{r}[G[\mathrm{Margin}(x(z)),\pi_{x},z]\cup\mathrm{Adh}(x(z))$. Using Lemma 4.5 and Lemma 4.6 we conclude that $x(z)$ can take at most $\binom{r+k}{k}$ different values, and hence, we have $\operatorname{WReach}_{r}[G,\pi,v]\leq\binom{r+k}{k}\cdot(\operatorname{wcol}_{r}(\mathscr{D})+k)$. ∎
4.1 H-minor-free graphs
We now turn to $H$-minor-free graphs. These graphs are tree decomposable over almost embeddable graphs by the famous structure theorem of Robertson and Seymour [137], however, we do not follow this route to obtain the best bounds for the generalized coloring numbers. We rather follow the approach of flat chordal partitions, which were also later the inspiration for the planar product structure theorem.
Recall Definition 2.4 up to Definition 2.7, defining simplicial vertices, perfect elimination orders, chordal graphs, the back-degree of a vertex and the width of an elimination order.
Definition 4.7.
Let $G$ be a graph. A partition $\mathcal{P}=(V_{1},\ldots,V_{\ell})$ of $V(G)$ is a connected partition if each $V_{i}$ induces a connected subgraph of $G$. We denote the minor of $G$ obtained from $G$ by contracting each $V_{i}$ to a single vertex $v_{i}$ by $G/\mathcal{P}$. We inherit all the above definitions for connected partitions by identifying a part $V_{i}$ with the vertex $v_{i}$ in (the ordered graph) $G/\mathcal{P}$.
With this definition, we derive for example the following lemma.
Lemma 4.8.
Let $G$ be a graph and $\mathcal{P}$ be a connected partition of $G$. Then the treewidth of $G/\mathcal{P}$ is bounded from above by the width of $\mathcal{P}$.
[…]
Definition 4.9.
Let $G$ be a graph and let $f:\mathbb{N}\to\mathbb{N}$ be a function. A set $W\subseteq V(G)$ $f$-spreads in $G$ if, for every $r\in\mathbb{N}$ and $v\in V(G)$, we have
$|N^{G}_{r}(v)\cap W|\leq f(r).$
A partition $V_{1},\ldots V_{\ell}$ of $G$ is $f$-flat if each $V_{i}$ $f$-spreads in $G-\bigcup_{1\leq j<i}V_{j}$.
Given an $f$-flat connected partition $\mathcal{P}=(V_{1},\ldots,V_{\ell})$ of a graph $G$ of small width we can now define orders with good generalized coloring properties. First choose an arbitrary linear order on the vertices of each set $V_{i}$. Now let $\pi$ be the linear extension of that order where for $v\in V_{i}$ and $w\in V_{j}$ with $i<j$ we define $\pi(v)<\pi(w)$.
Lemma 4.10 (Lemma 3.3 of [150]).
Let $\mathcal{P}=(V_{1},\ldots,V_{\ell})$ be a connected partition of a graph $G$, and let $\pi$ be an order defined as above. For an integer $i$, $1\leq i\leq\ell$, let $G^{\prime}=G-\bigcup_{1\leq j<i}V_{j}$. Then for every $r\in\mathbb{N}$ and every $v\in V(G)$
$\operatorname{SReach}_{r}[G,\pi,v]\cap V_{i}\>\subseteq\>N_{r}^{G^{\prime}}[v]\cap V_{i}\text{, and}$
$\operatorname{WReach}_{r}[G,L,v]\cap V_{i}\>\subseteq\>N_{r}^{G^{\prime}}[v]\cap V_{i}.$
Now it is easy to give upper bounds for $\operatorname{col}_{r}(G)$ and (using Theorem 4.1) for $\operatorname{wcol}_{r}(G)$ in terms of the width of a flat connected partition.
Lemma 4.11 (Lemma 3.4 of [150]).
Let $f:\mathbb{N}\to\mathbb{N}$ and let $r,k\in\mathbb{N}$. Let $G$ be a graph that admits an $f$-flat connected partition of width $k$. Then we have
$\operatorname{col}_{r}(G)\>\leq\>(k+1)\cdot f(r).$
Lemma 4.12 (Lemma 3.5 of [150]).
Let $f:\mathbb{N}\to\mathbb{N}$ and let $r,k\in\mathbb{N}$. Let $G$ be a graph that admits an $f$-flat connected partition of width $k$. Then we have
$\operatorname{wcol}_{r}(G)\>\leq\>\binom{r+k}{k}\cdot f(r).$
It remains to prove that $H$-minor-free graphs admit flat connected partitions of small width. The proof goes back to a decomposition of $H$-minor-free graphs due to Abraham [1], which is there called a cop-decomposition. This name is inspired by an earlier result of Andreae [6], which uses a cops-and-robber game argument to establish the existence of decompositions of $H$-minor-free graphs along isometric paths, which then leads to the desired flat partitions of small width.
Definition 4.13.
A path $P$ in a graph $G$ is an isometric path if $P$ is a shortest path between its endpoints. We call a partition $\mathcal{P}=(V_{1},\ldots,V_{\ell})$ an isometric paths partition if each $V_{i}$ induces an isometric path in $G-\bigcup_{1\leq j<i}V_{j}$.
Note that an isometric paths partition is in particular a connected partition. Note furthermore, that the vertex sets $V_{i}$ for $i>0$ not necessarily induce isometric paths in $G$, but only in $G$ after the earlier paths have been deleted. It will be the key to the planar product structure theory to partition the graph into globally isometric paths. The next lemma shows that isometric paths partitions are flat partitions.
Lemma 4.14.
[…]
Finally, based on the arguments of Abraham and Andreae [1, 6] it was shown by Van den Heuvel et al. [150] that for every graph that excludes $K_{t}$ as a minor one can find a flat connected partition where each part consists of at most $t-3$ isometric paths and which is of width at most $t-2$.
Lemma 4.15 ([150]).
Let $t\geq 4$ and let $f:\mathbb{N}\to\mathbb{N}$ be the function $f(r)=(t-3)(2r+1)$. Let $G$ be a graph that excludes $K_{t}$ as a minor. Then there exists a connected $f$-flat partition of $G$ of width at most $t-2$.
This immediately leads to the following theorems.
Theorem 4.16 ([150]).
Let $t\geq 4$. For every graph $G$ that excludes $K_{t}$ as a minor, we have
$\operatorname{col}_{r}(G)\>\leq\>\binom{t-1}{2}{}\cdot{}(2r+1).$
Theorem 4.17 ([150]).
Let $t\geq 4$. For every graph $G$ that excludes $K_{t}$ as a minor, we have
$\operatorname{wcol}_{r}(G)\>\leq\>\binom{r+t-2}{t-2}{}\cdot{}(t-3)(2r+1)\in\mathcal{O}(r^{\,t-1}).$
Van den Heuvel and Wood [151] started investigating whether the structure of the excluded minor can give even more insights on the generalized coloring numbers and they proved that for every $H$-minor free graph $G$ we have $\operatorname{wcol}_{r}(G)\in\mathcal{O}(r^{|vc(H)+1|})$, where $vc(H)$ is the vertex cover number of $H$. This result was recently improved by Dujmović et al. [46] who proved that there exists an exponential function $g$ such that all $H$-minor free graphs $G$ satisfy $\operatorname{wcol}_{r}(G)\in\mathcal{O}(r^{g(\operatorname{td}(H)})$. Hodor et al. [83] improved this result even more and showed that $H$-minor free graphs $G$ satisfy $\operatorname{wcol}_{r}(G)\in\mathcal{O}(r^{\operatorname{td}(H)})$. In fact, they prove that the exponent is bounded by the rooted $2$-treedepth of $H$ (a measure the authors introduced in [83]), and which is defined as follows.
Definition 4.18.
The rooted $2$-treedepth of a graph $H$, $\operatorname{rtd}_{2}(H)$, is defined recursively as follows: $\operatorname{rtd}_{2}(H)$ is
  • 1.
    00 if $H$ is empty,
  • 2.
    $1$ if $H$ is a single-vertex graph, and
  • 3.
    the minimum of $\max\{\operatorname{rtd}_{2}(H[A]),\operatorname{rtd}_{2}(H[B\setminus A])+|A\cap B|\}$ over all separations $(A,B)$ of $H$ of order at most one with $A\neq\emptyset$ and $B\setminus A\neq\emptyset$.
Theorem 4.19 ([83]).
For every positive integer $t$ and for every graph $H$ with $\operatorname{rtd}_{2}(H)\leq t$ there exists an integer $c$ such that for every $H$-minor free graph $G$ and every integer $r\geq 2$ we have
$\operatorname{wcol}_{r}(G)\leq c\cdot r^{t-1}\cdot\log r.$
4.2 Planar graph and graphs of bounded genus
In this section, we look at upper bounds for $\operatorname{col}_{r}(G)$ and $\operatorname{wcol}_{r}(G)$ when $G$ is a graph with bounded genus. Since for every genus $g$, there exists a $t$ such that every graph with genus at most $g$ does not contain $K_{t}$ as a minor, we could use Theorem 4.16 to obtain upper bounds for the generalized coloring numbers of such graphs. However, one can obtain significantly better bounds. First, it was shown that planar graphs have isometric paths partitions of width at most $2$.
Lemma 4.20 ([150]).
Every planar graph $G$ has an isometric paths partition of width at most 2.
For a graph of genus $g$, we first find isometric paths that we can delete to bring down the genus to 00, that is, to arrive at a planar graph. For every graph of genus $g>0$ there exists a non-separating cycle $C$ that consists of two isometric paths such that $G-C$ has genus $g-1$, see e.g. the textbook of Mohar [106]. By arranging these vertices first in the linear order, we obtain the following bounds.
Theorem 4.21 ([150]).
For a planar graph $G$, we have
$\operatorname{wcol}_{r}(G)\leq\binom{r+2}{2}\cdot(2r+1)\in\mathcal{O}(r^{3}),$
and for a graph $G$ with genus $g$, we have
$\operatorname{wcol}_{r}(G)\leq\biggl{(}2g+\binom{r+2}{2}\biggr{)}\cdot(2r+1)\in\mathcal{O}(r^{3}+gr).$
In case the treewidth of $G$ is bounded the following improved bounds hold, as proved by Hodor et al. [83].
Theorem 4.22 ([83]).
For all non-negative integers $g$ and $w$ there exists an integer $c$ such that for every graph $G$ of Euler genus at most $g$ and with treewidth at most $w$ we have
$\operatorname{wcol}_{r}(G)\leq c\cdot r^{2}\cdot\log r.$
This bound is tight for planar graphs of bounded treewidth. It is one of the main open problems whether in general the weak coloring numbers of planar graphs is bounded by $\mathcal{O}(r^{2}\cdot\log r)$.
Problem 4.23.
Is it true that $\operatorname{wcol}_{r}(G)\in\mathcal{O}(r^{2}\log r)$ for every planar graph $G$?
From the isometric paths partitions for planar graphs, one would already obtain good bounds also for the strong coloring numbers. A more careful construction (also based on isometric paths partitions) and analysis gives the following bounds.
Theorem 4.24 ([150]).
For a planar graph $G$, we have
$\operatorname{col}_{r}(G)\leq 5r+1,$
and for a graph $G$ with genus $g$, we have
$\operatorname{col}_{r}(G)\leq(2g+3)\cdot(2r+1).$
For quite a while it was an open question whether the $r$-admissibility of planar graphs could even be sublinear. This was eventually proved by Nederlof, Pilipczuk and Wegrzycki [109].
Theorem 4.25 ([109]).
For a planar graph $G$ we have
$\operatorname{adm}_{r}(G)\in\mathcal{O}(r/\log r).$
This bound is tight as stated in the next section. The proof of Theorem 4.25 is based on a theorem of Koebe, showing that every planar graph has a so-called coin model. A coin model associates to every vertex $v$ of a graph $G$ a disc (circle) $D(v)$ in the plane such that the discs are pairwise internally disjoint, and whenever two vertices $u$ and $v$ are adjacent in $G$, the discs $D(u)$ and $D(v)$ touch. Given a coin model, one can naturally derive an order by non-increasing radii of the associated discs, which yields the claimed bounds for the $r$-admissibility.
4.3 H-topological-minor-free graphs
To study $H$-topological-minor-free graphs we are coming back to tree decompositions. We are going to apply the following structure theorem, proved in this quantitative form by Erde and Weißauer [58]. A result of this type was first proved by Grohe and Marx [78], and then improved by Dvořák [49].
Theorem 4.26 ([58]).
[…]
  • 1.
    has fewer than $t^{2}$ vertices of degree at least $2t^{4}$, or
  • 2.
    excludes $K_{2t^{2}}$ as a minor.
We immediately derive the following corollary.
Corollary 4.27.
[…]
  • 1.
    $\operatorname{adm}_{r}(G)\leq t^{2}+(t^{2}+2t^{4})+4t^{4}(2r+1)$
  • 2.
    $\operatorname{col}_{r}(G)\leq t^{2}+(t^{2}+2t^{4r})+4t^{4}(2r+1)$
  • 3.
    $\operatorname{wcol}_{r}(G)\leq(4t^{4}(2r+2))^{r}$.
By a classical result of [15, 94] graphs that exclude $K_{t}$ as a topological minor have edge density $\mathcal{O}(t^{2})$, hence, satisfy ~$\widetilde{\nabla}_{r}(G)\in\mathcal{O}(t^{2})$ for all non-negative integers $r$. Hence, these bounds even improve the bounds we would get for the $r$-admissibility from Theorem 3.18.
4.4 Graph product structure theory
An isometric paths partition $\mathcal{P}=(V_{1},\ldots,V_{\ell})$ of small width provides already some insight into the treelike structure of a graph. The main shortcoming of these partitions is that a path induced by $V_{i}$ is not globally isometric but only isometric in the subgraph $G-\bigcup_{j<i}V_{j}$. In the study of $p$-centered colorings (which we will consider in Section 6) it was proved by Pilipczuk and Siebertz that planar graphs actually admit isometric paths partitions of width at most $8$ such that each path of the partition is a globally isometric path [127]. This result was further strengthened by Dujmović et al. [47], who showed the following result. Consider a tree $T$ rooted at a node $r$. A path $v_{1},\ldots,v_{t}$ in $T$ is called vertical if for some $d\geq 0$ and for all $1\leq i\leq t$ we have $\mathrm{dist}_{T}(v_{i},r)=d+i$.
Theorem 4.28 ([47]).
Let $T$ be a rooted spanning tree in a connected planar graph $G$. Then $G$ has a partition $\mathcal{P}$ into vertical paths in $T$ such that $G/\mathcal{P}$ has treewidth at most $8$.
Note that if $T$ is a BFS spanning tree, then all vertical paths in $T$ are isometric paths. Hence, the result yields a strengthening of the result of [127] that has led to the groundbreaking result of Dujmović et al. [47], the planar product structure theorem.
Definition 4.29.
[…]
Figure 14: The strong product of two paths.
Theorem 4.30 ([47]).
Every planar graph is a subgraph of
  • 1.
[…]
  • 2.
[…]
The planar product theorem had an immense impact and was the basis to solve major open problems and to make progress on longstanding open problems. For example, it was used by Dujmović et al. [47] to show that planar graphs have queue layouts with a bounded number of queues; this was a longstanding open problem. It was used by Dujmović et al. [45] that planar graphs can be nonrepetitively colored with a bounded number of colors; another longstanding open problem. We refer to the recent survey [52] for further applications and product structure theorems for various graph classes.
4.5 $\infty$-admissibility
Let $G$ be a graph, $v\in V(G)$ and $A\subseteq V(G)$. Recall that a $v-A$ fan is a set of paths $P_{1},\ldots P_{k}$ with one endpoint in $v$ and the other endpoint in $A$ and which are internally vertex disjoint from $A$ such that $V(P_{i})\cap V(P_{j})=\{v\}$ for all $i\neq j$. A $k$-fan set in $G$ is a set $A\subseteq V(G)$ such that for every $a\in A$, there exists a depth-$r$ $a-A$ fan of order $k$ in $G$. According to Theorem 3.14 the $\infty$-admissibility of $G$ is equal to the largest number $k$ such that there exists a $k$-fan set in $G$.
We define a related concept. Let $G$ be a graph and $k\in\mathbb{N}$. As set $X\subseteq V(G)$ of size at least $k$ is $k$-inseparable if no two vertices of $X$ can be separated in $G$ by deleting fewer than $k$ vertices. A maximal such set is called a $k$-block. The maximum integer $k$ such that $G$ contains a $k$-block is the block number of $G$, denoted by $\beta(G)$. Weissauer [155] proved the following.
Theorem 4.31 ([155]).
For every graph $G$
$\frac{\operatorname{adm}(G)+1}{2}\leq\beta(G)\leq\operatorname{adm}(G).$
Weissauer [155] proved the following decomposition theorems.
Theorem 4.32.
If $G$ has no $(k+1)$-block, then $G$ has a tree decomposition of adhesion less than $k$ in which every torso has at most $k$ vertices of degree at least $2k(k-1)$.
Theorem 4.33.
If $G$ has a tree decomposition in which every torso has at most $k$ vertices of degree at least $k$, then G has no $(k+1)$-block.
We conclude the following corollary, which was first proved by Dvořák [49]
Corollary 4.34.
A graph $G$ has bounded $\operatorname{adm}_{\infty}(G)$ if and only if it admits a tree decomposition with bounded adhesion over a class of bounded degree with a bounded number of apex vertices (vertices of arbitrary degree).
5 Lower bounds
We now turn to more sophisticated lower bounds for the generalized coloring numbers. The first result by Grohe et al. [76] shows that the upper bounds of Theorem 4.1 for the weak coloring numbers for graphs of bounded treewidth are optimal.
Theorem 5.1 ([76]).
For every $k\geq 1$, $r\geq 1$, there is a family of graphs $G^{k}_{r}$ with $\operatorname{tw}(G^{k}_{r})=k$, such that $\operatorname{wcol}_{r}(G^{k}_{r})=\binom{r+k}{k}$.
This result shows also that the upper bounds of Theorem 4.17 for $K_{t}$-minor free graphs are almost optimal. Furthermore, it verifies an exponential gap between $\operatorname{wcol}_{r}$ and $\operatorname{col}_{r}$.
Corollary 5.2 ([76]).
For every $k\geq 1$, $r\geq 1$, there is a graph $G^{k}_{r}$ such that $\operatorname{col}_{r}(G^{k}_{r})=k+1$ and $\operatorname{wcol}_{r}(G^{k}_{r})\geq\big{(}\frac{\operatorname{col}_{r}(G^{k}_{r})}{r}\big{)}^{r}$.
As recently shown by Hodor et al. [83], if a minor-closed class $\mathscr{C}$ contains graphs with $\operatorname{wcol}_{r}(G)\in\omega(r^{t}\log r)$, then it contains $G_{r}^{t}$ for every non-negative integer $r$. As a consequence they obtain that for minor-closed classes the class of graph $G_{r}^{t}$ is, up to an $\mathcal{O}(r\log r)$ factor, the unique construction of graphs with large weak coloring numbers.
By Lemma 3.11 for all graphs $G$ and positive integers $r$ we have $\operatorname{col}_{4r+1}(G)\geq\nabla_{r}(G)$. It is easy to see that every graph of girth greater than $8r$ and with minimum degree at least $d+1$ satisfies $\nabla_{2r}(G)\geq d^{r}$. As the $r$-admissibility is trivially bounded by the maximum degree, we conclude that there is also an exponential gap between $\operatorname{col}_{r}$ and $\operatorname{adm}_{r}$, as witnessed by $d$-regular high girth graphs. A stronger bound was established in [76].
Theorem 5.3 ([76]).
Let $G$ be a $d$-regular graph of girth at least $14g+$ 1, where $d\geq 7$. Then for every $r\leq g$,
$\operatorname{col}_{r}(G)\geq\frac{d}{2}\big{(}\frac{d-2}{4}\big{)}^{2^{\lfloor\log r\rfloor}-1}$
For the weak coloring numbers we have the following lower bounds on regular high-girth graphs.
Theorem 5.4 ([76]).
Let $G$ be a $d$-regular graph of girth at least $12g+$ 1, where $d\geq 4$. Then for every $r\leq g$,
$\operatorname{wcol}_{r}(G)\geq\frac{d}{d-3}\left(\Big{(}\frac{d-1}{2}\Big{)}^{r}-1\right)\,.$
We can also give a concrete example of a graph class with polynomial (and in fact even linear) expansion that has super-polynomial weak coloring numbers. For a graph $G$ denote by $G^{(r)}$ the exact $r$-subdivision of $G$, that is, the graph obtained from $G$ by replacing every edge by a path of length $r+1$ (with $r$ vertices on it). Recall the question of Joret and Wood, formulated in Problem 3.20, whether classes of polynomial expansion have polynomial strong coloring numbers.
Theorem 5.5 ([76]).
The class $\mathscr{C}=\{G^{(6\operatorname{tw}(G))}\leavevmode\nobreak\ :\leavevmode\nobreak\ G$ graph}}\}} has polynomial (even linear) expansion and super-polynomial weak coloring numbers.
Finally, let us mention that the bounds for the $r$-admissibility on planar graphs presented in Theorem 4.25 are tight. The construction of Dvořák and Siebertz [55] is presented explicitly in the work of Nederlof et al. [109].
Theorem 5.6 ([55]).
For every planar graph $G$ we have
$\operatorname{adm}_{r}(G)\in\Omega(r/\log r).$
6 Structural decompositions
Tree decompositions offer a way of globally decomposing a graph along small separators into well behaved pieces. On the other hand, the generalized coloring numbers offer rather local insights. In this section we will study another way of decomposing graphs into well behaved pieces. These turn out to be equivalent to the generalized coloring numbers but offer different insights. The structural decompositions we discuss here were first introduced by DeVos et al. [33] in terms of proper vertex colorings. For a positive integer $p$, a $p$-treewidth coloring of a graph $G$ is a vertex coloring such that the subgraph induced by any $i\leq p$ color classes has treewidth at most $i-1$. Hence, every single color class induces a graph of treewidth 00, that is, an edgeless graph. In particular, the coloring is a proper vertex coloring. A class of graphs admitting a $p$-treewidth coloring for every positive integer $p$ with a number of colors depending only on $p$ is said to admit low treewidth colorings. De Vos et al. [33] proved that every class excluding some graph as a minor admits low treewidth colorings. Not much later, Nešetřil and Ossona de Mendez proved that these classes even admit low treedepth colorings, then introduced classes with bounded expansion and proved [114] that they are characterized by the existence of low treedepth colorings. Finally, Zhu [158] provided an alternate characterization of classes with bounded expansion by the boundedness of generalized coloring numbers. The terminology has changed from structural colorings to structural decompositions and after presenting Zhu’s result we will switch to this terminology.
6.1 p-centered colorings
Let us begin with the two key definitions.
Definition 6.1.
For a positive integer $p$, a $p$-centered coloring of a graph $G$ is a vertex coloring of $G$ such that for any connected subgraph $H\subseteq G$, either some color appears exactly once in $H$, or $H$ gets at least $p$ colors. The least number of colors in a $p$-centered coloring of $G$ is denoted $\chi_{p}(G)$.
A centered coloring of $G$ is a vertex coloring such that, for any connected subgraph $H$ some color appears exactly once in $H$. It is easy to see that the treedepth of $G$ is equal to the minimum number of colors in a centered coloring of $G$.
Definition 6.2.
For a positive integer $p$, a $p$-treedepth coloring of a graph $G$ is a vertex coloring of $G$ such that the subgraph induced by any $i\leq p$ color classes has treedepth at most $i$.
Figure 15: A $3$-centered coloring of a grid with $5$ colors.
The following lemma by Nešetřil and Ossona de Mendez [113] shows that $(p+1)$-centered colorings yield $p$-treedepth colorings.
Lemma 6.3 ([113]).
A $q$-centered coloring for $q>p$ is a $p$-treedepth coloring.
The converse is not true, however, the existence of a $p$-treedepth coloring with few colors implies the existence of a $p$-centered coloring with few colors. We did not find a reference for this fact, but we believe that it is folklore in the community. We include a proof for completeness.
Lemma 6.4.
Let $p\geq 2$. If $G$ admits a $p$-treedepth coloring with $k$ colors, then $G$ admits a $(p+1)$-centered coloring with $k\bigl{(}2p\cdot\binom{k-1}{p-1}+1\bigr{)}<2pk^{p}$ colors.
Proof.
Fix a $p$-treedepth coloring $c$ of $G$. Let $\vec{H}$ be the directed graph with vertex set $V(G)$, where $uv\in E(\vec{H})$ if $v$ is an ancestor of $u$ for some choice $I$ of $p$ colors including $c(u)$ in the treedepth decomposition $Y_{I}$ for the colors $I$. There are $\binom{k-1}{p-1}$ such choices for $I$ and in each $Y_{I}$ vertex $u$ has at most $p$ ancestors. Then, every vertex has outdegree at most $p\cdot\binom{k-1}{p-1}$ hence has chromatic number at most $12p\cdot\binom{k-1}{p-1}+$ 1. Let $\gamma$ be a proper coloring of $H$. We consider the product coloring $v\mapsto(c(v),\gamma(v))$. Assume $H$ is a connected subgraph of $G$ with at most $p$ colors. Let $J=\{c(v)\colon v\in V(H)\}$. Then $|J|\leq p$, thus $H$ is included in the subgraph $G_{I}$ of $G$ induced by a set $I\supseteq J$ of $p$ $c$-colors. As $H$ is connected, it contains a vertex $r$ that is an ancestor of all the vertices of $H$. By construction, this vertex is adjacent in $H$ to all the other vertices of $H$, hence the $(c,\gamma)$-color of $r$ is unique in $H$. ∎
We now show how to obtain $p$-centered colorings from the generalized coloring numbers, proving that nowhere dense and bounded expansion classes admit $p$-centered colorings with a small number of colors.
Let $G$ be a graph, let $r$ be a positive integer and let $\pi$ be an order of $V(G)$. We define $G\left\langle\pi,r\right\rangle$ as the graph with vertex set $V(G)$ and for $u\neq v$ we have $uv\in E(G\left\langle\pi,r\right\rangle)\Leftrightarrow u\in\operatorname{WReach}_{r}[G,\pi,v]$ or $v\in\operatorname{WReach}_{r}[G,\pi,u]$.
Definition 6.5.
Let $G$ be a graph and let $r$ be a positive integer. We define $\operatorname{wcol}_{r}^{\star}(G)$ as the chromatic number of $G\left\langle\pi,r\right\rangle$, where $\pi$ is minimum over all linear orders of $V(G)$.
Of course, $\operatorname{wcol}_{r}^{\star}(G)\leq\operatorname{wcol}_{r}(G)$ for all graphs $G$ and all integers $r$, since we can just use the order $\pi$ witnessing that $\operatorname{wcol}_{r}(G)$ is small and the greedy coloring along the order. In some cases $\operatorname{wcol}_{r}^{\star}(G)$ may be surprisingly small. For example, $\operatorname{wcol}_{2}^{*}(G)$ is bounded by $3\nabla_{0}(G)^{2}+1+\min\{\nabla_{0}(G)\nabla_{1}(G),\nabla_{0}(G)^{2}\nabla_{1}(G)\}$, as shown Esperet and Wiechert [61].
We now turn to Zhu’s result [158].
Theorem 6.6 ([158]).
Let $G$ be a graph and $p$ a positive integer. Then $\chi_{p}(G)\leq\operatorname{wcol}_{2^{p-2}}^{\star}(G)$.
Proof.
Let $\pi$ be an order of $V(G)$ and $c$ be a coloring witnessing that $G\left\langle\pi,2^{p-2}\right\rangle$ has chromatic number $d$. Let $H$ be a connected subgraph of $G$ and let $v_{0}$ be the minimum vertex of $H$ with respect to $\pi$. If $c(v_{0})$ appears exactly once in $H$ we are done. Otherwise, we prove that $H$ gets at least $p$ colors.
As $c(v_{0})$ is not unique in $H$, there exists $u\in V(H)$, $u\neq v_{0}$, with $c(u)=c(v_{0})$. As $H$ is connected there exists a path $P_{0}=v_{0}v_{1}\ldots v_{\ell}=u$ in $H$ connecting $v_{0}$ and $u$. Note that $P_{0}$ must have length greater than $2^{p-2}$, as otherwise $v_{0}$ is weakly $2^{p-2}$-reachable from $u$ ($v_{0}$ is minimum in $H$ and the path uses only vertices of $H$), and hence would receive a color different from $u$.
Now $P_{1}=v_{1}\ldots v_{2^{p-2}}$ does not use color $c(v_{0})$ and has $2^{p-2}$ vertices. As $P_{1}$ has only $2^{p-2}$ vertices, every vertex weakly reaches the minimum vertex $u_{1}$ of $P_{1}$. Hence, $u_{1}$ gets a unique color. We repeat the process with the larger component $P_{2}$ of $P_{1}-u_{1}$, which has at least $2^{p-3}$ vertices. Hence, after $p-2$ steps we have found the vertices $u_{1},\ldots,u_{p-2}$ with pairwise different colors and we are left with $P_{p-2}$, which consists of a single vertex $u_{p-1}$ (with color distinct from $v_{0},u_{1},\dots,u_{p-1}$). Together with $v_{0}$ we found that $H$ gets at least $p$ colors. ∎
Corollary 6.7.
  1. 1.
    A class $\mathscr{C}$ of graphs has bounded expansion if and only if for every $p$ there exists $c(p)$ such that every $G\in\mathscr{C}$ satisfies $\chi_{p}(G)\leq c(p)$.
  2. 2.
    A class $\mathscr{C}$ of graphs is nowhere dense if and only if for every $p$ and every $\varepsilon>0$ there exists $c(p,\varepsilon)$ such that every $n$-vertex graph $H\subseteq G\in\mathscr{C}$ satisfies $\chi_{p}(H)\leq c(p,\varepsilon)\cdot n^{\varepsilon}$.
Unfortunately, the general bounds for the $p$-centered colorings depend on $\operatorname{wcol}_{2^{p-2}}^{\star}(G)$. By the large radius $2^{p-2}$, this leads to bounds that are exponential in $p$ even for very restricted graph classes. Direct constructions improve these bounds to polynomial bounds on many important graph classes. For example, we have polynomial bounds on classes that exclude a minor by the result of Pilipczuk and Siebertz [127], and, as mentioned before, this result not only improved the bounds for $p$-centered colorings, but it was the inspiration for the planar product structure theory [47]. The bounds for planar graphs were improved to $\mathcal{O}(p^{3}\cdot\log p)$ by Dębski et al. [31]. The proof is an easy corollary of a coloring procedure for strong products and the product structure theorem for planar graphs. Quite surprisingly, as shown by Dębski et al., also bounded degree graphs have polynomial $p$-centered colorings, and by the tree decomposition method, this result lifts to classes with excluded topological minors [31]. The beautiful proof for bounded degree graphs builds on the method of entropy compression. Furthermore, Dębski et al. [31] and Dubois et al. [44] provide several lower bounds.
Theorem 6.8 ([44, 31, 127]).
  1. 1.
    Graphs of treewidth at most $k$ admit $p$-centered colorings with $\binom{p+k}{k}$ colors [127] and this bound is tight [31].
  2. 2.
    If $\mathscr{C}$ is a class of graphs that is tree-decomposable with adhesion at most $k$ over a class $\mathscr{D}$ that admits $p$-centered colorings with $\mathcal{O}(p^{d})$ colors, then $\mathscr{C}$ admits $p$-centered colorings with $\mathcal{O}(p^{d+k})$ colors [127].
  3. 3.
    Planar graphs admit $p$-centered colorings with $\mathcal{O}(p^{3}\log p)$ colors and there is a family of planar graphs of treewidth $3$ that requires $\Omega(p^{2}\log p)$ colors in any $p$-centered coloring.
  4. 4.
    Graphs with Euler genus $g$ admit $p$-centered colorings with $\mathcal{O}(gp+p^{3}\log p)$ colors [31].
  5. 5.
    For every graph $H$ there is a polynomial $f$ such that the graphs excluding $H$ as a minor admit $p$-centered colorings with at most $f(p)$ colors [127].
  6. 6.
    Graphs with maximum degree $\Delta$ admit $p$-centered colorings with $\mathcal{O}(\Delta^{2-1/p}\cdot p)$ colors [31] and there are graphs of maximum degree $\Delta$ that require $\Omega(\Delta^{2-1/p}\cdot p\cdot\log^{-1/p}\Delta)$ colors in any $p$-centered coloring [44].
  7. 7.
    For every graph $H$ there is a polynomial $f$ such that the graphs excluding $H$ as a topological minor admit $p$-centered colorings with at most $f(p)$ colors [31].
The following question was posed by Hodor et al. [83].
Problem 6.9.
Does there exist a function $g$ such that for every fixed graph $H$, for every $H$-minor-free graph $G$ and every positive integer $p$, $G$ admits a $p$-centered coloring with $\mathcal{O}(r^{g(\operatorname{td}(H))})$ colors?
6.2 Structural decompositions
As discussed above, $p$-treedepth colorings can be seen as structural decompositions. The key property they provide is that every subgraph with at most $p$ vertices can be found as a subgraph of bounded treedepth. When $p$-treedepth colorings with few colors can be efficiently computed, this, for example, immediately leads to efficient algorithms for the subgraph isomorphism problem or the model checking problem for existential FO. These observations led to the study of colorings such that any $p$ color classes induce graphs with other good structural properties. Let us discuss this more general approach, which in particular opened the road to the study of dense but structurally sparse classes of graphs.
A class property $\Pi$ is a set of graph classes that is closed by subclasses (that is, if $\mathscr{C}\in\Pi$ and $\mathscr{D}\subseteq\mathscr{C}$, then $\mathscr{D}\in\Pi$). For a function $f:\mathbb{N}\rightarrow\mathbb{N}$ and a parameter $p$, a class $\mathscr{C}$ admits an $f$-bounded $\Pi$-decomposition with parameter $p$ if there exists a graph class $\mathscr{D}_{p}\in\Pi$, such that every $G\in\mathscr{C}$ has a vertex coloring with at most $f(|G|)$ colors with the property that every $p$ color classes induce a subgraph in $\mathscr{D}_{p}$. A class $\mathscr{C}$ has bounded-size decompositions (resp. quasibounded-size decompositions) if, for every positive integer $p$, it has an $f_{p}$-bounded $\Pi$-decomposition with parameter $p$, where $f_{p}(n)=\mathcal{O}(1)$ (resp. $f_{p}(n)=\mathcal{O}(n^{\varepsilon})$ for every $\varepsilon>0$). Hence, a class has bounded expansion if and only if it has bounded-size bounded treedepth decompositions and it is nowhere dense if and only if its hereditary closure has quasibounded-size bounded treedepth decompositions.
The study of dense but structurally sparse classes by structural decompositions was initiated by Kwon et al. [101], who proved that fixed powers of bounded expansion classes have bounded-size bounded cliquewidth decompositions. This result was strengthened and extended by Gajarský et al. [66] showing that even all classes that are transductions of bounded expansion classes have bounded-size shrubdepth decompositions. Such classes are called classes with structurally bounded expansion. We also define dependent, monadically dependent, stable, and monadically stable classes, which turn out to be important generalizations of nowhere dense classes.
  • 1.
    A class $\mathscr{D}$ has structurally bounded expansion if there exists a class $\mathscr{C}$ of bounded expansion such that $\mathscr{D}$ can be $\mathrm{FO}$-transduced from $\mathscr{C}$ [66].
  • 2.
    A class $\mathscr{D}$ is structurally nowhere dense if there exists a nowhere dense class $\mathscr{C}$ such that $\mathscr{D}$ can be FO-transduced from $\mathscr{C}$ [66].
  • 3.
    A class $\mathscr{C}$ is stable if from $\mathscr{C}$ one cannot interpret (with an FO-interpretation in powers) the class of all linear orders.
  • 4.
    A class $\mathscr{C}$ is monadically stable if every monadic lift of $\mathscr{C}$ is stable. This is equivalent to saying that from $\mathscr{C}$ one cannot FO-transduce the class of all linear orders (follows from work of Baldwin and Shelah [8]).
  • 5.
    A class $\mathscr{C}$ is dependent if from $\mathscr{C}$ one cannot interpret (with an FO-interpretation in powers) the class of all graphs.
  • 6.
    A class $\mathscr{C}$ is monadically dependent if every monadic lift of $\mathscr{C}$ is dependent. This is equivalent to saying that from $\mathscr{C}$ one cannot FO-transduce the class of all graphs (follows from work of Baldwin and Shelah [8]).
Also the recently introduced and highly influential notion of twinwidth, introduced by Bonnet et al. [18], can be defined via transductions, as shown by Bonnet et al. in [17]
  • 1.
    A class $\mathscr{C}$ has bounded twinwidth if and only if every $G\in\mathscr{C}$ can be linearly ordered such that the resulting class $\mathscr{D}$ of ordered graphs is monadically dependent.
Every class with structurally bounded expansion is structurally nowhere dense, which in turn is monadically stable, and which in turn is monadically dependent. Classes with bounded shrubdepth have bounded cliquewidth, which in turn are monadically dependent. Recall that a class of graphs is monotone if it is closed under taking subgraphs and hereditary if it is closed under taking induced subgraphs. On hereditary classes the notions of stability and monadic stability, as well as the notions of dependence and monadic dependence coincide [19]. On monotone classes the notions of stability, monadic, stability, dependence, monadic dependence and nowhere denseness coincide [3]. From this we easily get the following nice observation.
Observation 6.10.
  1. 1.
    A hereditary class $\mathscr{C}$ is monadically stable (dependent) if and only if the class of its incidence graphs is stable (dependent).
  2. 2.
    A monotone class $\mathscr{C}$ is nowhere dense if and only if the class of its incidence graphs is monadically stable.
Proof.
Let $\mathscr{C}$ be hereditary. Then $\mathscr{C}$ is monadically stable (dependent) if and only if it is stable (dependent) by the result of [19]. As every interpretation of a stable (dependent) class is stable (dependent), the statement follows from the fact that the class of incidence graphs of $\mathscr{C}$ is bi-interpretable with $\mathscr{C}$.
Let $\mathscr{C}$ be monotone. If $\mathscr{C}$ is nowhere dense, then it is stable by the result of [3]. Since the class of incidence graphs of $\mathscr{C}$ is an interpretation of $\mathscr{C}$ and every interpretation of a stable class is stable, it follows that the class of incidence graphs $\mathscr{D}$ of $\mathscr{C}$ is stable. As $\mathscr{C}$ is monotone it follows that $\mathscr{D}$ is hereditary [^3] Formally we also have to deal with incidence vertices that are connected to only one graph vertex, but it is easy to see that monadic stability is not affected by these vertices.. It follows that $\mathscr{D}$ is monadically stable by the result of [19]. Vice versa, if a class $\mathscr{D}$ of incidence graphs is monadically stable it follows that its hereditary closure $\mathscr{D}^{\prime}$ is hereditary and also monadically stable. Then the class $\mathscr{C}$ of graphs encoded in $\mathscr{D}^{\prime}$ is monotone and monadically stable, which implies that it is nowhere dense by the result of [3]. ∎
Structurally bounded expansion classes admit bounded-size bounded shrubdepth decompositions as shown by Gajarský et al. [66], structurally nowhere dense classes admit quasibounded-size bounded shrubdepth decompositions as shown by Dreier et al. [38] and in fact, a hereditary class admits quasibounded-size bounded shrubdepth decompositions if and only if it is stable as shown by Braunfeld et al. [21]. One interesting open problems in this direction is whether hereditary dependent classes admit quasibounded-size bounded twinwidth decompositions. We refer to [16, 20, 37, 38, 41, 42, 65, 66, 67, 68, 121, 118, 124, 147] for more background on the recent exciting development of a theory of structural sparsity.
Problem 6.11.
Does every monadically dependent class admit quasibounded-size bounded twinwidth decompositions?
6.3 Quantifier elimination revisited
According to Theorem 2.22 graphs with bounded treedepth admit quantifier elimination in the following sense. For every FO-formula $\varphi(\bar{x})$ and every class $\mathscr{C}$ of colored graphs with bounded treedepth there exists a quantifier-free formula $\tilde{\varphi}(\bar{x})$ and a linear time computable map $Y$ such that, for every $G\in\mathscr{C}$, $Y(G)$ is a guided expansion of $G$ such that for all tuples of vertices $\bar{v}$
$G\models\varphi(\bar{v})\quad\iff\quad Y(G)\models\tilde{\varphi}(\bar{v}).$
This quantifier elimination result easily lifts through bounded-size bounded treedepth decompositions. A similar statement was first proved by Dvořák et al. [53] and Grohe and Kreutzer [75]. Our proof sketch follows Gajarský et al. [66].
Theorem 6.12 ([53, 66, 75]).
For every FO-formula $\varphi(\bar{x})$ and every class $\mathscr{C}$ of guided pointer structures with bounded expansion there exists a quantifier-free formula $\tilde{\varphi}(\bar{x})$ and a linear time computable map $Y$ such that, for every $G\in\mathscr{C}$, $Y(G)$ is a guided expansion of $G$ such that for all tuples of vertices $\bar{v}$
$G\models\varphi(\bar{v})\quad\iff\quad Y(G)\models\tilde{\varphi}(\bar{v}).$
Proof sketch..
We show the statement by induction on the structure of $\varphi$. In the atomic case there is nothing to prove, and in the inductive step, the interesting case is the case of existential quantification. In this case we show how an existential quantifier can be eliminated at the cost of introducing guided unary functions and colors. Consider a formula
$\varphi(\bar{x})=\exists y(\psi(\bar{x},y)),$
where $\psi(\bar{x},y)$ is a quantifier-free formula using unary functions (obtained by induction).
As $\mathscr{C}$ has bounded expansion, it admits a bounded-size bounded treedepth decomposition with parameter $p$, where $p$ is the number of different terms occurring in $\psi(\bar{x},y)$. For any graph $G\in\mathscr{C}$ we consider the family $\mathcal{U}_{G}=\{U_{I}\mid I$ set of $p$ colors}}\}}, where $U_{I}$ is the set of vertices induced by the colors from $I$. By assumption, each $G[U_{I}]$ has bounded treedepth and there is a constant $N$ such that $|\mathcal{U}_{G}|\leq N$ for all $G\in\mathscr{C}$.
For each $G\in\mathscr{C}$, the quantified variable $y$ must be in one of the sets $U_{I}\in\cal U_{G}$. Therefore, we can rewrite $\varphi(\bar{x})$ as
$\bigvee_{I}\exists y(y\in U_{I}\wedge\psi(\bar{x},y)).$
Here, we make the colors available in the structure by adding them as unary predicates $P_{1},\ldots,P_{t}$ and write $y\in U_{I}$ as an abbreviation for $\bigvee_{i\in I}P_{i}(y)$. Write $\psi_{I}$ for $\exists y(x\in U_{I}\wedge\psi(\bar{x},y)$. By Theorem 2.22 there exists a there exists a quantifier-free formula $\tilde{\psi}_{I}(\bar{x})$ and a linear time computable map $Y_{I}$ such that, for every $G\in\mathscr{C}$, $Y_{I}(G[U_{I}])$ is a guided expansion of $G[U_{I}]$ such that $\psi_{I}$ is equivalent to $\tilde{\psi}_{I}$ on $G[U_{I}]$. By expanding $G$ with the colors and all the $Y_{I}$ and restricting quantification appropriately, we obtain the guided expansion $Y$ of $G$ such that
$\tilde{\varphi}(\bar{x})=\bigvee_{I}\tilde{\psi}_{I}(\bar{x})$
is equivalent to $\varphi$ on $Y(G)$. ∎
Theorem 6.12 extends to classes with structurally bounded expansion [66], as well as to first-order logic with modulo counting on classes with bounded expansion [119].
7 Game characterizations and wideness
In this section we are going to study game characterizations of the generalized coloring numbers as well as connections to a related notion called wideness. Just as treewidth can be characterized by cops-and-robber games in a very intuitive way, it turns out that the generalized coloring numbers can be characterized by games. Cops-and-robber games are played on a graph by a team of cops and a robber. The robber resides on the vertices of the graph. The cops aim to catch the robber who is trying to avoid capture. Depending on the concrete setting, e.g. visibility, speed of cops and robber, and local restrictions in the game rounds, we obtain different models that correspond to different width measures.
7.1 Cops-and-robber games
Let us first recall the classical cops-and-robber game introduced by Seymour and Thomas [141] characterizing treewidth. The game is played by two players on a graph $G$. The first player controls a robber that stands on a vertex of $G$ and can run at infinite speed along paths in the graph. The second player controls a team of $k$ cops that also occupy vertices of $G$. The goal of the cops is to catch the robber. The game proceeds in rounds. In each round, some of the cops may step into helicopters (these cops are temporarily removed from the game) and announce the next position where they want to be placed. Now the robber can run to its new position, however, it is not allowed to run through a path that contains a vertex occupied by a cop. The cops win if they can land on the position of the robber, that is, if it has no moves left when the cops announce to land on its position. Otherwise, the game continues forever and the robber wins. Seymour and Thomas proved that the cops-and-robber game can be used to characterize treewidth.
Theorem 7.1 (Seymour and Thomas [141]).
The least number of cops needed to catch the robber on $G$ is equal to the treewidth of $G$ plus one.
In the above game, the robber is visible, agile, and can run at infinite speed. In a different variant introduced by Dendris et al. [32] the robber is invisible and lazy, in the sense that it may move only when one of the cops plans to occupy the vertex it currently resides on. This variant turns out to be equivalent to the visible-and-agile variant of the game [32]. Dendris et al. also considered a bounded speed variant of the invisible-and-inert game. A robber with speed $r$ in its move is allowed to traverse a cop-free path of length at most $r$. They observed that for speed $1$ the game characterizes the degeneracy of the graph and they called the number of cops needed to capture a robber with speed $r$ the $r$-dimension of the graph, which corresponds exactly to the strong $r$-coloring number $\operatorname{col}_{r}(G)$. To the best of our knowledge (and to our greatest regret) they did not follow up on studying the $r$-dimension.
Theorem 7.2 ([32]).
The least number of cops needed to catch the robber in the invisible-and-inert game with speed $1$ on $G$ is equal to the degeneracy of $G$ plus $1$.
The least number of cops needed to catch the robber in the invisible-and-inert game with speed $r$ on $G$ is equal to the $r$-dimension of $G$, which is equal to $\operatorname{col}_{r}(G)$.
Surprisingly, to the best of our knowledge, the bounded speed variant of the visible-and-agile game was only very recently studied by Toruńczyk [147]. It does not seem to correspond one-to-one to an established width measure but Toruńczyk proved that the number of cops required to catch the robber in the speed-$r$ variant of the visible-and-agile game is sandwiched between the $r$-admissibility and the weak $2r$-coloring number.
Theorem 7.3 ([147]).
The least number of cops needed to catch the robber in the visible-and-agile game with speed $1$ on $G$ is equal to the degeneracy of $G$ plus $1$.
The least number of cops needed to catch the robber in the visible-and-agile game with speed $r$ on $G$ is bounded from below by $\operatorname{adm}_{r}(G)+1$ and bounded from above by $\operatorname{wcol}_{2r}(G)$.
We sketch the proof of the second statement.
Proof.
The lower bound follows easily from the characterization of $r$-admissibility by $(k,r)$-fan sets, which provide a shelter for the robber (Theorem 3.14).
Let us prove the upper bound. Let $\pi$ be a linear order witnessing that $\operatorname{wcol}_{2r}(G)=:c$. The following is a winning strategy for the cops: in each round, if the robber is placed at a vertex $v$, then the cops fly to the vertices $w\leq v$ that are weakly-$2r$ reachable from $v$.
We show that this is a winning strategy for the cops. Let $P_{1},P_{2},\ldots$ be the paths in $G$ such that $P_{i}$ is the path of length at most $r$ along which the robber moved from its $i$-th position $v_{i}$ to its $(i+1)$-st position $v_{i+1}$. Denote by $m_{i}$ the smallest vertex of the path $P_{i}$ with respect to $\pi$. We claim that $m_{i+1}>_{\pi}m_{i}$ for all $i$. Otherwise, $m_{i+1}$ is weakly-$2r$ reachable from $v_{i}$ (as witnessed by the concatenated path $P_{i}P_{i+1}$ cut at $m_{i+1}$).
Hence, when the robber traversed the path $P_{i+1}$ the vertex $m_{i+1}$ was occupied, which is not possible. Therefore, we have $m_{1}<_{\pi}m_{2}<_{\pi}\ldots$, so the cops win after at most $n$ rounds. ∎
Corollary 7.4.
A graph class $\mathscr{C}$ has bounded expansion if and only if for every $r\in\mathbb{N}$ there exists a constant $c(r)$ such that for every $G\in\mathscr{C}$ the least number of cops needed to catch the robber in the invisible-and-inert game with speed $r$ or in the visible-and-agile game with speed $r$ is bounded by $c(r)$.
A class $\mathscr{C}$ is nowhere dense if and only if for every $r\in\mathbb{N}$ and $\varepsilon>0$ there exists a constant $c(r,\varepsilon)$ such that for every $H\subseteq G\in\mathscr{C}$ the least number of cops needed to catch the robber in the invisible-and-inert game with speed $r$ or in the visible-and-agile game with speed $r$ on $H$ is bounded by $c(r,\varepsilon)\cdot|H|^{\varepsilon}$.
The lift-free game, introduced by Ganian et al. [70], is played like the visible-and-agile game but the cop player is not allowed to move a cop from a vertex once it has landed. That is, in the lift-free game the cops are placed on the vertices of the graph one after another and we count the number of cops needed to catch the robber.
Theorem 7.5 ([70]).
The least number of cops needed to catch the robber in the lift-free game on $G$ is equal to the treedepth of $G$.
In order to appropriately localize this game we take a different view on counting the number of rounds in the lift-free game. For $r\in\mathbb{N}\cup\{\infty\}$, in the radius-$r$ counter game every vertex is equipped with a counter that is initially set to 00. In the radius-$r$-counter game the cops are placed on the vertices of the graph one after the other until all vertices are occupied by a cop. Denote the set of placed cops in round $i$ by $X_{i}$ (and $X_{0}=\emptyset$. In round $i$ when a cop is placed on vertex $v_{i}$, we increase the counters of all vertices in the $r$-neighborhood of $v_{i}$ in $G-X_{i-1}$ by $1$ (also the counter of the vertex itself). The radius-$r$ counter-depth of a play is the largest number that a counter takes in the play and the radius-$r$ counter-depth of $G$ is the smallest counter-depth that the cops can ensure in the radius-$r$-counter game with optimal play.
Theorem 7.6.
The radius-$r$ counter-depth of $G$ is equal to $\operatorname{wcol}_{r}(G)$.
In particular, the radius-$1$ counter-depth of $G$ is equal to the degeneracy of $G$ plus $1$ and the radius-$\infty$ counter-depth of $G$ is equal to the treedepth of $G$.
Proof.
Let $G$ be a graph and let $\pi=v_{1},\ldots,v_{n}$ be an order of $V(G)$. Then $v_{i}\in\operatorname{WReach}_{r}[G,\pi,v_{j}]$ if and only if there is a path $P$ of length at most $r$ between $v_{j}$ and $v_{i}$ such that $v_{j}$ is the minimum vertex on $P$, which is exactly the set of vertices that increase the counter of $v_{j}$ when the cops occupy the vertices in order $v_{1},\ldots,v_{n}$.
Corollary 7.7.
A graph class $\mathscr{C}$ has bounded expansion if and only if for every positive integer $r$ there exists a constant $c(r)$ such that for every $G\in\mathscr{C}$ the radius-$r$ counter depth is bounded by $c(r)$.
A class $\mathscr{C}$ is nowhere dense if and only if for every positive integer $r$ and every $\varepsilon>0$ there exists a constant $c(r,\varepsilon)$ such that for every $H\subseteq G\in\mathscr{C}$ the radius-$r$ counter depth of $H$ is bounded by $c(r,\varepsilon)\cdot|H|^{\varepsilon}$.
7.2 Splitter Game and Quasi-Wideness
Another way of localizing lift-free games is to restrict the game graph in each round to a local neighborhood of the last cop move. This leads to the Splitter game, which characterizes nowhere dense classes, as shown by Grohe et al. [77]. We present an equivalent characterization in terms of uniform quasi-wideness and derive bounds for the length of the Splitter game as well as for uniform quasi-wideness in terms of the generalized coloring numbers.
Let $G$ be a graph. The radius-$r$ Splitter game on $G$ is played by two players: Connector and Splitter, as follows. The game starts on $G_{0}:=G$. In round $i+1$ of the game, Connector chooses a vertex $c_{i+1}\in V(G_{i})$. Then Splitter picks a vertex $s_{i+1}\in N_{r}^{G_{i}}(c_{i+1})$ and deletes it, that is, the game continues on $G_{i+1}:=G_{i}[N_{r}^{G_{i}}(c_{i+1})\setminus\{s_{i+1}\}]$. Splitter wins if $G_{i+1}=\emptyset$, otherwise, the game continues.
Observation 7.8.
Let $G$ be a graph. Then $\operatorname{td}(G)$ is equal to the least number such that Splitter wins the radius-$\infty$ Splitter game.
As shown in [77] the Splitter game characterizes nowhere dense classes.
Theorem 7.9 ([77]).
A class $\mathscr{C}$ of graphs is nowhere dense if and only if for every positive interger $r$ there exists a constant $\ell(r)$ such that Splitter wins the radius-$r$ Splitter game on every $G\in\mathscr{C}$ in at most $\ell(r)$ rounds.
The proof is based on the fact that nowhere dense classes are uniformly quasi-wide (defined below). It was observed by Kreutzer et al. [96] that the weak coloring number bounds the number of rounds that Splitter needs to win.
Lemma 7.10 ([96]).
Let $G$ be a graph. Then Splitter has a strategy to win the radius-$r$ Splitter game on $G$ in $\ell=\operatorname{wcol}_{2r}(G)+1$ rounds.
Proof.
Given a graph $G$ and an integer $r$. Let $\ell:=\operatorname{wcol}_{2r}(G)$ and let $\pi$ be a witnessing order. Splitter plays the following strategy: in any round $i+1$, Splitter picks $s_{i+1}=\min_{\pi}(N_{r}^{G_{i}}(c_{i+1}))$. Then for every remaining vertex $v\in V(G_{i+1})=N_{r}^{G_{i}}(c_{i+1})\setminus\{s_{i+1}\}$ we have $s_{i+1}\in\operatorname{WReach}_{2r}[G,\pi,v])$, as there is a path of length at most $2r$ connecting $v$ and $s_{i+1}$ (via the central vertex $c_{i+1}$) and $s_{i+1}$ is minimum on this path (compare this to the counter game). We conclude that the game must end after at most $\ell$ rounds. ∎
The fact that nowhere dense classes are characterized by the Splitter game is derived from a characterization of these classes by the notion of uniform quasi-wideness, which was introduced by Dawar [29]. A set $B\subseteq V(G)$ is called distance-$r$ independent if the vertices of $B$ pairwise have distance greater than $r$.
Definition 7.11.
A class $\mathscr{C}$ of graphs is uniformly quasi-wide if for every $m\in\mathbb{N}$ and every $r\in\mathbb{N}$ there exist numbers $N(m,r)$ and $s(r)$ such that the following holds.
Let $G\in\mathscr{C}$ and let $A\subseteq V(G)$ with $|A|\geq N(m,r)$. Then there exists a set $S\subseteq V(G)$ with $|S|\leq s(r)$ and a set $B\subseteq A\setminus S$ of size at least $m$ that is distance-$r$ independent in $G-S$.
Figure 16: Uniform quasi-wideness (case $r=2$). By deleting a subset $S$ of $s=3$ vertices (in gray on the right), the set $A$ of vertices on the circle will include a subset $B$ of $m=4$ vertices (in red on the right) that is $2$-independent.
Nešetřil and Ossona de Mendez [115] proved that the notions of nowhere denseness and uniform quasi-wideness coincide.
Theorem 7.12 ([115]).
A class $\mathscr{C}$ of graphs is nowhere dense if and only if it is uniformly quasi-wide.
Theorem 7.12 was proved in model theory by Podewski and Ziegler [133] already in 1978. Podewski and Ziegler introduced the notion of superflatness, which is equivalent to nowhere denseness on classes of finite graphs. This result was brought to the attention of the graph theory community by Adler and Adler [3] and inspired the recent investigation of connections between model-theoretic notions and structural sparsity.
It was shown by Kreutzer et al. [97] (improved and constructive bounds were provided by Pilipczuk et al. [130]) that the bounds for the function $N(m,r)$ can be made polynomial, which is important for many algorithmic applications. While the proofs for nowhere dense classes is quite technical, a quite simple proof that bounded expansion classes are uniformly quasi-wide is given by Nadara et al. [108].
Theorem 7.13 ([108]).
Assume we are given a graph $G$, a set $A\subseteq V(G)$, integers $r\geq 1$ and $m\geq 2$, and an order $\pi$ of $V(G)$ with $c=\operatorname{wcol}_{r}(G,\pi)$. Furthermore, assume that $|A|\geq 4\cdot(2cm)^{c+1}$. Then in polynomial time, one can compute sets $S\subseteq V(G)$ and $B\subseteq A\setminus S$ such that $|S|\leq c$, $|B|\geq m$, and $B$ is distance-$r$ independent in $G-S$.
Let us mention the following application of uniform quasi-wideness due to Nešetřil and Ossona de Mendez [117], showing that nowhere dense classes are locally well separable.
Definition 7.14.
A class $\mathscr{C}$ of graphs admits balanced neighborhood separators if for every non-negative integer $r$ and every real $\varepsilon>0$ there exists a number $s=s(r,\varepsilon)$ such that the following holds. For every graph $G\in\mathscr{C}$ and every subset $A\subseteq V(G)$ there exists a set $S\subseteq V(G)$ of size at most $s$ such that $|N_{r}^{G-S}(v)\cap A|\leq\varepsilon|A|$ for all $v\in V(G)\setminus S$.
Theorem 7.15 ([117]).
A class $\mathscr{C}$ of graphs is nowhere dense if and only if $\mathscr{C}$ admits balanced neighborhood separators.
7.3 Flipwidth, flipper-game, and flip-flatness
We conclude this section with a short overview of how the presented games and the notion of uniform quasi-wideness have recently been generalized to dense graphs with the hinted at beautiful connections to model theory. We begin with the flipwidth game, which was recently introduced by Toruńczyk [147]. The idea of game is to (temporarily) flip adjacencies between arbitrarily large sets of vertices (instead of occupying them by cops) – an operation that is well suited to simplify dense graphs. Note also the close connection between flips and quantifier-free transductions.
Definition 7.16.
Let $G$ be a graph and $X,Y\subseteq V(G)$. A flip between $X$ and $Y$ results in the graph $G^{\prime}$ obtained from $G$ by inverting the adjacency of every pair $(x,y)$ of vertices with $x\in X$ and $y\in Y$: every pair $(x,y)$ that is adjacent in $G$ becomes non-adjacent in $G^{\prime}$, and vice-versa. A $k$-flip of a graph $G$ is obtained by partitioning $V(G)$ into $k$ parts, and then performing flips between some pairs $X,Y$ of parts of the partition (possibly with $X=Y$).
The flipwidth game is defined as follows.
Definition 7.17.
The flipwidth game with radius $r$ and width $k$ on a graph $G$ is a game played by two players, Flipper and Runner, as follows. We let $G_{0}:=G$ and let Runner pick any vertex $v_{0}$ of $G$. In the $i$th round, Flipper announces a $k$-flip $G_{i}$ of the original graph $G$. Runner can move from their previous position $v_{i-1}$ to a new position $v_{i}$, by traversing a path of length at most $r$ in $G_{i-1}$. The game is won by Flipper if the runner’s new position $v_{i}$ is isolated in the announced $k$-flip $G_{i}$.
The flipwidth of radius $r$ of a graph $G$, denoted $\operatorname{fw}_{r}(G)$, is the smallest number $k$ for which Flipper has a winning strategy in the flipwidth game with radius $r$ and width $k$.
A class $\mathscr{C}$ of graphs has bounded flipwidth if $\operatorname{fw}_{r}(\mathscr{C})<\infty$ for all $r\in\mathbb{N}$. It has almost bounded flipwidth if for every $\varepsilon>0$ and $r\in\mathbb{N}$, we have $\operatorname{fw}_{r}(H)\in\mathcal{O}_{r,\varepsilon}(|H|^{\varepsilon})$ for every $n$-vertex graph $H$ which is an induced subgraph of a graph $G\in\mathscr{C}$.
Toruńczyk proved the following connections with graph theory and model theory. First, he proved that the flipwidth game is a robust generalization of the bounded-speed variants of the cops-and-robber game by considering weakly sparse graph classes. A graph class is weakly sparse if it excludes some fixed biclique $K_{t,t}$ (balanced complete bipartite graph with parts of size $t$) as a subgraph.
Theorem 7.18 ([147]).
Let $\mathscr{C}$ be a class of graphs. Then:
  1. 1.
    $\mathscr{C}$ has bounded degeneracy if and only if $\mathscr{C}$ is weakly sparse and $\operatorname{fw}_{1}(\mathscr{C})<\infty$,
  2. 2.
    $\mathscr{C}$ has bounded treewidth if and only if $\mathscr{C}$ is weakly sparse and $\operatorname{fw}_{\infty}(\mathscr{C})<\infty$,
  3. 3.
    $\mathscr{C}$ has bounded expansion if and only if $\mathscr{C}$ is weakly sparse and $\mathscr{C}$ has bounded flipwidth.
  4. 4.
    $\mathscr{C}$ is nowhere dense if and only if $\mathscr{C}$ is weakly sparse and $\mathscr{C}$ has almost bounded flipwidth.
On dense graphs he obtained the following wonderful results.
Theorem 7.19.
Let $\mathscr{C}$ be a class of graphs. Then
  1. 1.
    $\mathscr{C}$ has bounded cliquewidth if and only if $\operatorname{fw}_{\infty}(\mathscr{C})<\infty$.
  2. 2.
    if $\mathscr{C}$ has bounded twinwidth, then $\mathscr{C}$ has bounded flipwidth. Furthermore, if $\mathscr{C}$ is a class of ordered graphs then $\mathscr{C}$ has bounded twinwidth if and only if $\mathscr{C}$ has bounded flipwidth.
  3. 3.
    if $\mathscr{C}$ is structurally nowhere dense, then $\mathscr{C}$ has almost bounded flipwidth.
  4. 4.
    if $\mathscr{C}$ has bounded flipwidth and $\mathsf{T}$ is a first-order transduction, then $\mathsf{T}(\mathscr{C})$ has bounded flipwidth. Hence, every class of bounded flipwidth is monadically dependent.
Toruńczyk conjectures that a class has almost bounded flipwidth if and only if it is monadically dependent.
We now come to a second game, the flipper game, which was introduced by Gajarský et al. and which can be seen as a generalization of the Splitter game, and which characterizes monadically stable classes [67].
Definition 7.20.
The radius-$r$ flipper game is played by two players, Flipper and Connector, on a graph $G$ as follows. The game starts at $G_{0}:=G$. In round $i+1$ of the game, Connector chooses a vertex $v_{i+1}\in V(G_{i})$. Flipper chooses a flip and applies it to $G_{i}[N_{r}^{G_{i}}(v_{i+1})]$ and the game continues on the resulting graph $G_{i+1}$. Flipper wins if $G_{i+1}$ consists of a single vertex.
Theorem 7.21 ([67]).
A class $\mathscr{C}$ of graphs is monadically stable if and only if for every positive integer $r$ there exists a constant $\ell(r)$ such that for every graph $G\in\mathscr{C}$, Flipper can win the radius-$r$ Flipper game on $G$ within $\ell(r)$ rounds.
The theorem is proved by another characterization of monadically stable classes via uniform flip-flatness, which was introduced by Dreier et al. [41]. We remark that flip-flatness is not called flip-wideness (as it should be) in order to avoid confusion with flipwidth.
Definition 7.22.
A class $\mathscr{C}$ of graphs is uniformly flip-flat if for every $r\in\mathbb{N}$ there exists a function $N_{r}\colon\mathbb{N}\rightarrow\mathbb{N}$ and a constant $k_{r}\in\mathbb{N}$ such that for all $m\in\mathbb{N}$, $G\in\mathscr{C}$, and $A\subseteq V(G)$ with $|A|\geq N_{r}(m)$, there exists a $k_{r}$-flip $H$ of $G$ and $B\subseteq A$ with $|B|\geq m$ such that $B$ is distance-$r$ independent in $H$.
Theorem 7.23 ([41]).
A class of graphs is uniformly flip-flat if and only if it is monadically stable.
As observed by Dreier et al. [42], $\infty$-flip-flatness corresponds to bounded shrubdepth (here and below distance $\infty$ means that vertices lie in different connected components).
Theorem 7.24 ([42]).
A class of graphs is uniformaly $\infty$-flip-flat if and only if it has bounded shrubdepth.
Dreier et al. also introduced the new notion of flip-breakability and used it to characterize monadically dependent classes.
Definition 7.25.
A class of graphs $\mathscr{C}$ is uniformly flip-breakable if for every $r\in\mathbb{N}$ there exists a function $N_{r}\colon\mathbb{N}\rightarrow\mathbb{N}$ and a constant $k_{r}\in\mathbb{N}$ such that for all $m\in\mathbb{N}$, $G\in\mathscr{C}$, and $W\subseteq V(G)$ with $|W|\geq N_{r}(m)$, there exists a $k_{r}$-flip $H$ of $G$ and $A,B\subseteq W$ with $|A|,|B|\geq m$ such that in $H$ all vertices of $A$ have distance greater than $r$ from all vertices in $B$.
They also considered the notion of deletion-breakability, which relates to flip-breakability as uniform quasi-wideness relates to uniform flip-flatness.
Definition 7.26.
A class of graphs $\mathscr{C}$ is uniformly deletion-breakable if for every $r\in\mathbb{N}$ there exists a function $N_{r}\colon\mathbb{N}\rightarrow\mathbb{N}$ and a constant $k_{r}\in\mathbb{N}$ such that for all $m\in\mathbb{N}$, $G\in\mathscr{C}$, and $W\subseteq V(G)$ with $|W|\geq N_{r}(m)$, there exists a set $S$ with $|S|\leq k_{r}$ and $A,B\subseteq W\setminus S$ with $|A|,|B|\geq m$ such that all vertices of $A$ have distance greater than $r$ from all vertices in $B$ in $G-S$.
Dreier et al. [42] proved the following beautiful characterization theorems.
Theorem 7.27 (Dreier et al. [42]).
A class of graph is
  1. 1.
    uniformly flip-breakable if and only if it is monadically dependent.
  2. 2.
    uniformly $\infty$-flip-breakable if and only if it has bounded cliquewidth.
  3. 3.
    uniformly deletion-breakable if and only if it is nowhere dense.
  4. 4.
    uniformly $\infty$-deletion-breakable if and only if it has bounded treewidth.
8 Vapnik-Chervonenkis dimension
In this section, we study the classical notions of VC-dimension and order-dimension and their connection to the generalized coloring numbers. VC-dimension was originally introduced by Vapnik and Chervonenkis [152] as a measure of complexity of set systems. VC-dimension plays a key role in learning theory as it characterizes PAC-learnability by the result of Blumer et al. [12]. Independently it was introduced by Shelah in model theory where it defines dependent theories [142]. The connection between PAC-learning and dependence was observed by Laskowski [102].
A second related measure is Littlestone-dimension (we are going to work with order-dimension, which is functionally equivalent), which characterizes online-learning by the result of Littlestone [104]. In model theory it defines stability, as observed by Chase and Freitag [22].
As mentioned before, nowhere dense graph classes are both stable and dependent, hence, set systems definable by first-order formulas in nowhere dense classes have bounded order-dimension and bounded VC-dimension. Besides the direct applications of VC-dimension and order-dimension for definable set systems in nowhere dense classes, these connections have led to further interesting combinatorial and algorithmic applications that we will present in the following. The generalized coloring numbers as well as uniform quasi-wideness have played a big role in the development of the theory in this direction.
8.1 VC-dimension and number of types
The VC-dimension of a set system is defined as follows.
Definition 8.1.
Let $X$ be a set and let $\mathcal{F}\subseteq\mathcal{P}(X)$ be a family of subsets of $X$. A subset $A\subseteq X$ is shattered by $\mathcal{F}$ if $\{A\cap F\colon F\in\mathcal{F}\}=\mathcal{P}(A)$; that is, every subset of $A$ can be obtained as the intersection of some set from $\mathcal{F}$ with $A$. The VC-dimension, of $\mathcal{F}$ is the maximum size of a subset $A\subseteq X$ that is shattered by $\mathcal{F}$.
Definition 8.2.
Let $X$ be a set and let $\mathcal{F}\subseteq\mathcal{P}(X)$ be a family of subsets of $X$. The order-dimension of $\mathcal{F}$ is the largest integer $\ell$ such that there exist elements $x_{1},\ldots,x_{\ell}\in X$ and $F_{1},\ldots,F_{\ell}\in\mathcal{F}$ with $x_{i}\in F_{j}\Leftrightarrow i\leq j$.
The key property of VC-dimension is that it implies polynomial upper bounds on the number of different traces that a set system can have on a given parameter set. This is the statement of the famous Sauer-Shelah Lemma.
Theorem 8.3 (Sauer-Shelah Lemma, [152, 139, 143]).
For any family $\mathcal{F}$ of subsets of a set $X$, if the VC-dimension of $\mathcal{F}$ is $d$, then for every finite $A\subset X$,
|$|\{A\cap F\mid F\in{\cal F}\}|\leq c\cdot|A|^{d},\qquad\textit{where $c$ is a universal constant.}$.
Definition 8.4.
The vc-density (also called the vc-exponent) of a set system $\cal F$ on an infinite set $X$ is the infimum of all reals $\alpha>0$ such that |$|\left\{A\cap F\,\mid\,F\in\cal F\right\}|\in\mathcal{O}(|A|^{\alpha})$, for all finite $A\subset X$.
The vc-density of neighborhood set systems in bounded expansion classes was first studied by Reidl et al. [134, 135] under the name neighborhood complexity.
Theorem 8.5 ([135]).
Let $G$ be a graph and $A\subseteq V(G)$. Then
|$|\{N_{r}(v)\cap A\mid v\in V(G)\}|\leq\frac{1}{2}(2r+2)^{\operatorname{wcol}_{2r}(G)}\cdot\operatorname{wcol}_{2r}(G)\cdot|A|+1.$
The problem with extending this result to nowhere dense classes in a useful way was that the weak coloring number, which can only be bounded by $\mathcal{O}(n^{\varepsilon})$ in nowhere dense classes, appears in the exponent in the above inequality.
As shown by Pilipczuk and Siebertz [128], the VC-dimension of the $r$-neighborhoods set system in nowhere dense classes is very small.
Lemma 8.6 ([128]).
[…]
$\{N_{r}(v)\mid v\in V(G)\}$
of $r$-neighborhoods is at most $t-1$.
This yields already polynomial neighborhood complexity for nowhere dense graph classes. On the way to proving that neighborhood set systems in nowhere dense classes have in fact vc-density $1$, Eickmeyer et al. [56] first proved that also the set system of weak coloring numbers in nowhere dense classes has bounded VC-dimension.
Lemma 8.7 ([56]).
Let $\mathscr{C}$ be a nowhere dense class of graphs and let $G\in\mathscr{C}$. Let $\pi$ be a linear order of $V(G)$ and let
$\mathcal{W}_{r,\pi}=\{\operatorname{WReach}_{r}[G,\pi,v]\mid v\in V(G)\}$
be the set system of weak $r$-reachability sets. Then there exists a constant $vc(r)$, depending only on $\mathscr{C}$ and $r$ (and not on $G$ and $\pi$) such that $\mathcal{W}_{r,\pi}$ has VC-dimension at most $vc(r)$.
Based on this result, the proof of Theorem 8.5 could be lifted to nowhere dense classes.
Theorem 8.8 ([56]).
Let $\mathscr{C}$ be a nowhere dense class of graphs. Then there exists a function $f(r,\varepsilon)$ such that for every $G\in\mathscr{C}$, $A\subseteq V(G)$, $r\geq 1$ and $\varepsilon>0$
|$|\{N_{r}(v)\cap A\mid v\in V(G)\}|\leq f(r,\varepsilon)\cdot|A|^{1+\varepsilon}.$
The setting studied in model theory is to consider definable set systems. One of the key results of Shelah is that stable and dependent theories can be characterized by upper bounds on the number of types (as defined next).
For a first order formula $\varphi(\bar{x},\bar{y})$ with free variables partitioned into object variables $\bar{x}$ and parameter variables $\bar{y}$, a structure $\mathbb{A}$, and a subset of its domain $B\subseteq A$, define the set system of $\varphi$-types with parameters from $B$ that are realized in $\mathbb{A}$ as
$S^{\varphi}(\mathbb{A}/B)=\left\{\big{\{}\bar{v}\ \in B^{|\bar{y}|}\,\colon\,\mathbb{A}\models\varphi(\bar{u},\bar{v})\big{\}}\colon\,\bar{u}\in A)^{|\bar{x}|}\right\}\ \subseteq\ \mathcal{P}(B^{|\bar{y}|}).$
Every tuple $\bar{u}\in A^{|\bar{x}|}$ defines the set of those tuples $\bar{v}\in B^{|\bar{y}|}$ for which $\varphi(\bar{u},\bar{v})$ is true. The set $S^{\varphi}(\mathbb{A}/B)$ consists of all subsets of $B^{|\bar{y}|}$ that can be defined in this way. We can also work with the following definitions of stability and dependence.
Definition 8.9.
A class $\mathscr{C}$ of structures is dependent if for every $\varphi(\bar{x},\bar{y})$ there exists a constant $d$ such that for all $\mathbb{A}\in\mathscr{C}$ and all $B\subseteq A$ the VC-dimension of $S^{\varphi}(\mathbb{A}/B)$ is bounded by $d$.
Definition 8.10.
A class $\mathscr{C}$ of structures is stable if for every $\varphi(\bar{x},\bar{y})$ there exists a constant $d$ such that for all $\mathbb{A}\in\mathscr{C}$ and all $B\subseteq A$ the order-dimension of $S^{\varphi}(\mathbb{A}/B)$ is bounded by $d$.
A class $\mathscr{C}$ is called monadically dependent or monadically stable, respectively, if the class of all colorings of structures from $\mathscr{C}$ is dependent or stable, respectively. As proved by Baldwin and Shelah [8] in this case it suffices to consider formulas $\varphi(x,y)$ with only two free variables, which allows the treatment of monadically dependent and monadically stable classes via transductions, as presented in Section 6.2.
We can now define the VC-dimension, order-dimension and vc-density of formulas over structures in the natural way. For example, the vc-density of a formula is defined as follows.
Definition 8.11.
The vc-density of a formula $\varphi(\bar{x},\bar{y})$ over a class of structures $\mathscr{C}$ is the infimum of all reals $\alpha>0$ such that $|S^{\varphi}(\mathbb{A}/B)|\in\mathcal{O}(|B|^{\alpha})$, for all $\mathbb{A}\in\mathscr{C}$ and all finite $B\subset A$.
Using the locality of first-order logic, uniform quasi-wideness and neighborhood complexity, it was proved by Pilipczuk et al. [130] that the vc-density of first-order formulas in nowhere dense classes is very small.
Theorem 8.12 ([130]).
Let $\mathscr{C}$ be a class of graphs and let $\varphi(\bar{x},\bar{y})$ be a first-order formula.
  1. (1)
    If $\mathscr{C}$ is nowhere dense, then for every $\varepsilon>0$ there exists a constant $c$ such that for every $G\in\mathscr{C}$ and every nonempty $A\subseteq V(G)$, we have $|S^{\varphi}(G/A)|\leq c\cdot|A|^{|\bar{x}|+\varepsilon}.$
  2. (2)
    If $\mathscr{C}$ has bounded expansion, then there exists a constant $c$ such that for every $G\in\mathscr{C}$ and every nonempty $A\subseteq V(G)$, we have $|S^{\varphi}(G/A)|\leq c\cdot|A|^{|\bar{x}|}$.
As shown by Dreier et al. [37] as a key ingredient for their first-order model checking algorithm on monadically stable classes, also monadically stable classes behave very well.
Theorem 8.13 ([37]).
Let $\mathscr{C}$ be a monadically stable class of graphs and let $\varphi(x,y)$ be a first order formula with two free variables. Then for every $\varepsilon>0$ there exists a constant $c$ such that for every $G\in\mathscr{C}$ and every nonempty $A\subseteq V(G)$, we have $|S^{\varphi}(G/A)|\leq c\cdot|A|^{1+\varepsilon}.$
Dreier et al. [37] conjecture that the theorem generalizes to formulas $\varphi(\bar{x},\bar{y})$, just as Theorem 8.12 for nowhere dense classes. This conjecture would be implied by the conjecture that the notions of monadic stability and structural nowhere denseness coincide [125, 126]. They also conjecture that Theorem 8.13 generalizes to monadicallly dependent classes. This is known for some special cases, e.g. for classes with bounded twinwidth (even in the general setting of Theorem 8.12 and with the strong bounds as for classes with bounded expansion) as proved by Gajarský et al. [69].
8.2 Distance-r domination
A distance-$d$ dominating set in a graph $G$ is a set $D\subseteq V(G)$ such that every vertex of $G$ is at distance at most $d$ to a vertex of $D$. The algorithmic problem of deciding whether a graph admits a distance-$d$ dominating set of size at most $k$ is one of the benchmark problems on nowhere dense classes and many of the combinatorics we have seen so far have been developed to approach this algorithmic problem. Recall that a problem is fixed-parameter tractable with respect to a parameter $k$ if it can be solved in time $f(k)\cdot n^{\mathcal{O}(1)}$ for some computable function $f$. It admits a kernel of size $g(k)$ if one can compute in polynomial time an equivalent instance of size $g(k)$.
After Nešetřil and Ossona de Mendez [110] demonstrated first that the dominating set problem is fixed-parameter tractable based on low treedepth decompositions, Dawar and Kreutzer [30] extended these results to nowhere dense classes and the distance-$d$ dominating set problem. They introduced the notion of domination cores and developed an irrelevant vertex technique that was highly relevant for the further algorithmic development of nowhere dense classes based on uniform quasi-wideness. A $k$-domination core is a set $C\subseteq V(G)$ such that every set of size at most $k$ that dominates $C$ also dominates the whole graph $G$. Obviously, the existence of small domination cores is highly useful for the kernelization of the dominating set problem. The next step by Drange et al. was to establish linear kernels for distance-$d$ domination for bounded expansion classes based on neighborhood complexity combinatorics [35]. With the development of polynomial bounds for uniform quasi-wideness functions by Kreutzer et al. [97] first a polynomial kernel [97] and then an almost linear kernel for distance-$d$ domination on nowhere dense classes was developed by Eickmeyer et al. [56]. Telle and Villanger proved that small domination cores exist even for biclique-free graphs [146] and most generally for semi-ladder-free graphs as proved by Fabianski et al. [62] (however, these can only be computed in fpt time). Polynomial kernels for the independent set and dominating set problem in powers of nowhere dense classes can be efficiently computed as shown by Dreier et al. [39]. A unified view on kernelization for packing and covering problems in nowhere dense classes is provided by Ahn et al. [4] and by Einarson and Reidl [57]. The weak coloring numbers also play a key role in the constant factor approximation algorithms for the distance-$d$ dominating set problem [5, 50, 51, 98, 128]
8.3 Neighborhood covers
The dual of a set system $\mathcal{F}$ on a universe $X$ is the set system $\mathcal{F}^{*}$ that has one element for each set of $\mathcal{F}$ and contains the sets $S_{x}=\{A\in\mathcal{F}\mid x\in A\}$ for all $x\in X$. Intuitively, if we represent the set system $\mathcal{F}$ by its incidence matrix, then we obtain the dual set system $\mathcal{F}^{*}$ by the transposed incidence matrix. The dual VC-dimension of $\mathcal{F}$ is the VC-dimension of $\mathcal{F}^{*}$. Observe that the set system $\{N_{r}(v)\mid v\in V(G)\}$ is equal to its dual, hence, its VC-dimension is equal to its dual VC-dimension. In general, if the VC-dimension of $\mathcal{F}$ is $d$, then its dual VC-dimension is less than $2^{d+1}$ [7]. The dual set system of $\mathcal{W}_{r,\pi}$ considered in Lemma 8.7 is the system $\mathcal{W}^{*}_{r,\pi}=\{X_{v}\mid v\in V(G)\}$, where $X_{v}=\{w\mid v\in\operatorname{WReach}_{r}[G,\pi,w]\}$. This system hence also has bounded VC-dimension in nowhere dense classes. Its most interesting property however is that it forms an $r$-neighborhood cover.
Definition 8.14.
An $r$-neighborhood cover of $G$ is a set $\mathcal{X}$ of vertex subsets of $G$, called clusters such that the $r$-neighborhood of every vertex is contained in some cluster $X\in\mathcal{X}$. The radius of a cluster $X$ is the radius of $G[X]$. The cover $\mathcal{X}$ is said to have radius at most $d$ if all its clusters have radius at most $d$. The degree of the cover is the maximum number of clusters that intersect at some vertex.
The existence of sparse neighborhood covers for nowhere dense classes was one of the key ingredients for the first-order model checking algorithm on these classes by Grohe et al. [77].
Theorem 8.15 ([77]).
Let $\pi$ be an order of $V(G)$. Then $\mathcal{W}_{2r,\pi}^{*}$ is an $r$-neighborhood cover of radius at most $2r$ and degree at most $\operatorname{wcol}_{2r}(G,\pi)$.
This result was vastly generalized by Dreier et al. [37], who proved that sparse neighborhood covers can be constructed from linear orders with small crossing number.
Definition 8.16.
Let $(U,\mathcal{F})$ be a set system and $\pi$ be a linear order of $U$. The crossing number of $X\in\mathcal{F}$ is the number of pairs $(u,v)$ of elements of $U$ such that $v$ is an immediate successor of $u$ in $\pi$ and exactly one of $u$ and $v$ belongs to $X$.
Theorem 8.17 ([37]).
Let $G$ be a graph and assume the set system of closed neighborhoods has crossing number at most $k$. Then $G$ admits a distance-$1$ neighborhood cover of radius at most $4$ and degree at most $k+1$.
Welzl proved that small vc-density leads to linear orders with small crossing number [156].
Theorem 8.18 ([156]).
Let $(U,\mathcal{F})$ be a set system with vc-density at most $d$, where $d>1$ is a real. Then there exists a linear order of $U$ with crossing number bounded by $\mathcal{O}(|U|^{1-1/d}\log|U|)$.
As the $r$th-power of every monadically stable class is again monadically stable one obtains the following corollary.
Theorem 8.19 ([37]).
Let $\mathscr{C}$ be a monadically stable class of graphs, let $r$ be a positive integer and let $\varepsilon>0$. Then $\mathscr{C}$ admits $r$-neighborhood covers with radius at most $4r$ and degree $\mathcal{O}(n^{\varepsilon})$.
Furthermore, the neighborhood covers of the theorem can be efficiently computed.
Together with the flipper-game characterization of monadically stable classes this leads to the efficient model checking algorithm of Dreier et al. [37, 40] on monadically stable classes of graphs.
We conclude this section with one final result that Braunfeld et al. used to show that monadically stable classes have quasi-bounded bounded shrubdepth decompositions [21].
Definition 8.20.
An induced distance-$r$ neighborhood cover of a graph $G$ is a set $\mathcal{X}$ of induced subgraphs of $G$, called clusters, such that the $r$-neighborhood of every vertex is contained in some $X\in\mathcal{X}$. The radius of $\mathcal{X}$ is the maximum radius of the connected components of the clusters of $\mathcal{X}$. The size of $\mathcal{X}$ is the cardinality of $\mathcal{X}$.
Note that the connected components of any induced distance-r neighborhood cover of radius $d$ and size $k$ yield a distance-r neighborhood cover with radius $d$ and degree $k$.
Theorem 8.21 ([21]).
Let $\mathscr{C}$ be a monadically stable class of graphs, let $r$ be a positive integer and let $\varepsilon>0$. Then $\mathscr{C}$ admits induced $r$-neighborhood covers with radius at most $24r+$ 2 and size $\mathcal{O}(n^{\varepsilon})$.
9 Further applications
We conclude our survey with a few selected applications of the generalized coloring numbers.
9.1 Exact distance colourings
We begin with an elegant result of van den Heuvel at al. [149] bounding the chromatic number of exact distance graphs by the weak coloring numbers. Exact distance graphs were introduced by Nešetřil and Ossona de Mendez in [111]. We refer to [111, 149] for more background on colorings of power graphs and exact distance graphs.
Definition 9.1.
For a graph $G$ and positive integer $p$, the exact distance-$p$ graph $G^{[\sharp p]}$ is the graph with vertex set $V(G)$ that has an edge between vertices $u$ and $v$ if and only if $u$ and $v$ have distance exactly $p$ in $G$.
For odd $p$, Nešetřil and Ossona de Mendez proved that for a class $\mathscr{C}$ with bounded expansion, the chromatic number of $G^{[\sharp p]}$ is bounded by an absolute constant for all $G\in\mathscr{C}$ [111]. We provide the elegant proof of the following result (in fact, [149] proves a stronger result in terms of a measure called $\mathrm{dcol}$ that we refrain from defining here).
Theorem 9.2 ([149]).
For every odd positive integer $p$ and graph $G$ we have $\chi(G^{[\sharp p]})\leq\operatorname{wcol}_{2p-1}^{*}(G)$.
Proof.
Let $\pi$ be a linear ordering minimizing the chromatic number of $G\langle\pi,2p-1\rangle$ and let $\rho$ be a corresponding coloring witnessing that $\operatorname{wcol}_{2p-1}^{*}(G)=:c$, hence any two vertices $u,v\in V(G)$ have different colors under $\rho$ whenever $u\in\operatorname{WReach}_{2p-1}[G,\pi,v]$ or $v\in\operatorname{WReach}_{2p-1}[G,\pi,u]$. For a vertex $v$ write $m(v)$ for the minimum vertex in the $\lfloor p/2\rfloor$-neighborhood of $v$ with respect to $\pi$. Define the coloring $\chi$ such that every vertex $v$ receives the color $\rho(m(v))$. We claim that $\chi$ is a proper coloring of $G^{[\sharp p]}$.
Consider an edge $uv\in E(G^{[\sharp p]})$. By definition, $u$ and $v$ have distance exactly $p$ in $G$ and the $\lfloor p/2\rfloor$-neighborhoods of $u$ and $v$ are disjoint. We have $m(u)\neq m(v)$ and there exists a path of length at most $12p-$ 1 between $m(u)$ and $m(v)$. Assume without loss of generality that $m(u)<_{\pi}m(v)$. Then $m(u)\in\operatorname{WReach}_{2p-1}[G,\pi,m(v)]$, hence, $u$ and $v$ receive different colors under $\chi$. ∎
9.2 Dimension of posets
[…]
A convenient way of representing a poset is to draw its Hasse diagram: Each element is drawn as a point in the plane such that if $x<y$, then $x$ is drawn below $y$. Then, for each relation $x<y$ that is not implied by transitivity (these are called the cover relations), we draw a $y$-monotone curve going from $x$ to $y$. The diagram defines a corresponding undirected graph, where edges correspond to pairs of elements in a cover relation. This graph is called the cover graph of the poset. The height of a poset is the maximum size of a chain in the poset (a set of pairwise comparable elements).
There is a large body of literature connecting poset dimension with graph structure. Streib and Trotter proved that for every fixed heigh $h>1$, posets of height $h$ with a planar cover graph have bounded dimension [145]. In general, Kelly proved that there is no function bounding the dimension posets by their height and a planar cover graph also does not imply a bound on the dimension of the poset [90]. In the sequel, Joret et al. showed that posets have dimension upper bounded by a function of their height if their cover graphs have bounded treewidth, bounded genus, or exclude an apex-graph as minor [86], exclude a minor [154] (Walczak) or topological minor [105] (Micek and Wiechert), or come from a class with bounded expansion [88] (Joret et al.). The following nice result was proved by Joret et al. [87].
Theorem 9.3 ([87]).
Let $P$ be a poset of height at most $h$ and let $G$ be its cover graph. Then $\mathrm{dim}(P)\leq 4^{\operatorname{wcol}_{3h-3}(G)}$.
This line of research culminated in the following characterization of monotone graph classes.
Theorem 9.4 ([87]).
Let $\mathscr{C}$ be a monotone class of graphs. Then $\mathscr{C}$ is nowhere dense if and only if for every integer $h>1$ and real $\varepsilon>0$, $n$-element posets of height at most $h$ whose cover graphs belong to $\mathscr{C}$ have dimension $\mathcal{O}(n^{\varepsilon})$.
9.3 Directed graphs
With the great success of treewidth and related measures for undirected graphs, researchers have attempted to define similar concepts for directed graphs, see e.g. [9, 11, 70, 84, 123, 138]. However, the algorithmic impact of directed width measures has by far not been as successful as for undirected graphs. For example, Ganian et al. [71] call a width measure for digraphs algorithmically useful if all MSO definable problems admit polynomial-time algorithms on digraphs of bounded width. However, Courcelle’s Theorem [26], stating that every MSO property can be decided in linear time on every class of bounded treewidth, under standard complexity theoretic assumptions cannot be extended to graph classes of unbounded treewidth that additionally satisfy certain mild closure conditions [71, 99]. Nevertheless, these lower bounds still allow the search for classes of digraphs that are substantially different from classes whose underlying undirected graphs have bounded treewidth or are nowhere dense, and that still admit efficient algorithms for individual problems.
Kreutzer and Tazari [100] introduced generalizations of bounded expansion and nowhere dense classes of graphs to the directed setting. Excluding some tournament (a directed graph with exactly one edge between each two vertices) as a bounded depth minor does not turn out to be the right approach. Nowhere crownful classes of digraphs are defined by excluding one-subdivided cliques (which interestingly are the forbidden patterns in monadically dependent classes), and turn out to have a nice combinatorial theory [100]. In particular, nowhere crownful classes can alternatively be characterized by directed uniform quasi-wideness, with immediate algorithmic applications. Classes with directed bounded expansion were studied in [98, 95]. In particular, these classes can be characterized by directed generalized coloring numbers.
10 Concluding remarks and acknowledgments
The generalized coloring numbers are one of the most influential notions in the theory of bounded expansion and nowhere dense graph classes. In this survey I have tried to give an overview over some of the most important results and applications. Of course the presented material reflects my own preferences and I hope that I have not upset anyone by missing their results.
In the text I have included open questions about the generalized coloring numbers in the respective context. I conclude with some further explicit questions. Of course, there are many more, especially in the newly developing and very exciting area of graph product theory and in the study of dense but structurally sparse graph classes.
  1. 1.
[…]
  2. 2.
    Give a direct proof for good quality weak coloring orders from the existence of bounded-size bounded treedepth decompositions, or from sparse neighborhood covers and winning the splitter game.
  3. 3.
    Find a good notion of a dense analog of the generalized coloring numbers.
This work is dedicated to Jarik Nešetřil. Jarik, I want to express my heartfelt thanks for your inspiration and unwavering positivity, which I have had the privilege to experience since the early days of my PhD studies. Your good mood and enthusiasm have always made it a pleasure to work with you and Patrice. I hope that we will continue to explore the beauty of combinatorics for many more years to come!
Thank you Patrice for proofreading the manuscript, for many pointers to the literature, many comments on improving constructions, the beautiful figures, and the great time we always have when working together!

### v3

@@
 For a positive integer $r$, the strong $r$-coloring number $\operatorname{col}_{r}(G)$ and weak $r$-coloring number $\operatorname{wcol}_{r}(G)$ were introduced by Kierstead and Yang [94], and the $r$-admissibility $\operatorname{adm}_{r}(G)$ was introduced by Dvořák [53].
-All of these generalizations rely on linear orders and certain back-connectivity properties (formal definition will be given below).
-Intuitively, the generalized coloring numbers capture structural properties not only of subgraphs of a graph, as degeneracy does, but also of its bounded-depth minors.
+All of these generalizations rely on linear orders of the vertices and certain back-connectivity properties (formal definition will be given below).
+Intuitively, the generalized coloring numbers capture structural properties not only of subgraphs of a graph, as degeneracy does, but also of its bounded-depth minors (see Figure 2).
 Bounded-depth minors are the key concept in the theory of bounded expansion and nowhere dense graph classes, introduced by Nešetřil and Ossona de Mendez.
@@
 Hence, the classification of tractable first-order model checking on monotone graph classes is complete and leaves the question of a classification of tractability on hereditary graph classes.
+Figure 3 shows an inclusion diagram of relevant graph classes and the current frontiers of tractable model checking.
 The deep connections of nowhere dense classes to classical model theory (stability theory) are currently guiding an exciting development of dense but structurally sparse graph classes.
@@
 [48], which was inspired by the techniques developed for the construction of low treedepth colorings for classes with excluded minors by Pilipczuk and Siebertz [127].
-Figure 3: The map of the (structurally) sparse universe.
+Figure 3: The map of the (structurally) sparse universe and tractable model checking.
 In this survey, we overview combinatorial and algorithmic applications of the generalized coloring numbers, emphasizing new developments in this area.
@@
 Treewidth, as a measure of the global connectivity of a graph, plays an important role in graph theory.
-Treewidth and tree decompositions were introduced by Robertson and Seymour [136] as part of their graph minors project, however, several strongly related or equivalent notions were invented independently [10, 79].
+It was originally introduced by Bertelè and Brioschi under the name of dimension [10].
+It was rediscovered by Halin [80], and again rediscovered by Robertson and Seymour [136] as part of their graph minors project.
 We refer to [13, 82] for surveys on various characterizations of treewidth and equivalent parameters and to [14, 29] for surveys on algorithmic applications.
 Definition 2.1.
-A tree decomposition of a graph $G$ is a pair $\mathcal{T}=(T,\mathrm{Bag})$, where $T$ is a tree and $\mathrm{Bag}\colon V(T)\to\mathcal{P}(V(G))$ is a mapping that assigns each node $x$ of $T$ to its bag $\mathrm{Bag}(x)\subseteq V(G)$ so that the following conditions are satisfied:
+A tree decomposition of a graph $G$ is a pair $\mathcal{T}=(T,\mathrm{Bag})$, where $T$ is a tree and $\mathrm{Bag}\colon V(T)\to\mathcal{P}(V(G))$ is a mapping that assigns each node $x$ of $T$ to its bag $\mathrm{Bag}(x)\subseteq V(G)$ so that the following conditions are satisfied (see Figure 4):
   1. 1.
@@
 A rooted forest $F$ is a disjoint union of rooted trees.
-We write $\leq_{F}$ for the partial order that coincides with $\leq_{T}$ on each rooted tree in $F$.
+We write $\leq_{F}$ for the partial order that coincides with $\leq_{T}$ on each rooted tree $T$ in $F$.
 Tree decompositions capture the global connectivity properties of graphs, as made precise next.
@@
 Any linearization of $\leq_{\mathcal{T}}$ has very nice properties, which we will exploit later.
-In fact, we can associate a width to any linear order of $V(G)$ and thereby provide a characterization of treewidth.
+In fact, we can associate a width to any linear order of $V(G)$ and thereby provide another characterization of treewidth.
 Definition 2.4.
@@
 Finally, we let $G^{\pi}_{\textit{fill}}\coloneqq G_{1}$.
-It is immediate from the definition that $\pi$ is a perfect elimination order of $G^{\pi}_{\textit{fill}}$.
+It is immediate from the definition that $\pi$ is a perfect elimination order of $G^{\pi}_{\textit{fill}}$ (see Figure 5).
 Figure 5: An elimination order and fill-in graph.
@@
 Let $\pi$ be a linear order of $V(G)$.
-We say that a vertex $u\in V(G)$ is strongly $r$-reachable with respect to $\pi$ from a vertex $v\in V(G)$ if $u\leq_{\pi}v$ and there exists a path $P$ (possibly of length 00) between $u$ and $v$ of length at most $r$ with $w>_{\pi}v$ for all internal vertices $w\in V(P)$.
+We say that a vertex $u\in V(G)$ is strongly $r$-reachable with respect to $\pi$ from a vertex $v\in V(G)$ if $u\leq_{\pi}v$ and there exists a path $P$ (possibly of length 00) between $u$ and $v$ of length at most $r$ with $w>_{\pi}v$ for all internal vertices $w\in V(P)$ (see Figure 6).
 The set of vertices strongly $r$-reachable by $v$ with respect to the order $\pi$ is denoted $\operatorname{SReach}_{r}[G,\pi,v]$.
@@
 Let $G$ be a graph.
-The treedepth $\operatorname{td}(G)$ of $G$ is the minimum depth of a rooted forest $F$ on the same vertex set as $G$ such that whenever $uv\in E(G)$, then $u\leq_{F}v$ or $v\leq_{F}u$.
+The treedepth $\operatorname{td}(G)$ of $G$ is the minimum depth of a rooted forest $F$ on the same vertex set as $G$ such that whenever $uv\in E(G)$, then $u\leq_{F}v$ or $v\leq_{F}u$ (see Figure 7).
 Figure 7: The treedepth of a path $P_{n}$ with $n$ vertices is $\lceil\log_{2}(n+1)\rceil$.
@@
 Let $\pi$ be a linear order of $V(G)$.
-We say that a vertex $u\in V(G)$ is weakly $r$-reachable with respect to $\pi$ from a vertex $v\in V(G)$ if $u\leq_{\pi}v$ and there exists a path $P$ (possibly of length 00) between $u$ and $v$ of length at most $r$ with $w>_{\pi}u$ for all internal vertices $w\in V(P)$.
+We say that a vertex $u\in V(G)$ is weakly $r$-reachable with respect to $\pi$ from a vertex $v\in V(G)$ if $u\leq_{\pi}v$ and there exists a path $P$ (possibly of length 00) between $u$ and $v$ of length at most $r$ with $w>_{\pi}u$ for all internal vertices $w\in V(P)$ (see Figure 8).
 The set of vertices weakly $r$-reachable by $v$ with respect to the order $\pi$ is denoted $\operatorname{WReach}_{r}[G,\pi,v]$.
@@
 Monadic second-order formulas are defined as first-order formulas, but further allow the use of monadic quantifiers $\exists X$ and $\forall X$, and of a membership atomic formula $x\in X$, where $x$ is a first-order variable and $X$ a monadic second-order variable.
-A variable $x$ not in the scope of a quantifier is a free variable (we do not consider formulas with free set second-order variables).
+A variable $x$ not in the scope of a quantifier is a free variable (we do not consider formulas with free second-order variables).
 A formula without free variables is a sentence.
@@
 Theorem 2.22 ([55]).
-For every FO-formula $\varphi(\bar{x})$ and every class $\mathscr{C}$ of colored graphs with bounded treedepth there exists a quantifier-free formula $\tilde{\varphi}(\bar{x})$ and a linear-time computable map $Y$ such that, for every $G\in\mathscr{C}$, $Y(G)$ is a guided expansion of $G$ such that for all tuples of vertices $\bar{v}$
+For every FO-formula $\varphi(\bar{x})$ and every class $\mathscr{C}$ of colored graphs with bounded treedepth there exists a quantifier-free formula $\tilde{\varphi}(\bar{x})$ and a linear time computable map $Y$ such that, for every $G\in\mathscr{C}$, $Y(G)$ is a guided expansion of $G$ such that for all tuples of vertices $\bar{v}$
 $G\models\varphi(\bar{v})\quad\iff\quad Y(G)\models\tilde{\varphi}(\bar{v}).$
@@
 In this section, we present lower and upper bounds for the generalized coloring numbers.
-It turns out that the numbers are intimately linked to the density of bounded-depth minors and topological minors.
+It turns out that the numbers are intimately linked to the density of bounded-depth minors and bounded-depth topological minors.
 These are the central notions in the theory of bounded expansion and nowhere dense graph classes, which were introduced by Nešetřil and Ossona de Mendez [112, 114].
@@
 3.1 Graph expansion and simple lower bounds
-The following notions of bounded depth minors and bounded depth topological minors are the fundamental definitions in the theory of bounded expansion and nowhere dense graph classes introduced by Nešetřil and Ossona de Mendez [114, 116].
+The following notions of bounded-depth minors and bounded-depth topological minors are the fundamental definitions in the theory of bounded expansion and nowhere dense graph classes introduced by Nešetřil and Ossona de Mendez [112, 114].
 Definition 3.1.
@@
 Hence, expansion and topological expansion are functionally equivalent.
-The key observation for the proof is that in the branch set of a high-degree vertex in a bounded depth minor model, we will always find a vertex of large degree.
+The key observation for the proof is that in the branch set of a high-degree vertex in a bounded-depth minor model, we will always find a vertex of large degree.
 We can use such a vertex as the branching vertex of a high-degree vertex in a topological minor model.
@@
 A class $\mathscr{C}$ of graphs is nowhere dense if there exists a function $t$ such that for every non-negative integer $r$ we have $\omega_{r}(\mathscr{C})\leq t(r)$.
-Classes of bounded expansion are classes of sparse graphs by definition. It turns out that also nowhere dense classes are sparse.
+Bounded-depth minors in classes of bounded expansion have constant edge density by definition. It turns out that also bounded-depth minors in nowhere dense classes have small edge density.
 Theorem 3.7 ([114]).
@@
 [108] that deciding whether a graph contains a $1$-subdivision of density at least $d$ for $d\geq 2$ is also NP-complete.
-Problem 3.16.
-Let $r\geq 2$ be a positive integer.
-Does there exist an algorithm that, given an $n$-vertex graph $G$, computes $\operatorname{col}_{r}(G)$ and $\operatorname{wcol}_{r}(G)$ in time $n^{f(r,\operatorname{col}_{r}(G))}$, or $n^{f(r,\operatorname{wcol}_{r}(G))}$, for some function $f$?
+It was recently proved by Breen-McKay et al.
+[22] that determining if a graph has weak or strong $r$-coloring number at most $k$ is para-NP-hard when parameterized by $k$ for all $r\geq 2$.
+This answers an earlier question by the author and under standard complexity assumptions implies that there does not exist an algorithm that, given an $n$-vertex graph $G$, computes $\operatorname{col}_{r}(G)$ and $\operatorname{wcol}_{r}(G)$ in time $n^{f(r,\operatorname{col}_{r}(G))}$, or $n^{f(r,\operatorname{wcol}_{r}(G))}$, for any computable function $f$.
+Moreover, they showed that there exists a constant $c$ such that it is NP-hard to approximate the generalized coloring numbers within a factor of $c$.
 How can it possibly be hard to compute the values $b_{r}(S_{i},v_{i})$?
@@
 The following problem is attributed to Joret and Wood in [61].
-Problem 3.20.
+Problem 3.19.
 Does there exist a polynomial $p(x,y)$ such that for all graphs $G$ and all positive integers $r$ we have $\operatorname{col}_{r}(G)\leq p(r,\nabla_{r}(G))$?
 The problem appears already in bounded degree graphs.
-Problem 3.21.
+Problem 3.20.
 Does there exist a polynomial $p(x,y,z)$ such that for all graphs $G$ with maximum degree $\Delta$ and all positive integers $r$ we have $\operatorname{col}_{r}(G)\leq p(r,\nabla_{r}(G),\Delta)$?
@@
     or $\min\{w(u,v),w(v,u)\}>r$ and $\min_{z\in V\setminus\{u,v\}}(w(z,u)+w(z,v))>r$.
-If $w$ is an $r$-fraternity function for every value of $r$, then $w$ is called a fraternity function.
+If $w$ is an $r$-fraternity function for every value of $r$, then $w$ is called a fraternity function (see Figure 12).
 Figure 12: A fraternity function on a path $P_{6}$. All missing pairs are assigned value $\infty$.
@@
 We define $w(u,v)=d+1$ if $(u,v)\in E(\vec{G}_{d+1})$ and $w(u,v)=\infty$ if $(v,u)\in E(\vec{G}_{d+1})$. Finally, we define $w(u,v)=\infty$ for all $u,v$ with $(u,v),(v,u)\not\in E(\vec{G}_{r})$.
-Observation 3.33.
+Observation 3.32.
 The function $w$ as defined above is an $r$-fraternity function.
@@
 The margin of a node $x$ is the set $\mathrm{Margin}(x)=\mathrm{Bag}(x)\setminus\mathrm{Adh}(x)$.
-We say that $\mathcal{T}$ is a tree decomposition over a class $\mathscr{D}$ of graphs if $\mathrm{Torso}(x)\in\mathscr{D}$ for every node $x$ of $T$.
+We say that $\mathcal{T}$ is a tree decomposition over a class $\mathscr{D}$ of graphs if $\mathrm{Torso}(x)\in\mathscr{D}$ for every node $x$ of $T$ (see Figure 13).
 Figure 13: A tree decomposition of a graph $G$. The torso of $x$ is obtained from the graph induced by $\mathrm{Bag}(x)$ by turning the adhesions into cliques.
@@
 We denote the minor of $G$ obtained from $G$ by contracting each $V_{i}$ to a single vertex $v_{i}$ by $G/\mathcal{P}$.
-We inherit all the above definitions for connected partitions by identifying a part $V_{i}$ with the vertex $v_{i}$ in (the ordered graph) $G/\mathcal{P}$.
+Note that we care about the order of the $V_{i}$, that is, we consider $G/\mathcal{P}$ as an ordered graph.
+We inherit the above definitions of simplicial veritces, perfect elimination orders, etc.
+for connected partitions by identifying a part $V_{i}$ with the vertex $v_{i}$ in $G/\mathcal{P}$.
+For example, we call $V_{i}$ simplicial if $v_{i}$ is simplicial in $G/\mathcal{P}$ and the width of $\mathcal{P}$ is the width of the order $(v_{1},\ldots,v_{\ell})$ of $G/\mathcal{P}$.
 With this definition, we derive for example the following lemma.
@@
 For a graph $G$ denote by $G^{(r)}$ the exact $r$-subdivision of $G$, that is, the graph obtained from $G$ by replacing every edge by a path of length $r+1$ (with $r$ vertices on it).
-Recall the question of Joret and Wood, formulated in Problem 3.20, whether classes of polynomial expansion have polynomial strong coloring numbers.
+Recall the question of Joret and Wood, formulated in Problem 3.19, whether classes of polynomial expansion have polynomial strong coloring numbers.
 Theorem 5.5 ([77]).
@@
 Definition 6.2.
-For a positive integer $p$, a $p$-treedepth coloring of a graph $G$ is a vertex coloring of $G$ such that the subgraph induced by any $i\leq p$ color classes has treedepth at most $i$.
+For a positive integer $p$, a $p$-treedepth coloring of a graph $G$ is a vertex coloring of $G$ such that the subgraph induced by any $i\leq p$ color classes has treedepth at most $i$ (see Figure 15).
 Figure 15: A $3$-centered coloring of a grid with $5$ colors.
@@
 Definition 7.11.
-A class $\mathscr{C}$ of graphs is uniformly quasi-wide if for every $m\in\mathbb{N}$ and every $r\in\mathbb{N}$ there exist numbers $N(m,r)$ and $s(r)$ such that the following holds.
+A class $\mathscr{C}$ of graphs is uniformly quasi-wide if for every $m\in\mathbb{N}$ and every $r\in\mathbb{N}$ there exist numbers $N(m,r)$ and $s(r)$ such that the following holds (see Figure 16).
 Let $G\in\mathscr{C}$ and let $A\subseteq V(G)$ with $|A|\geq N(m,r)$.
@@
 Definition 8.20.
-An induced distance-$r$ neighborhood cover of a graph $G$ is a set $\mathcal{X}$ of induced subgraphs of $G$, called clusters, such that the $r$-neighborhood of every vertex is contained in some $X\in\mathcal{X}$.
+An induced $r$-neighborhood cover of a graph $G$ is a set $\mathcal{X}$ of induced subgraphs of $G$, called clusters, such that the $r$-neighborhood of every vertex is contained in some $X\in\mathcal{X}$.
 The radius of $\mathcal{X}$ is the maximum radius of the connected components of the clusters of $\mathcal{X}$.
 The size of $\mathcal{X}$ is the cardinality of $\mathcal{X}$.
-Note that the connected components of any induced distance-r neighborhood cover of radius $d$ and size $k$ yield a distance-r neighborhood cover with radius $d$ and degree $k$.
+Note that the connected components of any induced $r$-neighborhood cover of radius $d$ and size $k$ yield an $r$-neighborhood cover with radius $d$ and degree $k$.
 Theorem 8.21 ([21]).
@@
 In the sequel, Joret et al.
-showed that posets have dimension upper bounded by a function of their height if their cover graphs have bounded treewidth, bounded genus, or exclude an apex-graph as minor [86], exclude a minor [154] (Walczak) or topological minor [105] (Micek and Wiechert), or come from a class with bounded expansion [88] (Joret et al.).
+showed that posets have dimension upper bounded by a function of their height if their cover graphs have bounded treewidth, bounded genus, or exclude an apex-graph as a minor [87], exclude a minor [154] (Walczak) or topological minor [106] (Micek and Wiechert), or come from a class with bounded expansion [89] (Joret et al.).
 The following nice result was proved by Joret et al.